### v1

# Flock: A Knowledge Graph Foundation Model via Learning on Random Walks

Abstract
We study the problem of zero-shot link prediction on knowledge graphs (KGs), which requires models to generalize over novel entities and novel relations. Knowledge graph foundation models (KGFMs) address this task by enforcing equivariance over both nodes and relations, learning from structural properties of nodes and relations, which are then transferable to novel graphs with similar structural properties. However, the conventional notion of deterministic equivariance imposes inherent limits on the expressive power of KGFMs, preventing them from distinguishing structurally similar but semantically distinct relations. To overcome this limitation, we introduce probabilistic node-relation equivariance, which preserves equivariance in distribution while incorporating a principled randomization to break symmetries during inference. Building on this principle, we present Flock, a KGFM that iteratively samples random walks, encodes them into sequences via a recording protocol, embeds them with a sequence model, and aggregates representations of nodes and relations via learned pooling. Crucially, Flock respects probabilistic node-relation equivariance and is a universal approximator for isomorphism-invariant link-level functions over KGs. Empirically, Flock perfectly solves our new diagnostic dataset Petals where current KGFMs fail, and achieves state-of-the-art performances on entity- and relation prediction tasks on 54 KGs from diverse domains.
1 Introduction
Knowledge graph foundation models (KGFMs) (Lee et al., 2023; Geng et al., 2023; Galkin et al., 2024; Zhang et al., 2024; Cui et al., 2024; Huang et al., 2025) aim to infer missing links over novel knowledge graphs (KGs) that are not part of the training graphs or domains. This task requires generalization to both unseen nodes and unseen relation types. To achieve this, KGFMs rely on learning node and relation invariants: structural properties of nodes and relations that are transferable across KGs even when their relational vocabularies differ. This inductive bias is formalized by Gao et al. (2023) as double-equivariance — equivariance under permutations of both entities and relations — and used as a core principle in the design of KGFMs in the literature.
$\mathsf{Luke}$$\mathsf{Yoda}$$\mathsf{HanSolo}$$\mathsf{Emperor}$$\mathsf{Jabba}$$\mathsf{DarthVader}$$\mathsf{Chewbacca}$$\mathsf{Leia}$
Figure 1: A KG representing characters’ relationships in Star Wars movies. Blue arrows indicate $\mathsf{like}$, red arrows – $\mathsf{dislike}$, and green arrows indicate relation ($\mathsf{friendWith}$).
Problem statement. In this work, we challenge the fundamental assumption of existing KGFMs dictated by strict equivariance: structural isomorphism of relations implies semantic equivalence. Consider, for example the KG from Figure 1, where the relations like and dislike are structurally isomorphic, and yet they represent semantically opposite relations. In this motivating example, any KGFM that computes relation invariants is forced to assign the same representation to both like and dislike — losing the ability to distinguish between two entities with opposite relationships. This expressiveness limitation is an architectural one and cannot be resolved through finetuning, which further limits the downstream use of existing KGFMs. This raises a central question: How to design KGFMs that are both expressive and have the right inductive bias for generalization?
Our approach. We propose a new approach for KGFMs, which relies on probabilistic node-relation equivariance as inductive bias. Instead of enforcing deterministic equivariance over nodes and relations, these KGFMs respect probabilistic node-relation equivariance. This relaxes the hard constraint that “structurally isomorphic relations must have identical representations”, and requires only that “the representations of structurally isomorphic relations need to be equivalent in distribution” over a model’s stochastic processes. This way, the model retains crucial inductive bias needed for generalizing across different KGs, while the stochasticity of each forward pass ensures that structurally identical but semantically distinct relations are assigned different representations, allowing the model to distinguish between them (Srinivasan & Ribeiro, 2020; Abboud et al., 2021).
Inspired by the success of models that learn probabilistic invariants via random walks (Perozzi et al., 2014; Grover & Leskovec, 2016; Nikolentzos & Vazirgiannis, 2020; Kim et al., 2025), we introduce Flock, a KGFM that inherently computes probabilistic node-relation invariants. Given a (potentially unseen) KG, and a query, in each iteration, Flock first samples a set of random walks over KG based on the query, noting down both encountered nodes and relations with a recording protocol. To ensure the model can generalize to unseen entities and relation types, we anonymize all nodes and relations, enforcing that Flock only learn from their structural roles. These anonymized sequences are then fed into a sequence processor, and the representations for each node and relation are aggregated via a consensus protocol. Finally, we construct per-query (triple) features from the aggregated entity and relation embeddings and input them into a binary classifier for link prediction.
Key findings and contributions. The design of Flock offers several key advantages over existing KGFMs. First, it entirely abandons the conventional two-stage process of encoding relations and node representations via two separate networks, and does not rely on message-passing at all, thereby avoiding the well-known expressivity limitations of MPNNs on KGs (Barceló et al., 2022; Huang et al., 2023; 2025). Second, Flock is a universal approximator (see Proposition 4.1), and thus can approximate every link-level function defined on KGs of any bounded size. Finally, Flock architecture inherently respects the principle of probabilistic node-relation equivariance, which enables a strong generalization capacity. Our experimental results over both entity prediction and relation prediction validate the strength of this approach, demonstrating that Flock consistently outperforms state-of-the-art KGFMs on existing benchmarks. Our contributions can be summarized as follows:
  • We highlight a limitation in existing KGFMs: their over-reliance on deterministic node–relation equivariance prevents them from distinguishing between structurally similar but semantically different relations, limiting their expressivity.
  • We introduce probabilistic node-relation equivariance, a property for KGFMs that ensures invariance only in distribution, thus balancing the model expressivity and generalization.
  • We propose Flock, a KGFM that respects probabilistic node-relation equivariance. Flock replaces the conventional two-stage, message-passing paradigm with a direct sequence encoding approach based on random walks, and acts as a universal approximator of link-level functions.
  • We validate our approach on both entity and relation prediction tasks across 54 diverse KGs, where Flock consistently achieves state-of-the-art performance over existing KGFMs. We further design a synthetic dataset Petals to confirm our theoretical results empirically.
All proofs of the theoretical results are provided in Appendix C. The code is available at https://github.com/jw9730/flock-pytorch.
2 Related work
Link prediction and KGFMs. Early methods for inferring missing links in KGs  (Bordes et al., 2013; Sun et al., 2019; Balazevic et al., 2019; Abboud et al., 2020; Schlichtkrull et al., 2018; Vashishth et al., 2020) rely on learned embeddings, hence operating in the transductive setting, incapable of generalizing to unseen entities or relation types. Later GNN-based approaches based on the labeling trick (Teru et al., 2020; Zhang et al., 2021) or conditional message passing (Zhu et al., 2021; 2023; Zhang & Yao, 2022; Zhang et al., 2023b; Huang et al., 2023), unlocked the node inductive scenario, while remaining restricted to a fixed relational vocabulary. KGFMs eliminate this restriction and enable node-relation inductive link prediction over both unseen nodes and relation types through the use of a two-stage process by first encoding relations and then nodes. The first examples of this paradigm are InGram (Lee et al., 2023) and ULTRA (Galkin et al., 2024). Their ideas were extended by TRIX (Zhang et al., 2024) to build a more expressive framework. KG-ICL (Cui et al., 2024) achieved full inductivity by combining in-context learning with node-relation tokenization. ISDEA (Gao et al., 2023) and MTDEA (Zhou et al., 2023) highlighted the benefits of equivariance over both nodes and relations. MOTIF (Huang et al., 2025) was proposed as a general KGFM framework, supported by a theoretical analysis of the expressive power of KGFMs. Our work further advances the field with the first stochastic KGFM, which is invariant in probability and provably more expressive than all the existing methods. Notably, Flock achieves universality without any form of message passing, instead relying on random walks and sequence models to encode both nodes and relations anonymously to ensure generalization.
Random walks for graph representations. Random walks have attracted a lot of attention in graph learning, due to their simplicity and ability to gather context from neighborhoods. DeepWalk (Perozzi et al., 2014) and node2vec (Grover & Leskovec, 2016) were among the first to explore the potential of random walks for producing graph embeddings, treating walks as analogues of sentences in natural language and processing them using skip-gram models. Nikolentzos & Vazirgiannis (2020) generated graph-level task predictions by executing joint random walks on direct products of graphs with their extracted subgraphs. CRaWL (Tönshoff et al., 2021) represents the input graph as a collection of random walks and processes them with a 1-dimensional convolutional NN. WalkLM (Tan et al., 2023) samples random walks from graphs with textual features, passing them to a language model for embedding generation. RWNN (Kim et al., 2025) and RUM (Wang & Cho, 2024) anonymize the extracted walks and process them with sequence models and RNNs, respectively. NeuralWalker (Chen et al., 2025) aggregates embeddings derived by encoding random walks into message passing layers.
Probabilistic invariance. Neural architectures that enforce invariance to specific transformations often exhibit more stable training and improved performance (Bronstein et al., 2021), but this inductive bias can reduce their expressivity by preventing the model from distinguishing non-equivalent inputs. In graph learning, this trade-off is exemplified by MPNNs, whose power is limited by the 1-WL test (Xu et al., 2019; Morris et al., 2019). Randomization has emerged as a solution, enhancing expressivity through techniques such as noise injection (Abboud et al., 2021), vertex dropping (Papp et al., 2021), subgraph sampling (Bevilacqua et al., 2022; Zhang et al., 2023a), dynamic rewiring (Finkelshtein et al., 2024), and random walks (Kim et al., 2025; Wang & Cho, 2024). Despite their stochasticity, such methods can remain probabilistically invariant, ensuring that equivalent inputs yield identical expected outputs, or even identical output distributions. We extend the notion of probabilistic invariance to KGs and prove that Flock satisfies invariance in distribution.
3 Preliminary
Knowledge graphs. A knowledge graph (KG) is a tuple $G=(V,E,R)$, where $V$ denotes the set of entities (nodes), $R$ the set of relation types, and $E\subseteq V\times R\times V$ the set of labeled edges (facts). A fact is written as $(h,r,t)$ (or $h\xrightarrow{r}t$ interchangeably) with $r\in R$ and $h,t\in V$. A (potential) link in $G$ is any triple $(h,r,t)$ in $V\times R\times V$, regardless of whether it is present in $E$. We denote by $R^{-1}$ the set of inverses of relations $R$, defined as $\{r^{-1}\mid r\in R\}$, and mean $r$ when writing $(r^{-1})^{-1}$.
Isomorphism. An isomorphism between two knowledge graphs $G=(V,E,R)$ and $G^{\prime}=(V^{\prime},E^{\prime},R^{\prime})$ is a pair of bijections $\mu=(\pi,\phi)$, where $\pi:V\to V^{\prime}$ and $\phi:R\to R^{\prime}$, such that a fact $(h,r,t)$ belongs to $E$ if and only if the fact $\mu((h,r,t))=(\pi(h),\phi(r),\pi(t))$ belongs to $E^{\prime}$. Two KGs are isomorphic if such a mapping exists, in which case we write $G\simeq G^{\prime}$.
Link invariance. In this work, we focus on link-invariant functions. Let $\omega$ be a function assigning to each KG $G=(V,E,R)\in\mathbb{K}_{n,m}$ a map $\omega(G):V\times R\times V\rightarrow\mathbb{R}^{d}$. We say that $\omega$ is link invariant if for every pair of isomorphic KGs $G,G^{\prime}\in\mathbb{K}_{n,m}$, every isomorphism $(\pi,\phi)$ from $G$ to $G^{\prime}$, and every link $(h,r,t)$ in $G$, we have $\omega(G)((h,r,t))=\omega(G^{\prime})((\pi(h),\phi(r),\pi(t))).$
Probabilistic invariance. Let $\mathbb{K}_{n,m}$ be the space of knowledge graphs with $n$ vertices and $m$ relation types. A stochastic KG model $\varphi$ can be viewed as a function that takes a KG $G$ as the input and returns a random variable $\varphi(G)$. Following Kim et al. (2025), we call $\varphi$ invariant in probability if
$\forall G,G^{\prime}\in\mathbb{K}_{n,m}:\qquad G\simeq G^{\prime}\implies\varphi(G)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\varphi(G^{\prime})$
i.e. the distributions of $\varphi(G)$ and $\varphi(G^{\prime})$ are equal. In particular, this implies $\mathbb{E}[\varphi(G)]=\mathbb{E}[\varphi(G^{\prime})]$.
4 Methodology
[…]
Figure 2: Overall pipeline of Flock. In each updating step, Flock samples random walks on the KG, anonymizes the encountered nodes and relations via a recording protocol, and feeds the sequences in a sequence processor to compute node and relation representations. A consensus protocol then pools them back to the original KG’s nodes and relations.
We present Flock, a KGFM respecting probabilistic node-relation invariance. Flock is a randomized function $X_{\theta}(\cdot)$ which takes as input a KG $G=(V,E,R)$ and a link prediction query $q$. We consider two types of queries: entity prediction $q=(h,r,?)$ and relation prediction $q=(h,?,t)$. Flock outputs a random variable $\hat{{\bf y}}\sim X_{\theta}(G,q)$ which is suited for the task at hand. For entity prediction, it outputs $\hat{{\bf y}}:V\to[0,1]$ such that a potential link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(t)\in[0,1]$. For relation prediction, it outputs $\hat{{\bf y}}:R\to[0,1]$ such that a link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(r)$. In practice, we often ensemble repeated predictions $\hat{\bf y}_{1},...,\hat{\bf y}_{P}\sim X_{\theta}(G,q),P\in\mathbb{N}$ via averaging.
We describe the architecture of Flock in Section 4.1 focused on four main components, and then analyze its theoretical properties in Section 4.2, showing universality and probabilistic equivariance.
4.1 Flock
Internally, Flock has two lookup tables of hidden states, ${\bf v}:V\to\mathbb{R}^{d}$ for entities and ${\bf r}:R\to\mathbb{R}^{d}$ for relations, respectively. At each forward pass, it starts from trained initializations of these states ${\bf v}^{(0)}(\cdot)\coloneqq{\bf v}_{0}$ and ${\bf r}^{(0)}(\cdot)\coloneqq{\bf r}_{0}$, and updates them iteratively ${\bf v}^{(i)},{\bf r}^{(i)}$ for $i\leq I$. Each update is done residually using a randomized function $U_{\theta_{i}}$:
${\bf v}^{(i+1)}\coloneqq{\bf v}^{(i)}+\Delta{\bf v},\quad{\bf r}^{(i+1)}\coloneqq{\bf r}^{(i)}+\Delta{\bf r},\quad(\Delta{\bf v},\Delta{\bf r})\sim{\rm update}_{\theta_{i}}({\bf v}^{(i)},{\bf r}^{(i)}).$ (1)
The final hidden states ${\bf v}^{(I)}:V\to\mathbb{R}^{d}$ and ${\bf r}^{(I)}:R\to\mathbb{R}^{d}$ are then processed by a binary classifier ${\rm head}:\mathbb{R}^{d}\to[0,1]$ to produce the output $\hat{\bf y}$ which is $V\to[0,1]$ or $R\to[0,1]$ depending on task.
We now describe the randomized ${\rm update}_{\theta}$. We drop $i$ for brevity. It consists of four components:
  1. 1.
    Random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $l$ on the input KG.
  2. 2.
    Recording protocol $w:\eta_{j}\mapsto{\bf z}_{j}$ transforms each walk into a graph-agnostic sequence.
  3. 3.
    Sequence processor $f_{\theta}:{\bf z}_{j}\mapsto{\bf h}_{j}$ processes each sequence independently, outputting features.
  4. 4.
    Consensus protocol $c:({\bf h}_{1:N},\eta_{1:N})\mapsto(\Delta{\bf v},\Delta{\bf r})$ collects features of all walks and decides state updates for each entity and relation type.
An overview is presented in Figure 2. We note that $w$, $f_{\theta}$, and $c$ are all deterministic, and the random walk is the only source of stochasticity. We now discuss the design choice for each. For the ease of exposition, we explain for entity prediction tasks $q=(h,r,?)$, but relation prediction is similar.
Random walks. In Flock, random walks are central in two ways: they rewrite the connectivity of nodes and relations as sequences, and support generalization via probabilistic equivariance.
Formally, the random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $l$ on KG $G$. Each random walk $\eta$ is a chain of random variables, written as:n as:
$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l}}v_{l},\quad v_{s}\in V,r_{s}\in R,(v_{s-1},r_{s},v_{s})\in E,$ (2)
where the underlying transition mechanism and $l$ are hyperparameters.
To support probabilistic equivariance, we ask the walk algorithm to be invariant in probability. We say $\eta$ is invariant in probability if for any $G\simeq H$ in $\mathbb{K}_{n,m}$ with isomorphism $(\pi,\phi)$ from $G$ to $H$:
$\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\cdots\xrightarrow{\phi(r_{l})}\pi(v_{l})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\cdots\xrightarrow{s_{l}}u_{l}$ (3)
where $v_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{l}}v_{l}$ and $u_{0}\xrightarrow{s_{1}}\cdots\xrightarrow{s_{l}}u_{l}$ follow the distributions of $\eta(G,l)$ and $\eta(H,l)$, respectively. In such case, the isomorphism $(\pi,\phi)$ yields a natural translation from walks in $G$ to $H$.
In Flock, we use a simple random walk algorithm, which we show to be invariant in probability, while working robustly in practice. Specifically, given a base walk count $n$, for entity prediction queries $(h,r,?)$, we use $3n$ walks and choose three types of start locations. We start $n$ walks at query node $h$ by fixing $v_{0}=h$, start $n$ walks by choosing a random relation $s$, then a random edge $(v,s,u)\in E$, and then fixing the first step as $v\xrightarrow{s}u$, and start the rest of $n$ walks at random nodes in $V$. For relation prediction queries $(h,?,t)$, we additionally start $n$ walks at the tail node $t$ by fixing $v_{0}=t$, sampling a total of $4n$ walks.
For the transition mechanism, we use uniform walks with non-backtracking, with minor modifications to handle directed multi-edges of KGs. Despite the simplicity, we find that this choice works well in practice, consistent with findings of prior work (Tönshoff et al., 2021; Kim et al., 2025).
We lastly discuss how to choose the base walk count $n$. While this is a fixed hyperparameter $n_{\rm train}$ at pretraining, we find that scaling it adaptively to input KG at test-time benefits size generalization. We thus propose test-time adaptation of walk counts, and use:
$n=n_{\rm train}\times{\rm harmonic}\ {\rm mean}\left(\frac{|V|}{|V|_{\rm train}},\frac{|E|}{|E|_{\rm train}}\right)$ (4)
where $|V|_{\rm train},|E|_{\rm train}$ are average numbers of nodes and edges in pretraining KGs, respectively. Intuitively, this scales $n$ proportionally to the sizes of test KGs relative to pretraining. In practice, we clamp $n$ to the nearest power of $2$ in an interval to keep GPU memory usage in a range.
Recording protocol.
While random walks provide a basis for invariant sequence representations of KGs, two issues remain: (1) They reveal nodes $v_{s}$ and relations $r_{s}$ specific to each KG which obstructs transferability to unseen KGs. (2) They do not offer a way to condition on current states of entities $\bf v$}, relations $\bf r$}, and the query $q=(h,r,?)$ as often done in KGFMs via the labeling trick.
The recording protocol $w:\eta_{j}\mapsto{\bf z}_{j}$ resolves this by transforming each walk into a graph-agnostic sequence that only leaves structural information. Motivated by prior works on node anonymization for invariance (Kim et al., 2025; Wang & Cho, 2024), we propose an extension called node-relation anonymization: reserve separate namespaces for nodes and relations, respectively, and assign unique names in the order of discovery. For example, with $1,2,3,$... for nodes and $\alpha,\beta,$... for relations:
$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}v_{2}\xrightarrow{r_{1}^{-1}}v_{0}\qquad\mapsto\qquad 1\xrightarrow{\alpha}2\xrightarrow{\beta}3\xrightarrow{\alpha^{-1}}1,$ (5)
where $(\cdot)^{-1}$ marks direction of a relation. The protocol additionally employs a simple conditioning on current states $({\bf v},{\bf r})$ and query $q=(h,r,?)$, completing the record $\bf z$} as follows:
$w:\eta\mapsto{\bf z}=(1,{\bf v}(v_{0}),{\bf 1}_{h}(v_{0}))\xrightarrow{\alpha,{\bf r}(r_{1}),{\bf 1}_{r}(r_{1})}(2,{\bf v}(v_{1}),{\bf 1}_{h}(v_{1}))\xrightarrow{\beta,{\bf r}(r_{2}),{\bf 1}_{r}(r_{2})}\cdots,$ (6)
with indicator functions ${\bf 1}_{h}(\cdot),{\bf 1}_{r}(\cdot)$ at $h$ and $r$, respectively. As we will show, the recording protocol keeps node-relation invariance by hiding nodes and relations while leaving their structural roles.
Sequence processor.
Now that the recordings $\bf z$} only encode structural information of KG, we can safely process them with an arbitrary neural network $f_{\theta}:{\bf z}\mapsto{\bf h}$ without the risk of losing invariance. Since $\bf z$} are sequences, we choose sequence networks to leverage their inductive bias. Specifically, we use bidirectional GRU (Cho et al., 2014) equipped with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU feedforward network (Shazeer, 2020), which provided robust results.
Given that $f_{\theta}$ is a sequence network, it is convenient to interpret its output $\bf h$} as positionally aligned with each step of the walk $\eta$ or record $\bf z$}. Specifically, for the example in Equation 6, we obtain:
$f_{\theta}:{\bf z}\mapsto{\bf h}=(\Delta{\bf v}_{0},a_{0})\xrightarrow{\Delta{\bf r}_{1},b_{1}}(\Delta{\bf v}_{1},a_{1})\xrightarrow{\Delta{\bf r}_{2},b_{2}}\cdots.$ (7)
where $\Delta{\bf v}_{s},\Delta{\bf r}_{s}\in\mathbb{R}^{h\times d_{h}}$ and $a_{s},b_{s}\in\mathbb{R}^{h}$ are the decoded outputs at each position using linear projections. Intuitively, $\Delta{\bf v}_{s},\Delta{\bf r}_{s}$ encode proposals of state updates for entities and relations by $f_{\theta}$, and $a_{s},b_{s}$ encode respective confidences of $f_{\theta}$ for the proposed updates. This separation is useful due to the localized, pure-structure nature of the recordings $\mathbf{z}$. If a random walk $\eta$ densely visited a cycle-like region and then terminated in a dangling manner, it is natural to assign more confidence to the cycle-like region of the structural encodings $\bf h$}, and less confidence to the dangling region.
Consensus protocol.
After sequence processing, we are left with a handful of state update proposals ${\bf h}_{1:N}$ from $f_{\theta}$, that are positionally aligned with random walks $\eta_{1:N}$ on KG $G=(V,E,R)$. The consensus protocol $c$ uses the information to decide final state updates $\Delta{\bf v}:V\to\mathbb{R}^{d}$ and $\Delta{\bf r}:R\to\mathbb{R}^{d}$.
Since $c$ can access how each $\Delta{\bf v}_{s}$ within ${\bf h}_{j}$ is associated to a node $v_{s}\in V$ (and how each $\Delta{\bf r}_{s}$ is associated to a relation $r_{s}\in R$) through the random walk $\eta_{j}$, a simple way to form a consensus is by finding all proposals $\{\Delta{\bf v}_{s}\}$ associated to each node $v$, and all $\{\Delta{\bf r}_{s}\}$ associated to each relation $r$, and take averages of these proposals. The drawback is that uninformative proposals from e.g., dangling regions of walks are not directly suppressed, and can affect the state updates.
We can leverage the confidences $a_{s},b_{s}$ from $f_{\theta}$ to alleviate this issue. For each node $v\in V$ or relation $r\in R$, we first find all respective associated pairs $\{(\Delta{\bf v}_{s},a_{s})\}$ or $\{(\Delta{\bf r}_{s},b_{s})\}$ of proposals and confidences, and compute a multi-head softmax-normalized weighted average:
$\Delta{\bf v}(v)\coloneqq\left[\sum\exp(a_{s})\odot\Delta{\bf v}_{s}\right]\oslash\sum\exp(a_{s})\quad\Delta{\bf r}(r)\coloneqq\left[\sum\exp(b_{s})\odot\Delta{\bf r}_{s}\right]\oslash\sum\exp(b_{s}),$
[…]
Again, we can show that the consensus protocol does not operate in a way specific to particular KGs, and hence retains node-relation equivariance.
4.2 Theoretical properties of Flock
Expressivity. Following the notion of probabilistic expressivity introduced by Abboud et al. (2021), we say that a Flock model $X_{\theta}$ is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ if for any link invariant $\varphi:\mathbb{K}_{n,m}\rightarrow(V\times R\times V\rightarrow[0,1])$ and any $\epsilon,\delta>0$, there exists a choice of the network parameters $\theta$ and the length of the sampled random walks $l$, such that:
${\mathbb{P}}(|\varphi(G)((h,r,t))-X_{\theta}(G,(h,r,?))(t)|<\epsilon)>1-\delta$
for all graphs $G=(V,E,R)\in\mathbb{K}_{n,m}$ and all links $(h,r,t)\in V\times R\times V$.
Proposition 4.1.
With a powerful enough sequence processor $f_{\theta}$, the Flock framework described above is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
Invariance. Despite the stochastic nature of our framework, beyond randomized node embeddings (Abboud et al., 2021), Flock can be provably guaranteed to satisfy probabilistic invariance:
Proposition 4.2.
Suppose that the walk sampling protocol $\eta$ is invariant in probability and both the recording protocol $w$ and the consensus protocol $c$ are invariant. Then, regardless of the choice of the deterministic sequence processor $f_{\theta}$, the corresponding Flock model is invariant in probability.
Moreover, the designs of Flock’s components provided earlier in this section satisfy the conditions of Proposition 4.2. Therefore, the suggested pipeline is indeed invariant in probability:
Proposition 4.3.
Any Flock model with components as outlined in this section, and detailed in Appendix B is invariant in probability.
5 Experiments
We evaluate Flock over a wide range of KGs with both entity and relation prediction tasks, aiming to answer the following questions:
  • Q1.
    Can Flock approximate functions that existing KGFMs cannot?
  • Q2.
    How does Flock generalize to unseen entities and relations compared to existing KGFMs?
  • Q3.
    How does performance scale with the pretraining graph mix and the size of test-time ensemble?
  • Q4.
    What is the impact of test-time adaptation of walk counts?
  • Q5.
    How do current KGFMs perform under noise injection, and how do they compare with Flock? (see Appendix F)
In addition to these experiments, we report the computational complexity of Flock in Appendix D and conduct a detailed scalability analysis in Appendix E. We report ensembled predictions by averaging $P$ independent stochastic forward passes to reduce variance. Please refer to the detailed per-dataset choices in Table 22.
5.1 Synthetic datasets experiments
[…]
Figure 3: Example KG from Petals. KGFMs with relational invariants must equate blue $r_{1}$ and red $r_{2}$, thus predicting the same scores for both dashed queries with $r_{0}$.
Setup. We construct a synthetic benchmark Petals to validate the limitations of KGFMs following the node-relation equivariance (Q1). The Petals benchmark contains $220$ instances, with each instance including the following elements: 1) a KG $G=(V,E,R)$, which consists of a ‘central’ node $s\in V$, a ‘stem’ $T\subset V$ with query relation $r_{0}\in R$, and multiple cyclic ‘petals’, each ‘colored’ with a different pair of relations in $R\setminus\{r_{0}\}$, 2) an entity prediction query $(h,r_{0},?)$ with $h\in\{s\}\cup T$, and 3) two candidate targets $t_{1}$ and $t_{2}$ from the same ‘petal’, located at the same distance from $s$. An example is shown in Figure 3. See Appendix A for more details.
Petals is designed such that each instance always admits non-trivial automorphisms, meaning that swapping relations occurring in the same ‘petal’ will result in an isomorphic KG. Consequently, any model computing relation invariants will not be able to distinguish between potential links $(s,r_{0},t_{1})$ and $(s,r_{0},t_{2})$. However, the samples are constructed so that these links are not isomorphic from the graph perspective, making them distinguishable for general link-invariant functions. We say a model successfully solves an instance if it can classify $(s,r_{0},t_{1})$ as true and $(s,r_{0},t_{2})$ as false, and we measure the accuracy in the experiment.
Table 1: Accuracy of KGFMs of our Petals benchmark.
Model Petals
Ultra 50%
Motif(${\mathcal{F}}_{\text{Path}}^{3}$) 50%
Trix 50%
Flock 100%
We trained Ultra (Galkin et al., 2024), $\textsc{Motif}({\mathcal{F}}_{\text{Path}}^{3})$ (Huang et al., 2025), Trix (Zhang et al., 2024), and Flock from scratch and validated them on the training instances.
Results. We present the results in Table 1. As expected, all existing KGFMs relying on learning deterministic relational invariants fail to distinguish between the candidate target triplets completely, achieving $50\%$ accuracy due to random guesses. In contrast, Flock succeeds on all considered instances, displaying that, while remaining invariant in probability, it can differentiate between non-isomorphic links, even with isomorphic relations.
5.2 Entity and relation prediction over knowledge graphs
Setup. We follow the protocol of Galkin et al. (2024); Zhang et al. (2024) and pretrain Flock on FB15k-237 (Toutanova & Chen, 2015), WN18RR (Dettmers et al., 2018), and CoDEx Medium (Safavi & Koutra, 2020). We then evaluate its zero-shot and finetuned inference performance with the test set of 54 KG datasets (see Table 15 for details). These KG datasets are extracted from diverse domains across three settings: inductive on nodes and relations (Inductive $e,r$), inductive on nodes (Inductive $e$), and transductive. Note that these settings differ only during finetuning setup; in zero-shot setup, all entities and relations are unseen. We choose two state-of-the-art KGFMs Ultra (Galkin et al., 2024) and Trix (Zhang et al., 2024) as baselines, since they are pretrained on the same pretraining graph mix to ensure a fair comparison. Following conventions in the literature (Zhu et al., 2021; Huang et al., 2023), for each triple $(h,r,t)$, we add the corresponding inverse triple $(t,r^{-1},h)$, where $r^{-1}$ is a fresh relation symbol. For entity prediction, we report both head and tail results for each triple on all datasets, except for three from Lv et al. (2020) where only tail results are available, following Zhang et al. (2024). For evaluation, we use the filtered ranking protocol (Bordes et al., 2013), reporting Mean Reciprocal Rank (MRR) and Hits@10 for entity prediction, and Hits@1 for relation prediction, as some datasets have fewer than $10$ relations. Detailed per-dataset results are shown in Appendix G.
Entity prediction results. We present the average entity prediction results in Table 2 (Q2). In the zero-shot setting, Flock consistently outperforms Ultra and Trix on all metrics, demonstrating its strong generalization on KGs over diverse domains. Notably, on Metafam (Zhou et al., 2023), a dataset designed to challenge models with conflicting and compositional relational patterns, Flock roughly doubles MRR over Ultra and achieves about a 40% MRR gain over Trix in zero-shot performance. These gains align with our hypothesis that probabilistic node–relation equivariance improves expressivity without sacrificing generalization. In the finetuning setting, we observe a similar pattern: Flock maintains a consistent improvement over all datasets except transductive splits, where the underlying KGs are generally larger. We hypothesize that this gap stems from random walk coverages. Unlike Ultra and Trix whose message passing guarantees a full $k$-hop neighborhood coverage over the queried node, Flock relies on sampling random walks, which may not fully cover the target nodes of interest.
Table 2: Average entity prediction MRR and Hits@10 over 54 KGs from distinct domains.
Inductive $e,r$ Inductive $e$ Transductive Total Avg Pretrained
Model (23 graphs) (18 graphs) (13 graphs) (54 graphs) (3 graphs)
MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10
Ultra (zero-shot) 0.345 0.513 0.431 0.566 0.312 0.458 0.366 0.518 - -
Trix (zero-shot) 0.368 0.540 0.455 0.592 0.339 0.500 0.390 0.548 - -
Flock (zero-shot) 0.369 0.554 0.456 0.604 0.340 0.509 0.391 0.560 - -
Ultra (finetuned) 0.397 0.556 0.440 0.582 0.379 0.543 0.408 0.562 0.407 0.568
Trix (finetuned) 0.401 0.556 0.459 0.595 0.390 0.558 0.418 0.569 0.415 0.564
Flock (finetuned) 0.417 0.576 0.473 0.619 0.383 0.544 0.427 0.582 0.415 0.561
Table 3: Average relation prediction MRR and Hits@1 over 54 KGs from distinct domains.
Inductive $e,r$ Inductive $e$ Transductive Total Avg Pretrained
Model (23 graphs) (18 graphs) (13 graphs) (54 graphs) (3 graphs)
MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1
Ultra (zero-shot) 0.785 0.691 0.714 0.590 0.629 0.507 0.724 0.613 - -
Trix (zero-shot) 0.842 0.770 0.756 0.611 0.752 0.647 0.792 0.687 - -
Flock (zero-shot) 0.898 0.846 0.864 0.782 0.873 0.813 0.881 0.817 - -
Ultra (finetuned) 0.823 0.741 0.716 0.591 0.707 0.608 0.759 0.659 0.876 0.817
Trix (finetuned) 0.850 0.785 0.759 0.615 0.785 0.693 0.804 0.706 0.879 0.797
Flock (finetuned) 0.929 0.889 0.887 0.808 0.897 0.844 0.907 0.851 0.977 0.959
Relation prediction results. Table 3 presents the average MRR and Hits@1 for relation prediction results (Q2). We observe that Flock substantially outperforms all existing KGFMs across all categories in the zero-shot setting, achieving an $11.2\%$ relative improvement in MRR compared to the best baseline Trix. Flock shows a further performance boost of $12.8\%$ in the finetuned setting. We attribute this huge performance gain to Flock’s joint encoding of entities and relations during the updating step using the sequence encoder, while existing KGFMs, Ultra and Trix, require separate updating steps for entities and relations. This joint updating mechanism yields more holistic representations of both entities and relations with minimal information loss.
5.3 Scaling behavior of Flock
Pretraining graph mixing scaling. To assess whether Flock benefits from more pretraining graph and data (Q3), we follow the setup of Galkin et al. (2024), and pretrain Flock on an increasing number of KGs. We then evaluate them on all $41$ inductive benchmarks for a fair comparison. We present the detailed pretraining graph mix in Table 19. As shown in Figure 4(a), Flock’s generalization improves consistently as the number of pretraining KGs increases, exhibiting clear scaling behavior, which is a core characteristic of being a foundation model.
(a) Zero-shot MRR vs. # pretraining graphs.
(b) Zero-shot MRR vs. # ensembled predictions.
Figure 4: Pretraining and test-time scaling of Flock on 41 inductive KG datasets.
Number of ensembled predictions. To assess how test-time ensemble size $P$ affects performance (Q3), we take the pretrained Flock and run zero-shot entity prediction on 41 inductive KGs by increasing the number of ensembled passes. As shown in Figure 4(b), performance improves from 1 to 8 passes and then begins to plateau beyond 12. This indicates a clear scaling behavior: larger ensembles provide a more accurate estimate of the underlying node and relation distributions.
5.4 Impact of Test-Time Adaptation of Walk Counts
Table 4: Zero-shot entity prediction of Flock with and without adaptive test-time walks. We show the average number of entities $|V|$, triples $|E|$, base walks $n$, MRR, and Hits@10.
Dataset split Statistics Flock Flock w/o Adap.
$|V|$ $|E|$ $n$ MRR Hits@10 $n$ MRR Hits@10
Inductive $e,r$ 5,303 10,838 28.40 0.369 0.554 128 0.357 0.551
Inductive $e$ 7,578 29,090 18.08 0.456 0.604 128 0.441 0.596
Transductive 47,810 387,491 214.15 0.340 0.509 128 0.334 0.493
Recall that we employ test-time adaptation of walk counts, which adaptively selects the base walk count $n$ based on the graph size, computed via the harmonic-mean rule shown in Equation 4 during inference. To answer Q4, we conduct an ablation study over this adaptive mechanism. Table 4 compares this adaptive setting with a fixed setting that uses $128$ base walks per sample for all datasets, matching the pretraining setup ($n_{\text{train}}=128$). As expected, the average selected base count $n$ is smaller on both inductive splits and larger on the transductive split, yet the adaptive mechanism improves performance across all settings. This is consistent with the intuition that adaptive $n$ scales up walks on larger KGs to improve coverage while allocating fewer walks on smaller KGs to reduce redundant visits; Flock maintains comparable visiting rates and coverage to those seen during pretraining, thereby producing representations closer to the pretraining distribution and resulting in consistent performance gains.
6 Conclusions
We introduced Flock, as a knowledge graph foundation model that respects probabilistic node-relation equivariance. Flock iteratively samples query-conditioned random walks from the input KG, records encountered nodes and relations via a recording protocol, and relies on a sequence processor and consensus protocol to obtain node and relation representations. We empirically evaluate Flock over 54 KGs across different domains for both entity and relation prediction tasks, demonstrating its superior zero-shot and finetuned performances. We further construct a synthetic dataset petals to validate our theoretical findings. One limitation is scalability (discussed in Appendix E): ensuring coverage of the sampled random walk in a large KG requires an extensive number of longer walks, which can quickly become computationally infeasible. A future direction is to develop approximation strategies (Chamberlain et al., 2023) that reduce the cost of random walk sampling while retaining Flock’s downstream performance.
Acknowledgment
This work is supported in part by the National Research Foundation of Korea grant (No.RS-2024-00351212) and “HPC support” project funded by the Korea government (MSIT and NIPA). MB is supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1.
Appendix A Details of the Petals benchmark
$b_{0}$$a^{(1)}_{2}$$a^{(1)}_{4}$$a^{(1)}_{1}$$a^{(1)}_{3}$$a^{(1)}_{5}$$a^{(2)}_{1}$$a^{(2)}_{3}$$a^{(2)}_{2}$$a^{(2)}_{4}$$a^{(2)}_{5}$$a^{(3)}_{2}$$a^{(3)}_{4}$$a^{(3)}_{1}$$a^{(3)}_{3}$$a^{(3)}_{5}$$a^{(4)}_{1}$$a^{(4)}_{3}$$a^{(4)}_{2}$$a^{(4)}_{4}$$a^{(4)}_{5}$$b_{1}$$b_{2}$$b_{3}$
Figure 5: An example of a graph from Petals with $c=4$, $l=2$ and $t=3$, and the associated link prediction instances (dashed). The relation types ‘red’, ‘blue’, ‘pink’ and ‘yellow’ are structurally isomorphic, hence become equated in the eyes of the existing KGFMs.
State-of-the-art knowledge graph foundation models (KGFMs) typically impose relational invariance. Formally, given two knowledge graphs $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$, if there exists an isomorphism $(\pi,\phi)$ from $G$ to $H$, then for any $r\in R$, the model enforces identical representations for $r$ and its image $\phi(r)\in R^{\prime}$. This design promotes generalization across different graphs, as it aligns analogous relations, but reduces expressivity within a single graph ($G=H$), where relations related by automorphisms are forced to be indistinguishable. Concretely, if an automorphism $(\pi,\phi)$ of $G$ maps $r_{1}$ to $r_{2}$, then the model must treat $r_{1}$ and $r_{2}$ as identical during inference. While some approaches mitigate this limitation via the labeling trick, assigning distinct embeddings to query-specific nodes and relations, this only isolates the queried relation type and does not resolve the underlying issue in general.
Motivated by this limitation, we introduce the Petals benchmark. Petals comprises 220 graphs, each paired with a link prediction query $(h,r,?)$ and a target set $\{t_{1},t_{2}\}$. While $t_{1}$ and $t_{2}$ are non-isomorphic, KGFMs enforcing relational invariance are unable to distinguish them, producing identical predictions. We empirically validate this property by evaluating the classification accuracy of marking $t_{1}$ as true and $t_{2}$ as false, reported in Table 1.
A.1 Structure of the studied KGs
Knowledge graphs in Petals follow a flower-like structure, parametrized by the number $c$ of ‘petals’, their length $l$ and the length $t$ of the ‘stem’ (see Figure 5 for visualization).
Vertices. Each ‘petal’ is a set $A^{(i)}$ of $2l+1$ vertices $A^{(i)}=\left\{a_{1}^{(i)},a_{2}^{(i)},\dots,a_{2l+1}^{(i)}\right\}$, while the stem $B$ consists of $t+1$ nodes $B=\{b_{0},b_{1},\dots,b_{t}\}$. The full set of entities is then:
$V=B\cup\bigcup_{i=1}^{c}A(i)=\{b_{0},b_{1},\dots,b_{t}\}\cup\left\{a_{j}^{(i)}\mid 1\leq i\leq c,1\leq j\leq 2l+1\right\}$
We call $b_{0}$ the ‘central’ node, as it is connected to every petal, as described below.
Edges. The nodes of the stem are connected in a consecutive manner by the same relation type $r_{0}$. Precisely, for each $i\in{1,\cdots,t}$, there exists an edge $(b_{i-1},r_{0},b_{i})$. Each petal $A^{(i)}$ is associated with two edge types $r^{(i)}_{1},r^{(i)}_{2}$, and is connected to the central node $b_{0}$ with links $\left(b_{0},r^{(i)}_{1},a_{1}^{(i)}\right)$ and $\left(b_{0},r^{(i)}_{2},a_{2}^{(i)}\right)$. The rest of the petal is connected with edges of type $r_{1}^{(i)}$ only, going from $a^{(i)}_{2j-1}$ to $a^{(i)}_{2j+1}$, and from $a^{(i)}_{2j}$ to $a^{(i)}_{2j+2}$. Finally, there are also edges linking $a^{(i)}_{2l-1}$ and $a^{(i)}_{2l}$ to $a^{(i)}_{2l+1}$. Therefore, the full set of edges can be characterized as:
$E=\left(\{(b_{i-1},r_{0},b_{i})\mid 1\leq i\leq t\}\right)$ $\cup\left(\bigcup_{i=1}^{c}\left\{(b_{0},r^{(i)}_{1},a^{(i)}_{1}),(b_{0},r^{(i)}_{2},a^{(i)}_{2})\right\}\right)$
$\cup\left(\bigcup_{i=1}^{c}\bigcup_{j=1}^{l-1}\left\{\!\left(a^{(i)}_{2j-1},r^{(i)}_{1},a^{(i)}_{2j+1}\right),\!\left(a^{(i)}_{2j},r^{(i)}_{1},a^{(i)}_{2j+2}\right)\right\}\right)$
$\cup\left(\bigcup_{i=1}^{c}\left\{\!\left(a^{(i)}_{2l-1},r^{(i)}_{1},a^{(i)}_{2l+1}\right),\!\left(a^{(i)}_{2l},r^{(i)}_{1},a^{(i)}_{2l+1}\right)\right\}\right)$
We select each of the types $r^{(i)}_{1}$ and $r^{(i)}_{2}$ from the set of considered relations $R=\{r_{1},\dots,r_{|R|}\}$ so that any relation-invariant model will equate all petals (i.e. so that for each pair of petals, there is an automorphism taking one to another). For instance, Figure 5 displays a cyclic pattern, in which $r^{(i)}_{2}=r^{(i+1)}_{1}$. Such symmetry causes all petals to be isomorphic, and leads to the inability of KGFMs to distinguish between the relations inside them.
Link prediction instances. Although the petals are isomorphic to each other, given the asymmetry of edge types from $b_{0}$ to $a^{(i)}_{1}$ and $a^{(i)}_{2}$, the nodes within a single petal generally can be distinguished. Therefore, for each graph with the structure as described above, we randomly sample one of the stem nodes $b_{s}$, and ask the link prediction query $(b_{s},r_{0},?)$. For the target nodes, we randomly select petal index $i$ and distance $j$ from the central node $b_{0}$, and consider the predictions for $a^{(i)}_{2j-1}$ and $a^{(i)}_{2j}$. For example, Figure 5 shows the case when $b_{s}=b_{0}$, $i=1$ and $j=1$, where the query is $(b_{0},r_{0},?)$ and we are interested in the scores for $a^{(1)}_{1}$ and $a^{(1)}_{2}$.
A.2 Parameters and generation
We construct Petals by manually designing 11 relation-assignment schemes that guarantee isomorphism across all petals. For each such selection, which already determines the number $c$ of petals, we generate 20 graphs corresponding to all combinations of $t\in\{1,2,3,4\}$ and $l\in\{1,2,3,4,5\}$. Each graph is paired with a link prediction query and two target nodes, sampled as described above. This yields $11\cdot 20=220$ instances that constitute the Petals benchmark.
Appendix B Methodology - details
In this section, we expand on the descriptions of individual components of Flock summarized in Section 4: the random walk algorithm, the recording protocol, the sequence processor, and the consensus protocol.
B.1 Uniform random walk
Let $G=(V,E,R)$ be a knowledge graph, and let $l$ be the length of random walks. For each node $v\in V$, we will denote by $\mathcal{N}(v)$ the set of neighbors of $v$:
$\mathcal{N}(v)=\{w\in V:\exists r\in R.(v,r,w)\in E\lor(w,r,v)\in E\}$
and by E(v,w)E{(v,w)}, the set of relational edges from $v$ to $w$ (allowing for the inverse direction):
$E{(v,w)}=$ $\{(v,r,w)\in R\times\{v\}\times\{w\}:(v,r,w)\in E\}$
$\cup\{(v,r^{-1},w)\in R^{-1}\times\{v\}\times\{w\}:(w,r,v)\in E\}$
where $R^{-1}$ is the set symbolizing the inverses of relation types in $R$. The uniform random walk with no backtracking $\eta(G,l)$ of length $l$ over $G$, represented as:
$V_{0}\xrightarrow{R_{1}}V_{1}\xrightarrow{R_{2}}\cdots\xrightarrow{R_{l}}V_{l}$
is a second-order Markov process that follows the rules:
[…]
for all $i\geq 0,j\geq 1$. Intuitively, at each step of the walk, we first select a neighbor (except for the vertex chosen one step ago) of the current node uniformly at random (disregarding multi-edges and edge directions), and then sample an edge between these two nodes uniformly at random. If the current node has only one neighbor, we are forced to return to it.
The initial conditions depend on the selected scenario. Given a query $q=(h,r,?)$ over $G$, we can describe the process of selecting the first step $V_{0}\xrightarrow{R_{1}}V_{1}$ as setting either (each with probability $\frac{1}{3}$):
  • $V_{0}=h$ and selecting the first step uniformly at random as described above, meaning:
[…]
  • setting $R_{1}=r$ and selecting $V_{0}\xrightarrow{R_{1}}V_{1}$ uniformly at random from edges with type $r$.
  • choosing $V_{0}$ uniformly at random, and then sampling the first step at random as well:
    ${\mathbb{P}}(V_{0}=w)$ $=\frac{1}{|V|}$
[…]
For the relation prediction objective, we add one more scenario, similar to the first one described above, but substituting $V_{0}=t$ instead. For that problem, each scenario is chosen with probability $\frac{1}{4}$.
B.2 Recording function
Given a KG $G=(V,E,R)$, a query $q=(h_{q},r_{q},?)$, a walk $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}$ of length $l$ over $G$, and a set of embeddings $\mathbf{v}$ of nodes $V$ and $\mathbf{r}$ of relations $R$, our recording function $w$ first splits the walk into a sequence of $l+1$ steps:
$(r_{0},v_{0}),(r_{1},v_{1}),\dots(r_{l},v_{l})$
with $r_{0}=r_{\varnothing}$ being a special marker for no relation. Each step $(r_{i},v_{i})$ is transformed into a 7-tuple:
$S_{i}=\left(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{i};\bar{\eta}),\operatorname{dir}_{i},\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}},\mathbf{v}(v_{i}),\mathbf{r}(r_{i})\right)$
where:
  • $\operatorname{id}_{V}(v_{i};\bar{\eta})$ and $\operatorname{id}_{R}(r_{i};\bar{\eta})$ are the anonymized id’s of the node $v_{i}$ and relation $r_{i}$, evaluated as:
    $\operatorname{id}_{V}(v_{i};\bar{\eta})$ $=\operatorname*{arg\,min}_{t}[v_{t}=v_{i}]$
    $\operatorname{id}_{R}(r_{i};\bar{\eta})$ $=\operatorname*{arg\,min}_{t}\left[r_{t}=r_{i}\lor r_{t}=r_{i}^{-1}\right]$
  • $\operatorname{dir}_{i}$ denotes the direction in which we follow the edge. We set $\operatorname{dir}_{i}\!=\!0$ if $r_{i}\!\in\!R$ (the edge is traversed from head to tail) and $\operatorname{dir}_{i}=1$ if $r_{i}\in R^{-1}$ (the edge is taken in the reverse direction).
  • $\delta_{v_{i}=h_{q}}$ and $\delta_{r_{i}\sim r_{q}}$ are binary flags representing whether the current node $v_{i}$ is the query head $v_{q}$ and if the relation $r_{i}$ is either the queried relation $r_{q}$ or its inverse $r_{q}^{-1}$.
  • $\mathbf{v}(v_{i}),\mathbf{r}(r_{i})$ are the embeddings of $v_{i}$ and $r_{i}$, respectively.
The output of $w$ for $\bar{\eta}$ given $G,q,\mathbf{v},\mathbf{r}$ is then:
$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{l})$
B.3 Sequence processor
Once the sampled walks are anonymized by the recording protocol $w$, the output for each walk $\bar{\eta}_{i}$:
$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{l})$
is passed through the sequence processor $f_{\theta}$, parametrized by the following modules:
  • $\mathbf{A}_{v},\mathbf{A}_{r}\in\mathbb{R}^{(l+1)\times d}$: embedding tables for anonymized vertices and relations, respectively,
  • $\mathbf{D}\in\mathbb{R}^{2\times d}$: look-up table for the direction embedding,
  • $\mathbf{Q}_{h},\mathbf{Q}_{r}\in\mathbb{R}^{2\times d}$: embedding tables for the binary query labels,
  • $\mathbf{V},\mathbf{R}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$: linear maps applied to the passed embeddings of vertices and relations,
  • $\mathbf{\Omega}$: a bi-directional GRU (Cho et al., 2014) cell equipped with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU (Shazeer, 2020) activation function.
For each step, encoding $S_{i}$ of the form:
$S_{i}=\left(\operatorname{id}_{V}(v_{i};\bar{\eta}_{i}),\operatorname{id}_{R}(r_{i};\bar{\eta}_{i}),\operatorname{dir}_{i},\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}},\mathbf{v}(v_{i}),\mathbf{r}(r_{i})\right)$
we evaluate the processed embedding $\mathbf{c_{i}}$ of $S_{i}$ as a sum of the corresponding encoded components:
$\mathbf{c}_{i}=$ $\mathbf{A}_{v}(\operatorname{id}_{V}(v_{i};\bar{\eta}_{i}))+\mathbf{A}_{r}(\operatorname{id}_{R}(r_{i};\bar{\eta}_{i}))+\mathbf{D}(\operatorname{dir}_{i})$
$+\mathbf{Q_{h}}(\delta_{v_{i}=h_{q}})+\mathbf{Q}_{r}(\delta_{r_{i}=r_{q}})+\mathbf{V}(\mathbf{v}(v_{i}))+\mathbf{R}(\mathbf{r}(r_{i}))$
These are then passed to the GRU cell $\mathbf{\Omega}$, which fuses the features across the whole walk and produces multi-head embeddings of vertices and relations, as well as the associated weights:
$\left(\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i},\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\right)=\mathbf{\Omega}([\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{l}])$
where $\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i}\in\mathbb{R}^{(l+1)\times h\times d_{h}}$ and $\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\in\mathbb{R}^{(l+1)\times h}$. Stacking all $N$ of them gives us the final output of the sequence processor.
B.4 Consensus protocol
Given walks $\bar{\eta}_{1:N}$ over $G=(V,E,R)$ and the outputs $\mathbf{s}^{(V)},\mathbf{s}^{(R)},\mathbf{a}^{(V)},\mathbf{a}^{(R)}$ of the sequence processor, the consensus protocol $c$ aggregates the signal for each node by evaluating a weighted sum over the appearances of this node across the walks. More precisely, for each node $v\in V$, we find all pairs of indices $(i,j)$, such that the $j^{\text{th}}$ node visited in $\bar{\eta}_{i}$ was $v$, and concatenate the weighted sums of embeddings produced by each head, with weights exponentially proportional to the scores $\mathbf{a}^{(V)}$:
$\Delta\mathbf{v}(v)=\bigoplus_{k=1}^{h}\frac{\sum\limits_{\begin{subarray}{c}i,j\\ \bar{\eta}_{i}(v_{j})=v\end{subarray}}\exp\left(\mathbf{a}^{(V)}_{i,j,k}\right)\cdot\mathbf{s}^{(V)}_{i,j,k}}{\sum\limits_{\begin{subarray}{c}i,j\\ \bar{\eta}_{i}(v_{j})=v\end{subarray}}\exp\left(\mathbf{a}^{(V)}_{i,j,k}\right)}$
Similarly, we aggregate the encodings for relations, considering their occurrences in both directions:
$\Delta\mathbf{r}(r)=\bigoplus_{k=1}^{h}\frac{\sum\limits_{\begin{subarray}{c}i,j\\ \bar{\eta}_{i}(r_{j})\in\{r,r^{-1}\}\end{subarray}}\exp\left(\mathbf{a}^{(R)}_{i,j,k}\right)\cdot\mathbf{s}^{(R)}_{i,j,k}}{\sum\limits_{\begin{subarray}{c}i,j\\ \bar{\eta}_{i}(r_{j})\in\{r,r^{-1}\}\end{subarray}}\exp\left(\mathbf{a}^{(R)}_{i,j,k}\right)}$
In both formulas above, $\bigoplus$ denotes concatenation.
Additionally, we say that a consensus protocol $c$ is invariant if for any pair of isomorphic KGs $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$, any isomorphism $\mu=(\pi,\phi)$ from $G$ to $H$, any list of embeddings $\mathbf{h}_{1:N}$ with $\mathbf{h}_{i}\in\mathbb{R}^{d}$, and any sequence of sampled walks $\bar{\eta}_{1:N}$ over $G$, the outputs
$(\Delta\mathbf{v},\Delta\mathbf{r})$ $=c(\mathbf{h}_{1:N},\bar{\eta}_{1:N})$
$(\Delta\mathbf{v}^{\prime},\Delta\mathbf{r}^{\prime})$ $=c(\mathbf{h}_{1:N},\mu(\bar{\eta}_{1:N}))$
satisfy:
$\Delta\mathbf{v}(v)$ $=\Delta\mathbf{v}^{\prime}(\pi(v))\quad$ $\forall v\in V$
$\Delta\mathbf{r}(r)$ $=\Delta\mathbf{r}^{\prime}(\phi(r))\quad$ $\forall r\in R$
Appendix C Proofs
C.1 Expressivity
The main proposition of this section formalizes the fact that Flock can approximate any link-invariant function over fixed-size knowledge graphs in probability. Intuitively, when the length of the sampled walks $l$ becomes higher, the probability of a single walk witnessing all the edges grows to 1. Once a walk visits all the edges, a sufficiently powerful sequence processor can derive the whole graph structure from its anonymized representation, recreating the graph in its entirety, up to isomorphism. Then, the processor can return the value of the approximated function for that graph.
We start by showing that the edge cover time $C_{E}(\cdot)$ of graphs in $K_{n,m}$ is bounded:
Lemma C.1.
Let $G\in\mathbb{K}_{n,m}$ for some $n,m$. The edge cover time $C_{E}(G)$ of $G$, using the algorithm from Section B.1, is finite.
Proof.
Let $G=(V,E,R)\in\mathbb{K}_{n,m}$ be a graph. For any edge $e\in E$ and any vertex $v\in V$, let $H_{v}(e)$ denote the expected number of steps of the random walk algorithm $\eta$ described in Section B.1. Then, the edge cover time $C_{E}(G)$ of $G$ with $\eta$, i.e. the expected number of steps that $\eta$ needs to take before visiting every edge in $G$, is bounded above by:
$C_{E}(G)\leq\sum\limits_{e\in E}\max_{v\in V}H_{v}(e)\leq m\cdot\max_{\begin{subarray}{c}e\in E\\ v\in V\end{subarray}}H_{v}(e)$
Indeed, consider the event of visiting all these edges in order $e_{1},\dots,e_{m}$:
$C_{E}(G)$ $=\mathbb{E}[\text{\#steps to visit all }e_{1},\dots,e_{m}]$
$\leq\mathbb{E}[\text{\#steps to visit }e_{1},\text{ then }e_{2},\dots,\text{ then }e_{m}]$
[…]
$\leq\max_{v\in V}H_{v}(e_{1})+\sum_{i=1}^{m-1}\max(H_{h_{i}}(e_{i+1}),H_{t_{i}}(e_{i+1}))$
$\leq\max_{v\in V}H_{v}(e_{1})+\sum_{i=1}^{m-1}\max_{v\in V}H_{v}(e_{i+1})$
$=\sum_{i=1}^{m}\max_{v\in V}H_{v}(e_{i})$
where $h_{i}$ and $t_{i}$ are the head and tail of the edge $e_{i}$, respectively. Therefore, to show that $C_{E}(G)$ is finite, it suffices to prove that $H_{v}(e)$ is bounded for all $v\in V,e\in E$.
Fix $v\in V$ and $e\in E$. Consider an infinite random walk generated with $\eta$ over $G$, starting at $v$:
$v=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}v_{2}\xrightarrow{r_{3}}\dots$
We want to bound the expected first index $t$, such that $e$ is the edge traversed in step $v_{t-1}\xrightarrow{r_{t}}v_{t}$. Denote by $\Delta$ a maximum degree of a vertex in $G$ (counted as the number of connected vertices $\mathcal{N}(v)$), by $\rho$ the maximum number of edges between any single pair of nodes and by $d$ – the diameter of the graph, i.e. the length of the longest shortest path between two vertices (in the undirected version of $G$). Consider the series of events $A_{0},A_{1},\dots$ where $A_{i}$ is characterized as:
[…]
Let $e=(h_{e},r_{e},t_{e})$. For all values of $i$, by definition, the length of the shortest path from $v_{i(d+2)}$ to $h_{e}$ or $t_{e}$ is at most $d$. Therefore, the whole part of the walk described in $A_{i}$ is at most $d+1$ steps long. By the definition of the used random walk algorithm, which only looks at the previously taken edge, we can deduce that the events $A_{i}$ are all mutually independent.
Moreover, let $v_{i(d+2)}=u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\dots\xrightarrow{s_{l}}u_{l}\in\{h_{e},t_{e}\}$ be a shortest path from $v_{i(d+2)}$ to one of $h_{e},t_{e}$. Note that by minimality, there cannot be any backtracking while following this path. Therefore, the probability of the next visited node is dependent only on the value of the previous one, and we can bound the probability $P(A_{i})$ of $A_{i}$ from below by:
${\mathbb{P}}(A_{i})\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{l-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
The first term on the right hand side is the probability of selecting $e$ while being at $h_{e}$ or $t_{e}$, which is the probability of first selecting the other endpoint (out of at most $\Delta$ neighbors) and then picking $e$ over other edges between $h_{e}$ and $t_{e}$ (of which there is at most $\rho$). Hence:
${\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\geq\frac{1}{\Delta}\cdot\frac{1}{\rho}=\frac{1}{\Delta\cdot\rho}$
As we never reach a backtracking situation by minimality of the shortest path, we can also write:
${\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})=\frac{1}{|\mathcal{N}(v_{i(d+2)+j})|}\geq\frac{1}{\Delta}$
Combining these observations, we can derive a bound for ${\mathbb{P}}(A_{i})$ in terms of $\Delta,\rho$ and $d$:
${\mathbb{P}}(A_{i})$ $\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{l-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
$\geq\frac{1}{\Delta\cdot\rho}\cdot\prod_{j=0}^{l-1}\frac{1}{\Delta}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{l}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{d}\quad=\quad\frac{1}{\rho\Delta^{d+1}}$
Finally, note that if $A_{i}$ is true, then the first index $t$ such that $v_{t-1}\xrightarrow{r_{t}}v_{t}$ traverses $e$ is at most $(i+1)(d+2)$. We can therefore bound the expectation of such $t$, being $H_{v}(e)=H_{v_{0}}(e)$ by:
$H_{v}(e)$ $\leq(d+2)\cdot{\mathbb{P}}(A_{0})+2(d+2)\cdot{\mathbb{P}}(\neg A_{0}\land A_{1})+3(d+2)\cdot{\mathbb{P}}(\neg A_{0}\land\neg A_{1}\land A_{2})+\dots$
$=(d+2)\cdot{\mathbb{P}}(A_{0})+2(d+2)\cdot{\mathbb{P}}(\neg A_{0})\cdot P(A_{1})+3(d+2)\cdot{\mathbb{P}}(\neg A_{0})\cdot{\mathbb{P}}(\neg A_{1})\cdot{\mathbb{P}}(\land A_{2})+\dots$
$=(d+2)+{\mathbb{P}}(\neg A_{0})\cdot(d+2+{\mathbb{P}}(\neg A_{1})\cdot(d+2+{\mathbb{P}}(\neg A_{2})\cdot(\dots)))$
$\leq(d+2)+\left(1-\frac{1}{\rho\Delta^{d+1}}\right)\cdot\left(d+2+\left(1-\frac{1}{\rho\Delta^{d+1}}\right)\cdot\left(d+2+\left(1-\frac{1}{\rho\Delta^{d+1}}\right)\cdot(\dots)\right)\right)$
$=\rho(d+2)\Delta^{d+1}$
Since $\rho\leq m,d+2\leq n$ and $\Delta\leq n$, we have $H_{v}(e)\leq m(n+2)n^{n}$, which completes the proof. ∎
Remark C.2.
The bound obtained in the proof of Lemma C.1 is very crude. In fact, we could transform the given knowledge graph into a simple graph (undirected, with no multi-edges) by substituting each edge $u\xrightarrow{r}v$ with two undirected edges $u\leftrightarrow v_{(u,r,v)}\leftrightarrow v$. The augmented graph will then have $n+m$ vertices, and our random walk algorithm naturally translates to a weighted random walk on the transformed graph. This hints at an assumption that in practice, the edge cover time of the used random walk algorithm is of the magnitude $O((n+m)^{3})=O(n^{3}+m^{3})$.
Let us now prove a fact about the number of distinct, up to isomorphism, graphs in $\mathbb{K}_{n,m}$.
Lemma C.3.
For any $n,m$, the number of isomorphism classes in $\mathbb{K}_{n,m}$ is finite.
Proof.
Since the number of distinct relation types that a graph in $\mathbb{K}_{n,m}$ is at most $m$, it suffices to show that the number of isomorphism classes of graphs in $\mathbb{K}_{n,m}$ with exactly $k$ relation types is bounded, for all $k\in\{1,2,\dots,m\}$.
Fix the number $k\in\{1,2,\dots,m\}$ and consider $G=(V,E,R)\in\mathbb{K}_{n,m}$ with $|R|=k$. We will show that, up to isomorphism, there are finitely many such choices of $G$. Firstly, as renaming does not change the graph structure, without loss of generality, we can assume that:
$V=\{v_{1},v_{2},\dots,v_{n}\}\qquad\text{and}\qquad R=\{r_{1},r_{2},\dots,r_{k}\}$
Then, there are exactly $n^{2}k$ possible relational edges $e\in(V\times R\times V)$, and $E\subseteq V\times R\times V$ is a set of $m$ elements. Hence, there are $n^{2}k\choose{m}$ possible choices of $E$, and hence, at most $n^{2}k\choose{m}$ non-isomorphic choices of $G$. Since $k$ was chosen arbitrarily, this completes the proof. ∎
Lemma C.4.
For each pair $(n,m)$, there exists a number $C_{n,m}$ such that the edge cover time, using the algorithm from Section B.1, of any knowledge graph in $\mathbb{K}_{n,m}$ is at most $C_{n,m}$.
Proof.
The result follows from Lemmas C.1 and C.3. As two isomorphic graphs have identical cover time, we can set $C_{n,m}$ to be the maximum of cover times of representatives of all isomorphic classes, which, by finiteness of both, is well-defined. ∎
Lemma C.5.
Let $G\in\mathbb{K}_{n,m}$ be a graph, $q=(h_{q},r_{q},?)$ be a link query over $G$, and $\bar{\eta}$ be a walk over $G$. If $\bar{\eta}$ traverses all edges of $G$, then using only the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording function $w$ detailed in Section B.2, we can construct a graph-query pair $(H,q^{\prime})$ isomorphic to $(G,q)$.
Proof.
Suppose that $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}$ visits all edges of $G=(V,E,R)$ and let $l$ be its length. Recall the anonymization functions $\operatorname{id}_{V}(\cdot;\bar{\eta})$ and $\operatorname{id}_{R}(\cdot;\bar{\eta})$ as defined in Section B.2. The output $w(\bar{\eta};G,q,\cdot,\cdot)$ (the embedding functions provided as the last two arguments are irrelevant) is a sequence of tuples $S_{0},S_{1},\dots,S_{l}$ with each $S_{i}$ equal to:
$S_{i}=\left(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{i};\bar{\eta}),\operatorname{dir}_{i},\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}},\cdot,\cdot\right)$
Consider a graph $H=(V^{\prime},E^{\prime},R^{\prime})$ constructed as follows:
  • the vertices $V^{\prime}$ correspond to the anonymized node ids $\operatorname{id}_{V}(v_{i};\bar{\eta})$:
    $V^{\prime}=\{\operatorname{id}_{V}(v;\bar{\eta})\mid v\in V\}$
    Since each vertex must have been visited by $\bar{\eta}$, this is well-defined.
  • the relation types $R^{\prime}$ are the anonymized relation ids $\operatorname{id}_{R}(r_{i};\bar{\eta})$:
    $R^{\prime}=\{\operatorname{id}_{V}(r;\bar{\eta})\mid r\in R\}$
    Again, this is well-defined, as each relation must have been noticed by $\bar{\eta}$.
  • the edges $E^{\prime}$ are reconstructed from the consecutive step encodings using the anonymized vertex and relation indices and the direction $\operatorname{dir}_{i}$:
    $E^{\prime}=$ $\left\{(\operatorname{id}_{V}(v_{i-1}),\operatorname{id}_{R}(r_{i}),\operatorname{id}_{V}(v_{i}))\mid\operatorname{dir}_{i}=0,1\leq i\leq l\right\}$
    $\>\cup\left\{(\operatorname{id}_{V}(v_{i}),\operatorname{id}_{R}(r_{i}),\operatorname{id}_{V}(v_{i-1}))\mid\operatorname{dir}_{i}=1,1\leq i\leq l\right\}$
and a query $q^{\prime}=(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{j},\bar{\eta}),?)$ for $i,j$ such that $\delta_{v_{i}=h_{q}}=1$ and $\delta_{r_{j}=r_{q}}=1$.
Then by the definition of $w$ (Section B.2), it is straightforward to check that the pair $(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$ defines an isomorphism from $(G,q)$ to $(H,q^{\prime})$. Indeed, both these functions are injective by construction, and as $\bar{\eta}$ witnesses all nodes and relations, they are well-defined bijections. For each unique edge traversed by $\bar{\eta}$, there exists a unique edge in $E^{\prime}$ translated to the anonymized space, which implies an isomorphism between $E$ and $E^{\prime}$. Finally, by utilizing the flags $\delta_{v_{i}=h_{q}}$ and $\delta_{r_{j}=r_{q}}$, we can identify the query head node and relation in the new graph. All things considered, we can reconstruct the pair $(G,q)$, up to isomorphism, from the output of $w(\bar{\eta};G,q,\cdot,\cdot)$. ∎
We are now ready to prove the main result regarding the universality of Flock as an approximation of link invariant functions. The outline of the proof is as follows: 1) Using the upper-bound on the edge cover time of graphs in $\mathbb{K}_{n,m}$ derived in Lemma C.4, we can bound the probability of sampling a walk that visits all edges, 2) Once such a walk is sampled, we can recover the graph and query, up to isomorphism, from its anonymized form (Lemma C.5), 3) Lastly, we can return the value of the approximated function for the derived isomorphic instance. Since the approximated function is link invariant, if the reconstructed graph matches the original one, we return precisely the correct value.
Proposition 4.1.
With a powerful enough sequence processor $f_{\theta}$, the Flock framework described in Section 4 is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
Proof.
Let $\varphi:\mathbb{K}_{n,m}\rightarrow(V\times R\times V\rightarrow[0,1])$ be a link invariant function over $\mathbb{K}_{n,m}$ returning values from the interval $[0,1]$. Let $G=(V,E,R)\in\mathbb{K}_{n,m}$, $q=(h,r,?)$ be a link prediction query over $G$ and $t\in V$ be a target node. Pick some $\epsilon,\delta>0$. Our goal is to show that:
${\mathbb{P}}(|\varphi(G)((h,r,t))-X_{\theta}(G,(h,r,?))(t)|<\epsilon)>1-\delta$ (9)
For simplicity, let us consider a situation where only a single walk $\bar{\eta}$ of length $l$ is sampled by the model (otherwise, omit additional walks). We will also restrict the argument to a single refinement case – the result can be extended to multiple refinement steps by returning $\Delta\mathbf{v},\Delta\mathbf{r}=0$ during all additional iterations. Consider a sequence processor $f_{\theta}$ that given the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording protocol, creates a graph-query pair $(H,q^{\prime})$ with $q^{\prime}=(h_{q^{\prime}},r_{q^{\prime}},?)$ using the strategy described in the proof of Lemma C.5, and returns a vector $\mathbf{h}\in\mathbb{R}^{l+1}$ whose $i^{\text{th}}$ entry is equal $\mathbf{h}_{i}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(v_{i};\bar{\eta}))$ where $v_{i}$ is the $i^{\text{th}}$ node visited by $\bar{\eta}$. The consensus protocol $c$, provided $t$ was visited by $\bar{\eta}$, can then identify $t$ as one of the $v_{j}$ and pull the corresponding embedding $\mathbf{h}_{j}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$, returning it as the output $\mathbf{v}(t)=\mathbf{h}_{j}$ (note that no matter which specific value of $j$ is chosen, this value will be the same). Finally, the classification head can work as an identity operation, returning $X_{\theta}(G,q)(t)=\mathbf{v}(t)=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$.
We claim that if the sampled walk $\bar{\eta}$ traverses all edges of $G$, then the output of the Flock model described above satisfies:
$\varphi(G)((h,r,t))=X_{\theta}(G,(h,r,?))(t)$
By Lemma C.5, in such case, the reconstructed pair $(H,q^{\prime})$ is isomorphic to $(G,q)$ by the isomorphism $id=(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$. Since $\varphi$ is link invariant, we can write:
$\varphi(G)((h,r,t))$ $=\varphi(\operatorname{id}(G))((\operatorname{id}_{V}(h;\bar{\eta}),\operatorname{id}_{R}(r;\bar{\eta}),\operatorname{id}_{V}(t;\bar{\eta})))$
$=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta})))$
$=X_{\theta}(G,(h,r,?))(t)$
Therefore, whenever the walk $\bar{\eta}$ witnesses all edges of $G$, the output of the Flock model satisfies:
$\varphi(G)((h,r,t))=X_{\theta}(G,(h,r,?))(t)$
Hence, to show (9), it suffices to prove that we can uniformly choose the length $l$ of the random walk so that the probability of $\bar{\eta}$ covering all the edges is greater than $1-\delta$. By Markov’s inequality:
[…]
$=\frac{C_{E}(G)}{l}$
But by Lemma C.4, $C_{E}(G)\leq C_{n,m}$ for some constant $C_{n,m}$. Hence, taking $l>\frac{C_{n,m}}{\delta}$, we get:
${\mathbb{P}}(\bar{\eta}\text{ does not cover all edges})\leq\frac{C_{E}(G)}{l}\leq\frac{C_{n,m}}{l}<\delta$
This means that for such a choice of $l$:
${\mathbb{P}}(\bar{\eta}\text{ witnesses all edges of }G)>1-\delta$
which leads to the conclusion that for $l>\frac{C_{n,m}}{\delta}$, the proposed Flock framework satisfies:
${\mathbb{P}}(|\varphi(G)((h,r,t))-X_{\theta}(G,(h,r,?))(t)|<\epsilon)>1-\delta$
for any choice of $G=(V,E,R)\in\mathbb{K}_{n,m}$ and $(h,r,t)\in V\times R\times V$. ∎
C.2 Invariance
First, let us recall the definition of invariance for the context of knowledge graphs and the associated notion of invariance in probability. We say that a function $\varphi$ taking KGs as input is invariant if for any pair of isomorphic KGs $G\simeq H$ it produces the same input, i.e. $G\simeq H\implies\varphi(G)=\varphi(H)$.
We extend the notion of invariance for further types of inputs, not limited to full knowledge graphs, particularly to random walks and link prediction queries. Let $G=(V,E,R)\in\mathbb{K}_{n,m}$ and let $H=(V^{\prime},E^{\prime},R^{\prime})\simeq G$ be a KG isomorphic to $G$ via the isomorphism $\mu=(\pi,\phi)$. For any $h\in V$ and $r\in R$, we identify the link prediction query $q=(h,r,?)$ in $H$ using the isomorphism $\mu$ as:
$\mu(q)=\mu((h,r,?))=(\pi(h),\phi(r),?)$
Similarly, let $\eta=v_{0}\xrightarrow{r_{1}}\dots\xrightarrow{r_{l}}v_{l}$ be a walk of length $l$ in $G$. The view of $\eta$ with $\mu$ is defined as:
$\mu\left(v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\right)=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{l})}\pi(v_{l})$
Let $f$ be a function taking inputs drawn from KGs. We call $f$ invariant if for any pair of isomorphic graphs $G\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$\mu$}}}}}}{{\simeq}}H$ and an associated isomorphism $\mu=(\pi,\phi)$, $f$ satisfies
$f(x)=f(\mu(x))$
where $x$ can be, e.g., a walk or link prediction query. In words, invariance means that the function preserves output under the re-identifications of the input graph and the induced transformations of queries and walks.
This notion extends to functions with multiple inputs, where we enforce the transformation on each graph-related input. For example, a function $\varphi$ taking a KG, query and a $d$-dimensional vector is invariant if it satisfies:
$\forall G\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$\mu$}}}}}}{{\simeq}}H,q,\mathbf{v}\in\mathbb{R}^{d}:\quad\varphi(G,q,\mathbf{v})=\varphi(\mu(G),\mu(q),\mathbf{v})$
Following the definition of invariance in probability, provided in Section 3, we extend all the definitions above to the stochastic case, replacing equality (==) with equality in distribution ($\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}$).
We can now prove the main propositions stated in Section 4.2. Let’s begin with the more general:
Proposition 4.2.
Suppose that the walk sampling protocol $\eta$ is invariant in probability and both the recording protocol $w$ and the consensus protocol $c$ are invariant. Then, regardless of the choice of the deterministic sequence processor $f_{\theta}$, the corresponding Flock model is invariant in probability.
Proof.
Let $(V,E,R)=G\simeq H=(V^{\prime},E^{\prime},R^{\prime})$ be isomorphic knowledge graphs with isomorphism $\mu=(\pi,\phi)$ transforming $G$ into $H$. Our goal is to show that when the statement conditions are met for a Flock model $X_{\theta}$ with $I$ refinement steps, then for any link prediction query $q=(h,r,?)$ and any target node $t\in V$, the prediction of Flock for $t$ over $(G,q)$ is an identical random variable to the prediction for $\pi(t)$ over $(H,\mu(q))$, i.e.
$X_{\theta}(G,q)(t)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}X_{\theta}(H,\mu(q))(\pi(t))$
where $\mu(q)=(\pi(h),\phi(r),?)$. Recall that these predictions are defined as:
$X_{\theta}(G,q)(t)$ $\coloneqq{\rm head}({\bf v}^{(I)}(t)+{\bf r}^{(I)}(r))$
$X_{\theta}(H,\mu(q))(\pi(t))$ $\coloneqq{\rm head}({\bf v^{\prime}}^{(I)}(\pi(t))+{\bf r^{\prime}}^{(I)}(\phi(r)))$
As $\rm head$} is a deterministic map, it suffices to show that the final embeddings $\mathbf{v}^{(I)},\mathbf{r}^{(I)}$ for $(G,q)$ and $\mathbf{v^{\prime}}^{(I)},\mathbf{r^{\prime}}^{(I)}$ for $(H,\mu(q))$ satisfy:
$\mathbf{v}^{(I)}(v)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mathbf{v^{\prime}}^{(I)}(\pi(v))\quad\text{and}\quad\mathbf{r}^{(I)}(r)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mathbf{r^{\prime}}^{(I)}(\phi(r))\qquad\forall v\in V,r\in R$
We will prove this result by induction on the number of layers $i$. The base case $i=0$ is trivial, as we initialize the embeddings of all nodes with a pretrained vector $\mathbf{v}_{0}$, and all relations with $\mathbf{r}_{0}$.
For the induction step, suppose the claim holds for $i$. We drop the superscript $(i)$ for readability. The result for $i+1$ becomes apparent by unfolding the definitions of invariance of the considered components. Since $\eta$ is invariant in probability, we have
$\mu(\eta(G))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\eta(H)$ (10)
Let $\eta_{1},\dots,\eta_{n}$ be the random walks over $G$ using $\eta$ and $\eta_{1}^{\prime},\dots,\eta_{n}^{\prime}$ be random walks over $H$. Now, $\eta_{1},\dots,\eta_{n}$ are independent and identically distributed random variables, each following the distribution $\eta_{j}\sim\eta(G)$. Similarly, using (10):
$\eta_{j}^{\prime}\sim\eta(H)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mu(\eta(G))\implies\eta_{j}^{\prime}\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mu(\eta_{j})$ (11)
As the recording protocol $w$ is invariant, $w(\eta_{j})\!\!=\!w(\mu(\eta_{j}))$ for all $j$, which with (11) yields:
$\mathbf{z}_{j}:=w(\eta_{j})=w(\mu(\eta_{j}))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}w(\eta_{j}^{\prime}):=\mathbf{z^{\prime}_{j}}$ (12)
Then, $f_{\theta}$ is a deterministic map, so (12) implies:
$\mathbf{h}_{j}:=f_{\theta}(\mathbf{z}_{j})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}f_{\theta}(\mathbf{z}^{\prime}_{j}):=\mathbf{h}_{j}^{\prime}$
Let $(\Delta\mathbf{v},\Delta\mathbf{r})=c(\mathbf{h}_{1:N},\eta_{1:N})$, $(\Delta\mathbf{v}^{\prime},\Delta\mathbf{r}^{\prime})=c(\mathbf{h}^{\prime}_{1:N},\eta^{\prime}_{1:N})$ be the outputs of the consensus protocol. We will denote by $c_{\mathbf{v}}$ and $c_{\mathbf{r}}$, the restrictions to the first and second output, e.g. $\Delta\mathbf{v}=c_{\mathbf{v}}(\mathbf{h}_{1:N},\eta_{1:N})$. Let $\mathbf{x}\in\mathbb{R}^{d}$ be a vector, and denote by $\mathcal{W}(G)$ the space of walks over $G$. For any vertex $v\in V$, the probability that $\Delta\mathbf{v}(v)=\mathbf{x}$ equals:
${\mathbb{P}}(\Delta\mathbf{v}(v)=\mathbf{x})$ $=\sum_{\bar{\eta}\in\mathcal{W}(G)^{n}}{\mathbb{P}}(\Delta\mathbf{v}(v)=\mathbf{x}|\eta_{1:N}=\bar{\eta})\cdot{\mathbb{P}}(\eta_{1:N}=\bar{\eta})$
$=\sum_{\bar{\eta}\in\mathcal{W}(G)^{n}}{\mathbb{P}}(c_{\mathbf{v}}(\mathbf{h}_{1:N},\eta_{1:N})=\mathbf{x}|\eta_{1:N}=\bar{\eta})\cdot{\mathbb{P}}(\eta_{1:N}=\bar{\eta})$
$=\sum_{\bar{\eta}\in\mathcal{W}(G)^{n}}{\mathbb{P}}(c_{\mathbf{v}}(f_{\theta}(w(\eta_{1:N})),\eta_{1:N})=\mathbf{x}|\eta_{1:N}=\bar{\eta})\cdot{\mathbb{P}}(\eta_{1:N}=\bar{\eta})$
$=\sum_{\bar{\eta}\in\mathcal{W}(G)^{n}}{\mathbb{P}}(c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\bar{\eta})=\mathbf{x}|\eta_{1:N}=\bar{\eta})\cdot{\mathbb{P}}(\eta_{1:N}=\bar{\eta})$
$=\sum_{\begin{subarray}{c}\bar{\eta}\in\mathcal{W}(G)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\bar{\eta})(v)=\mathbf{x}\end{subarray}}{\mathbb{P}}(\eta_{1:N}=\bar{\eta})$
Similarly, we can derive:
${\mathbb{P}}(\Delta\mathbf{v}^{\prime}(\pi(v))=\mathbf{x})=\sum_{\begin{subarray}{c}\bar{\eta}^{\prime}\in\mathcal{W}(H)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta}^{\prime})),\bar{\eta}^{\prime})(\pi(v))=\mathbf{x}\end{subarray}}{\mathbb{P}}(\eta^{\prime}_{1:N}=\bar{\eta}^{\prime})$
Using the invariance of the consensus protocol and the invariance of $f_{\theta}\circ w$, we can write:
$c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta}^{\prime})),\bar{\eta}^{\prime})(\pi(v))$ $=c_{\mathbf{v}}(f_{\theta}(w(\mu(\bar{\eta}))),\mu(\bar{\eta}))(\pi(v))$
$=c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\mu(\bar{\eta}))(\pi(v))$
$=c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\bar{\eta})(v)$
The graph isomorphism $\mu$ defines a bijection between walks $\mathcal{W}(G)$ in $G$ and walks $\mathcal{W}(H)$ in $H$, so we can use this correspondence to deduce:
${\mathbb{P}}(\Delta\mathbf{v}^{\prime}(\pi(v))=\mathbf{x})\;\;\;=$ $\sum_{\begin{subarray}{c}\bar{\eta}^{\prime}\in\mathcal{W}(H)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta}^{\prime})),\bar{\eta}^{\prime})(\pi(v))=\mathbf{x}\end{subarray}}\!\!\!\!\!{\mathbb{P}}(\eta^{\prime}_{1:N}=\bar{\eta}^{\prime})$ (13)
$\;\;\;=$ $\sum_{\begin{subarray}{c}\mu(\bar{\eta})\in\mathcal{W}(H)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\mu(\bar{\eta}))),\mu(\bar{\eta}))(\pi(v))=\mathbf{x}\end{subarray}}\!\!\!\!\!{\mathbb{P}}(\eta^{\prime}_{1:N}=\mu(\bar{\eta}))$
$\;\;\;=$ $\sum_{\begin{subarray}{c}\bar{\eta}\in\mathcal{W}(G)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\bar{\eta})(v)=\mathbf{x}\end{subarray}}\!\!\!\!\!{\mathbb{P}}(\eta^{\prime}_{1:N}=\mu(\bar{\eta}))$
Since $\eta$ is invariant in probability, ${\mathbb{P}}(\eta_{1:N}\!=\!\bar{\eta})={\mathbb{P}}(\eta^{\prime}_{1:N}\!=\!\mu(\bar{\eta}))$. Applying this to (13) yields:
${\mathbb{P}}(\Delta\mathbf{v}^{\prime}(\pi(v))=\mathbf{x})$ $=\sum_{\begin{subarray}{c}\bar{\eta}\in\mathcal{W}(G)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\bar{\eta})(v)=\mathbf{x}\end{subarray}}\!\!\!\!\!{\mathbb{P}}(\eta^{\prime}_{1:N}=\mu(\bar{\eta}))$
$=\!\!\!\!\sum_{\begin{subarray}{c}\bar{\eta}\in\mathcal{W}(G)^{n}\\ c_{\mathbf{v}}(f_{\theta}(w(\bar{\eta})),\bar{\eta})(v)=\mathbf{x}\end{subarray}}\!\!\!\!\!\!\!{\mathbb{P}}(\eta_{1:N}=\bar{\eta})={\mathbb{P}}(\Delta\mathbf{v}(v)=\mathbf{x})$
As $\mathbf{x}$ was chosen arbitrarily, we can conclude that $\Delta\mathbf{v}(v)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\Delta\mathbf{v}^{\prime}(\pi(v))$. The proof for relations follows analogously, considering $c_{\mathbf{r}}$ instead of $c_{\mathbf{v}}$. This allows us to write:
$\Delta\mathbf{v}(v)$ $\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\Delta\mathbf{v}^{\prime}(\pi(v))$ $\quad\forall v\in V$ (14)
$\Delta\mathbf{r}(r)$ $\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\Delta\mathbf{r}^{\prime}(\phi(r))$ $\quad\forall r\in R$
By the induction hypothesis, $\mathbf{v}^{(i)}(v)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mathbf{v^{\prime}}^{(i)}(\pi(v))$ for all $v\in V$ and $\mathbf{r}^{(i)}(r)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mathbf{r^{\prime}}^{(i)}(r)$ for all $r\in R$. Therefore, by (14), combined with properties of sums of random variables:
$\mathbf{v}^{(i+1)}(v)=\mathbf{v}^{(i)}(v)+\Delta\mathbf{v}(v)$ $\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mathbf{v^{\prime}}^{(i)}(\pi(v))+\Delta\mathbf{v}^{\prime}(\pi(v))=\mathbf{v^{\prime}}^{(i+1)}(\pi(v))$ $\quad\forall v\in V$
$\mathbf{r}^{(i+1)}(r)=\mathbf{r}^{(i)}(r)+\Delta\mathbf{r}(r)$ $\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\mathbf{r^{\prime}}^{(i)}(\phi(r))+\Delta\mathbf{r}^{\prime}(\phi(r))=\mathbf{r^{\prime}}^{(i+1)}(\phi(r))$ $\quad\forall r\in R$
which completes the induction step, and hence the proof. ∎
We can use the conclusion from Proposition 4.2 to prove the probabilistic invariance of the architecture proposed in Section 4. To be able to apply it, we first need to verify the invariance of all used components, which we formalize in the following lemmas.
Lemma C.6.
The choice of the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ of the uniform random walk algorithm described in Section B.1 is invariant.
Proof.
Let $G=(V,E,R)$ be a graph and let $H\simeq G$ be an isomorphic graph, with the isomorphism $\mu=(\pi,\phi)$ taking $G$ to $H$. Consider a link prediction query $q=(h,r,?)$ over $G$, and its identification $q^{\prime}=\mu(q)=(\pi(h),\phi(r),?)$. The goal is to show that when using $\eta$ described in Section B.1 for $(G,q)$ and $(H,q^{\prime})$, the first steps:
$V_{0}\xrightarrow{R_{1}}V_{1}\qquad\text{and}\qquad U_{0}\xrightarrow{S_{1}}U_{1}$
of the execution of $\eta$ over $G$ and $H$, respectively, satisfy the following property:
$\pi(V_{0})\xrightarrow{\phi(R_{1})}\pi(V_{1})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}U_{0}\xrightarrow{S_{1}}U_{1}$
By definition of $\eta$, there are three scenarios of choosing the first step, each with probability $\frac{1}{3}$. Hence, it suffices to show that within each scenario, the selection process is invariant in probability:
  • Scenario 1: selecting the query head as the first node, then proceeding by random. First, $\pi$ takes the head node of $q$ to the head node of $q^{\prime}$. Secondly, as isomorphisms preserve the number of neighboring nodes and number of edges between a pair of nodes, we have:
[…]
    and
[…]
    $={\mathbb{P}}(S_{1}=\phi(r)\mid U_{1}=\pi(w))$
  • Scenario 2: selecting an edge with query relation type at random. Here, we use the fact that isomorphisms preserve the number of edges of a given type. Hence, $\mu$ defines a bijection between the sets of edges with type $r$ in $G$ and type $\phi(r)$ in $H$, which allows us to conclude that this scenario is also invariant in probability.
  • Scenario 3: selecting the first step completely at random. This case is similar to Scenario 1 – using the invariance of the number of neighboring nodes under isomorphism, we can repeat similar calculations in a straightforward manner to show probabilistic invariance.
Either way, we find that the selection process of the first step of $\eta$ over $G$ translates naturally via $\mu$ to the choice of the first step over $H$, proving the desired statement. ∎
Lemma C.7.
Suppose that the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ is chosen in an invariant manner. Then, the uniform random walk with no backtracking algorithm $\eta$ is invariant in probability.
Proof.
Let $G=(V,E,R)$ be a knowledge graph, and let $l$ be the length of random walks. Let $H$ be a KG isomorphic to $G$ via the isomorphism $\mu=(\pi,\phi)$. We aim to show that:
$\mu(\eta(G,l))=\pi(V_{0})\xrightarrow{\phi(R_{1})}\pi(V_{1})\xrightarrow{\phi(R_{2})}\dots\xrightarrow{\phi(R_{l})}\pi(V_{l})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}U_{0}\xrightarrow{S_{1}}U_{1}\xrightarrow{S_{2}}\dots\xrightarrow{S_{l}}U_{l}=\eta(H,l)$
Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\in\mathcal{W}(G)$ be a walk of length $l$ over $G$. It suffices to show that the probability of sampling $\bar{\eta}$ from $G$ is identical to the probability of sampling $\mu(\bar{\eta})$ from $H$:
${\mathbb{P}}(\eta(G,l)=\bar{\eta})={\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))$
To see this, let us expand the definitions of ${\mathbb{P}}(\eta(G,l)=\bar{\eta})$:
${\mathbb{P}}(\eta(G,l)=\bar{\eta})=$ ${\mathbb{P}}(V_{0}=v_{0})$ (15)
$\cdot{\mathbb{P}}(V_{1}=v_{1}\mid V_{0}=v_{0})$
$\cdot\prod^{l-2}_{i=0}{\mathbb{P}}(V_{i+2}=v_{i+2}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
$\cdot\prod^{l-1}_{i=0}{\mathbb{P}}(R_{i+1}=r_{i+1}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
and $P(\eta(H,l)=\mu(\bar{\eta}))$:
${\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))=$ ${\mathbb{P}}(U_{0}=\pi(v_{0}))$ (16)
$\cdot{\mathbb{P}}(U_{1}=\pi(v_{1})\mid U_{0}=\pi(v_{0}))$
$\cdot\prod^{l-2}_{i=0}{\mathbb{P}}(U_{i+2}=\pi(v_{i+2})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
$\cdot\prod^{l-1}_{i=0}{\mathbb{P}}(S_{i+1}=\phi(r_{i+1})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
Given that the graph isomorphism preserves the number of neighbors for each node and is a bijection, we can easily verify using the definitions from (8) that the following indeed hold:
${\mathbb{P}}(V_{i+2}=v_{i+2}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$ $={\mathbb{P}}(U_{i+2}=\pi(v_{i+2})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$ (17)
${\mathbb{P}}(R_{j+1}=r_{j+1}\mid V_{j+1}=v_{j+1},V_{j}=v_{j})$ $={\mathbb{P}}(S_{j+1}=\phi(r_{j+1})\mid U_{j+1}=\pi(v_{j+1}),U_{j}=\pi(v_{j}))$
for all $i\in\{0,1,\dots,l-2\},j\in\{1,\dots,l-1\}$. Moreover, by the assumption that the first step $V_{0}\xrightarrow{R_{1}}V_{1}$ is invariant, we have:
${\mathbb{P}}((V_{0},R_{1},V_{1})=(v_{0},r_{1},v_{1}))={\mathbb{P}}((U_{0},S_{1},U_{1})=(\pi(v_{0}),\phi(r_{1}),\pi(v_{1})))$ (18)
But by the laws of conditional probability:
${\mathbb{P}}((V_{0},R_{1},V_{1})=(v_{0},r_{1},v_{1}))$ $={\mathbb{P}}(R_{1}=r_{1}\mid V_{0}=v_{0},V_{1}=v_{1})\cdot{\mathbb{P}}(V_{0}=v_{0},V_{1}=v_{1})$
$={\mathbb{P}}(R_{1}=r_{1}\mid V_{0}=v_{0},V_{1}=v_{1})\cdot{\mathbb{P}}(V_{1}=v_{0}\mid V_{0}=v_{0})\cdot{\mathbb{P}}(V_{0}=v_{0})$
and analogously:
${\mathbb{P}}((U_{0},$ $S_{1},U_{1})=(\pi(v_{0}),\phi(r_{1}),\pi(v_{1})))$
$={\mathbb{P}}(S_{1}=\phi(r_{1})\mid U_{0}=\pi(v_{0}),U_{1}=\pi(v_{1}))\cdot{\mathbb{P}}(U_{1}=\pi(v_{0})\mid U_{0}=\pi(v_{0}))\cdot{\mathbb{P}}(U_{0}=\pi(v_{0}))$
Substituting these equalities into (18) and multiplying both sides by the equalities from (17) for all choices of $i\in\{0,1,\dots,l-2\},j\in\{1,\dots,l-1\}$, we get precisely the equality of the right sides of equations (15) and (16). Hence,
${\mathbb{P}}(\eta(G,l)=\bar{\eta})={\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))$
and we can conclude that $\mu(\eta(G,l))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\eta(H,l)$, and the algorithm $\eta$ is invariant in probability. ∎
Corollary C.8.
The random walk algorithm presented in Section B.1 is invariant in probability.
Lemma C.9.
The recording protocol $w$, as described in Section B.2, is invariant, provided that the embedding functions $\mathbf{v}$ and $\mathbf{r}$ are invariant.
Proof.
Let $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$ be isomorphic knowledge graphs with the isomorphism $\mu=(\pi,\phi)$ taking $G$ to $H$. Let $q=(h_{q},r_{q},?)$ be a link prediction query over $G$, and $\mu(q)=(\pi(h_{q}),\phi(r_{q}),?)$ be its identification in $H$. Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\in\mathcal{W}(G)$ be a walk over $G$, and $\bar{\eta}^{\prime}=\mu(\bar{\eta})=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{l})}\pi(v_{l})$ be the analogous walk over $H$. To prove that the recording protocol $w$ outlined in Section B.2 is invariant, it suffices to show that the encoding of each step:
$S_{i}=\left(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{i};\bar{\eta}),\operatorname{dir}_{i},\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}},\mathbf{v}(v_{i}),\mathbf{r}(r_{i})\right)$
is identical for $\bar{\eta}$ and $\bar{\eta}^{\prime}$. We will show this for each component:
  • since $\pi$ defines a bijection between nodes in $G$ and $H$, for any $i$, we have:
    $\operatorname{id}_{V}(v_{i};\bar{\eta})=\operatorname*{arg\,min}_{t}[v_{t}=v_{i}]=\operatorname*{arg\,min}_{t}[\pi(v_{t})=\pi(v_{i})]=\operatorname{id}_{V}(\pi(v_{i});\bar{\eta}^{\prime})$
  • similarly to the point above, $\phi$ is a bijection between relations of $G$ and $H$, so we can write:
    $\operatorname{id}_{R}(r_{i};\bar{\eta})$ $=\operatorname*{arg\,min}_{t}\left[r_{t}=r_{i}\lor r_{t}=r_{i}^{-1}\right]$
    $=\operatorname*{arg\,min}_{t}\left[\phi(r_{t})=\phi(r_{i})\lor\phi(r_{t})=\phi(r_{i})^{-1}\right]$
    $=\operatorname{id}_{R}(\phi(r_{i});\bar{\eta}^{\prime})$
  • $\operatorname{dir}_{i}$ is clearly preserved, as the isomorphism $\mu$ preserves the directions of edges,
  • as $\pi,\phi$ are bijections the masks $\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}}$, representing whether the $i$’th node and relation match the types in the query, satisfy:
    $v_{i}=h_{q}\iff\pi(v_{i})=\pi(h_{q})$ $\implies\quad\delta_{v_{i}=h_{q}}=\delta_{\pi(v_{i})=\pi(h_{q})}$
    $r_{i}=r_{q}\iff\phi(r_{i})=\phi(r_{q})$ $\implies\quad\delta_{r_{i}=r_{q}}=\delta_{\phi(r_{i})=\phi(r_{q})}$
  • $\mathbf{v}$ and $\mathbf{r}$ are invariant by assumption, so:
    $\mathbf{v}(v_{i})=\mathbf{v}(\pi(v_{i}))\qquad\text{and}\qquad\mathbf{r}(r_{i})=\mathbf{r}(\phi(r_{i}))$
Combining all these observations, we can conclude that $w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=w(\mu(\bar{\eta});H,\mu(q),\mathbf{v},\mathbf{r})$ and $w$ is indeed invariant. ∎
Lemma C.10.
The consensus protocol $c$, as described in Section B.4, is invariant.
Proof.
Let $G=(V,E,R)$ be a knowledge graph and $H$ be isomorphic to $G$ via an isomorphism $\mu=(\pi,\phi)$. Let $\bar{\eta}_{1:N}\in\mathcal{W}(G)$ be a sequence of walks in $G$. To show that the output of the consensus protocol is invariant, we need to prove that for each $v\in V$ and $r\in R$, the following holds:
$\Delta\mathbf{v}(v)=\mathbf{v}^{\prime}(\pi(v))\qquad\text{and}\qquad\Delta\mathbf{r}(r)=\Delta\mathbf{r}^{\prime}(\phi(r))$ (19)
where $(\Delta\mathbf{v},\Delta\mathbf{r})=c(\mathbf{h},\bar{\eta}_{1:N})$ and $(\Delta\mathbf{v}^{\prime},\Delta\mathbf{r}^{\prime})=c(\mathbf{h},\mu(\bar{\eta}_{1:N}))$ for $\mathbf{h}=(\mathbf{s}^{(V)},\mathbf{s}^{(R)},\mathbf{a}^{(V)},\mathbf{a}^{(R)})$.
The result follows from the fact that $\pi$ and $\phi$ are bijections – whenever $v$ is the $j^{\text{th}}$ vertex visited in the walk $\bar{\eta}_{i}$, the $j^{\text{th}}$ node of $\mu(\bar{\eta}_{i})$ must be $\pi(v)$ (and vice versa). An analogous result holds for the relations. Hence, the aggregation performed by $c$ for $v$ (resp. $r$) over $\bar{\eta}_{1:N}$ is equivalent to the aggregation for $\pi(v)$ (resp. $\phi(r)$) over $\mu(\bar{\eta}_{1:N})$, and (19) is indeed satisfied. ∎
Proposition 4.3.
Flock with components as described in Section 4 is invariant in probability.
Proof.
The result follows naturally from aggregating the results of Corollary C.8 and Lemmas C.9 and C.10, followed by applying Proposition 4.2. ∎
Appendix D Computational Complexity
Recall that $I$ is the iterations in each forward pass of Flock; $n$ is the base walk count; $\ell$ is the walk length; $L$ is the number of linear sequence-model layers (such as GRU); and $d$ is the hidden dimension for the sequence processor. Note that in practice, we perform $P$ forward passes and ensemble their outputs to reduce variance. For a single pass (P=1P{=}1), walk sampling and recording cost $O(n\ell)$, while the sequence processor with $L$ layers of hidden dimension $d$ costs $O(n\ell Ld^{2})$. The consensus protocol costs $O(n\ell d)$. In total, the time complexity is $O\big(PIn\ell Ld^{2}\big)$, which scales linearly with the number of (base) walks $n$, the length of walks $\ell$, and the number of ensembled predictions $P$. We empirically verified this in Appendix E.
Compared with message-passing KGFMs like Ultra and Trix, Flock’s complexity is independent of the graph size and average degree; empirically, however, using more walks (increasing $n$) and longer walks (increasing $\ell$) improves coverage and yields more fine-grained representation.
The space complexity of Flock per forward pass is $O(n\ell d)$ plus model parameters $O(Ld^{2})$. Note that running ensembles sequentially keeps peak memory near this bound, while parallel ensembling increases it by a factor of $P$.
Appendix E Scalability analysis
Table 5: Training scalability analysis on a single NVIDIA RTX A6000 (48 GB) with batch size = 8. Flock using $16$ number of base walks and 11 ensemble.
Model Parameters Time / batch (s) GPU memory (GB)
Ultra 168,705 0.117 2.110
Trix 87,138 0.690 3.442
Flock 801,969 1.30 27.89
Table 6: Inference scalability on a single NVIDIA RTX A6000 (48 GB); batch size =8=8. Left columns specify base walks $n$ and ensembled passes $P$. Dashes indicate not applicable.
Model # Base Walks $n$ Ensemble $P$ Time /batch (s) GPU memory (GB)
Ultra 1 0.073 0.848
Trix 1 0.500 1.382
Flock 16 1 1.26 2.868
16 2 1.99 2.864
16 4 3.24 3.938
16 8 5.45 5.172
16 16 9.45 8.892
128 1 1.77 5.000
128 2 2.80 7.880
128 4 5.00 14.42
128 8 10.05 43.68
To investigate the scalability of the proposed method Flock, we report the training and inference time per batch and peak GPU memory for Ultra, Trix, and variants of Flock on a single RTX A6000 (48 GB) in Tables 5 and 6.
Training.
During training, we fix Flock to $n=16$ base walks and with an ensemble size of $P=1$, which yields higher cost than Ultra/Trix but remains feasible on a single GPU. In addition, unlike Ultra/Trix, Flock does not rely on GNN message passing where highly optimized fused sparse kernels (e.g., RSPMM kernel developed in Zhu et al. (2021)) accelerate computation; instead, Flock’s runtime is dominated by walk sampling and sequence encoding, making time per batch the main bottleneck. As a result, pretraining typically takes about three days. One avenue for future work is to develop similarly highly optimized kernels for random-walk sampling to speed up the process.
Inference.
Additionally, we report the inference results in Table 6, where we vary the number of walks $n$ and ensembled passes $P$. We observe near-linear growth of latency and VRAM with $n$, reflecting the dominant costs of walk sampling and sequence processing. Note that during inference, ensembled predictions are parallelizable, meaning that with sufficient GPU memory, these $P$ stochastic passes can be executed concurrently, so the latency grows sublinearly in $P$, while peak VRAM scales roughly linearly with $P$. In practice, reducing $n$ (walks) or $P$ (ensembled passes) lowers both memory and latency, while larger $n/P$ settings trade extra cost for better coverage and stability on harder KGs.
Appendix F Noise injection over existing KGFMs
Setup.
To answer Q5, we apply noise injection over the best performing KGFMs baselines Trix. Specifically, in each forward pass, we subtract an element-wise noise sampled from a uniform distribution $\epsilon\sim{\mathcal{U}}[0,0.5]$ to all relation and entity embeddings after the initialization stage. Note that the addition of noise will technically break deterministic node-relation equivariance, but the resulting model (Trix + noise) still respects probabilistic node-relation equivariance. We then pretrain Trix using the same experimental setup shown in Section 5.2, and compare with Trix without noise injection and Flock. To minimize the variance induced by injected noise and to ensure a fair comparison, we report ensembled prediction results with 16 samples for both Trix + noise and Flock.
(a) Zero-shot performance for entity prediction.
MRR Hits@10
Trix 0.366 0.518
+ noise 0.385 0.545
Flock 0.391 0.560
(b) Zero-shot performance for relation prediction.
MRR Hits@1
Trix 0.792 0.687
+ noise 0.739 0.643
Flock 0.881 0.817
Results.
We report the average zero-shot performance for entity prediction and relation prediction over 54 KGs in Tables 7(a) and 7(b), respectively. Across all tasks, Trix with naive noise injection fails to close the gap between Flock. In particular, Trix + noise degrades compared with vanilla Trix without noise injection in relation prediction, while boosting the performance in the entity prediction task. We hypothesize that such a difference lies in the added randomization breaks symmetry among entity embeddings more than among relation embeddings, and entity prediction depends more on having distinguishable entity representations than relation prediction does. Additionally, we attribute this performance gap between Flock and Trix + noise to the source of randomization. Flock introduces stochasticity through random walks, which induces structure-informed perturbations that respect the underlying topology. In contrast, Trix with naive noise injection attempts to break deterministic node-relation equivariance by introducing structure-agonistic noise, which might, in turn, hurt the model’s generalization. Together, these findings suggest that simply adding structure-agonistic noise is insufficient; performance gains only arise when stochasticity is topology-aware and is induced from the graph structure in a principled way.
Appendix G Further experimental details
Table 8: Zero-shot inductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Inductive $e,r$
FB-25 0.388 0.640 0.393 0.650 0.404 0.664
FB-50 0.338 0.543 0.334 0.547 0.352 0.566
FB-75 0.403 0.604 0.401 0.611 0.418 0.622
FB-100 0.449 0.642 0.436 0.635 0.452 0.663
WK-25 0.316 0.532 0.305 0.496 0.280 0.491
WK-50 0.166 0.324 0.166 0.313 0.136 0.278
WK-75 0.365 0.537 0.368 0.513 0.382 0.538
WK-100 0.164 0.286 0.188 0.299 0.187 0.304
NL-0 0.342 0.523 0.385 0.549 0.381 0.606
NL-25 0.395 0.569 0.377 0.589 0.345 0.590
NL-50 0.407 0.570 0.404 0.548 0.366 0.565
NL-75 0.368 0.547 0.351 0.525 0.311 0.524
NL-100 0.471 0.651 0.486 0.676 0.452 0.692
MT1 tax 0.224 0.305 0.358 0.452 0.282 0.383
MT1 health 0.298 0.374 0.376 0.457 0.385 0.481
MT2 org 0.095 0.159 0.091 0.156 0.100 0.163
MT2 sci 0.258 0.354 0.323 0.465 0.318 0.458
MT3 art 0.259 0.402 0.284 0.441 0.301 0.466
MT3 infra 0.619 0.755 0.655 0.797 0.684 0.821
MT4 sci 0.274 0.449 0.290 0.460 0.301 0.463
MT4 health 0.624 0.737 0.677 0.775 0.680 0.780
Metafam 0.238 0.644 0.341 0.815 0.476 0.935
FBNELL 0.485 0.652 0.473 0.660 0.502 0.700
Inductive $e$
FB-v1 0.498 0.656 0.515 0.682 0.500 0.697
FB-v2 0.512 0.700 0.525 0.730 0.535 0.737
FB-v3 0.491 0.654 0.501 0.669 0.511 0.685
FB-v4 0.486 0.677 0.493 0.687 0.505 0.702
WN-v1 0.648 0.768 0.699 0.791 0.698 0.803
WN-v2 0.663 0.765 0.678 0.781 0.696 0.790
WN-v3 0.376 0.476 0.418 0.541 0.467 0.608
WN-v4 0.611 0.705 0.648 0.723 0.653 0.729
NL-v1 0.785 0.913 0.806 0.898 0.658 0.863
NL-v2 0.526 0.707 0.569 0.768 0.588 0.797
NL-v3 0.515 0.702 0.558 0.743 0.590 0.783
NL-v4 0.479 0.712 0.538 0.765 0.555 0.786
HM 1k 0.059 0.092 0.072 0.128 0.069 0.119
HM 3k 0.037 0.077 0.069 0.119 0.067 0.118
HM 5k 0.034 0.071 0.062 0.110 0.064 0.116
HM Indigo 0.440 0.648 0.436 0.645 0.423 0.638
ILPC Small 0.302 0.443 0.303 0.455 0.309 0.459
ILPC Large 0.290 0.424 0.307 0.428 0.318 0.438
Table 9: Zero-shot transductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Transductive
NELL995 0.406 0.543 0.472 0.629 0.494 0.655
NELL23k 0.239 0.408 0.290 0.497 0.233 0.398
WDsinger 0.382 0.498 0.511 0.609 0.410 0.528
ConceptNet100k 0.082 0.162 0.193 0.345 0.248 0.453
CoDEx Small 0.472 0.667 0.472 0.670 0.441 0.644
CoDEx Large 0.338 0.469 0.335 0.469 0.342 0.464
YAGO310 0.451 0.615 0.409 0.627 0.414 0.674
AristoV4 0.182 0.282 0.181 0.286 0.308 0.443
DBpedia100k 0.398 0.576 0.426 0.603 0.450 0.627
Hetionet 0.257 0.379 0.279 0.420 0.246 0.371
FB15k-237(10) 0.248 0.398 0.246 0.393 0.246 0.402
FB15k-237(20) 0.272 0.436 0.269 0.430 0.273 0.444
FB15k-237(50) 0.324 0.526 0.321 0.521 0.319 0.518
Table 10: Finetuned inductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Inductive $e,r$
FB-25 0.383 0.635 0.393 0.650 0.405 0.666
FB-50 0.334 0.538 0.334 0.547 0.357 0.570
FB-75 0.400 0.598 0.401 0.611 0.425 0.630
FB-100 0.444 0.643 0.436 0.633 0.460 0.668
WK-25 0.321 0.535 0.300 0.493 0.298 0.506
WK-50 0.140 0.280 0.166 0.313 0.127 0.260
WK-75 0.380 0.530 0.368 0.513 0.405 0.556
WK-100 0.168 0.286 0.188 0.299 0.187 0.306
NL-0 0.329 0.551 0.385 0.549 0.418 0.619
NL-25 0.407 0.596 0.377 0.589 0.405 0.626
NL-50 0.418 0.595 0.405 0.555 0.391 0.562
NL-75 0.374 0.570 0.351 0.525 0.344 0.544
NL-100 0.458 0.684 0.482 0.691 0.486 0.714
MT1 tax 0.330 0.459 0.397 0.508 0.413 0.497
MT1 health 0.380 0.467 0.376 0.457 0.394 0.493
MT2 org 0.104 0.170 0.098 0.162 0.107 0.174
MT2 sci 0.311 0.451 0.331 0.526 0.366 0.525
MT3 art 0.306 0.473 0.289 0.441 0.330 0.483
MT3 infra 0.657 0.807 0.672 0.810 0.709 0.838
MT4 sci 0.303 0.478 0.305 0.482 0.324 0.509
MT4 health 0.704 0.785 0.702 0.785 0.711 0.790
Metafam 0.997 1.000 0.997 1.000 0.992 1.000
FBNELL 0.481 0.661 0.478 0.655 0.531 0.714
Inductive $e$
FB-v1 0.509 0.670 0.515 0.682 0.549 0.721
FB-v2 0.524 0.710 0.525 0.730 0.553 0.754
FB-v3 0.504 0.663 0.501 0.669 0.528 0.696
FB-v4 0.496 0.684 0.493 0.687 0.510 0.702
WN-v1 0.685 0.793 0.705 0.798 0.715 0.811
WN-v2 0.679 0.779 0.682 0.780 0.702 0.795
WN-v3 0.411 0.546 0.425 0.543 0.494 0.627
WN-v4 0.614 0.720 0.650 0.722 0.665 0.741
NL-v1 0.757 0.878 0.804 0.899 0.762 0.928
NL-v2 0.575 0.761 0.571 0.764 0.612 0.806
NL-v3 0.563 0.755 0.571 0.759 0.606 0.803
NL-v4 0.469 0.733 0.551 0.772 0.572 0.801
HM 1k 0.042 0.100 0.072 0.128 0.071 0.153
HM 3k 0.030 0.090 0.069 0.119 0.067 0.153
HM 5k 0.025 0.068 0.074 0.118 0.061 0.130
HM Indigo 0.432 0.639 0.436 0.645 0.418 0.633
ILPC Small 0.303 0.453 0.303 0.455 0.305 0.454
ILPC Large 0.308 0.431 0.310 0.431 0.320 0.441
Table 11: Finetune transductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Pretrained
FB15k-237 0.368 0.564 0.366 0.559 0.343 0.532
WN18RR 0.480 0.614 0.514 0.611 0.550 0.656
CoDEx Medium 0.372 0.525 0.365 0.521 0.351 0.496
Transductive
NELL995 0.509 0.660 0.506 0.648 0.531 0.665
NELL23k 0.268 0.450 0.306 0.536 0.280 0.465
WDsinger 0.417 0.526 0.502 0.620 0.435 0.543
ConceptNet100k 0.310 0.529 0.340 0.564 0.352 0.580
CoDEx Small 0.490 0.686 0.484 0.676 0.463 0.648
CoDEx Large 0.343 0.478 0.348 0.481 0.342 0.467
YAGO310 0.557 0.710 0.541 0.702 0.552 0.700
AristoV4 0.343 0.496 0.345 0.499 0.383 0.523
DBpedia100k 0.436 0.603 0.457 0.619 0.470 0.623
Hetionet 0.399 0.538 0.394 0.534 0.314 0.465
FB15k-237(10) 0.254 0.411 0.253 0.408 0.260 0.420
FB15k-237(20) 0.274 0.445 0.273 0.441 0.284 0.459
FB15k-237(50) 0.325 0.528 0.322 0.522 0.317 0.517
Table 12: Zero-shot inductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Inductive $e,r$
FB-25 0.687 0.565 0.805 0.724 0.895 0.839
FB-50 0.696 0.575 0.780 0.699 0.880 0.820
FB-75 0.698 0.555 0.822 0.747 0.903 0.844
FB-100 0.830 0.728 0.921 0.880 0.962 0.938
WK-25 0.857 0.760 0.881 0.823 0.952 0.929
WK-50 0.865 0.793 0.868 0.818 0.921 0.882
WK-75 0.911 0.875 0.916 0.883 0.962 0.944
WK-100 0.887 0.812 0.907 0.869 0.963 0.937
NL-0 0.632 0.502 0.658 0.519 0.714 0.574
NL-25 0.688 0.562 0.742 0.614 0.729 0.632
NL-50 0.680 0.569 0.755 0.636 0.813 0.728
NL-75 0.795 0.692 0.788 0.699 0.833 0.756
NL-100 0.743 0.564 0.884 0.796 0.939 0.889
MT1 tax 0.985 0.976 0.975 0.958 0.998 0.997
MT1 health 0.721 0.561 0.973 0.949 0.991 0.983
MT2 org 0.974 0.951 0.986 0.973 0.991 0.984
MT2 sci 0.976 0.961 0.964 0.941 0.995 0.992
MT3 art 0.881 0.798 0.885 0.825 0.944 0.907
MT3 infra 0.962 0.935 0.940 0.905 0.989 0.980
MT4 sci 0.933 0.891 0.966 0.944 0.974 0.957
MT4 health 0.826 0.719 0.937 0.898 0.990 0.983
Metafam 0.124 0.000 0.291 0.011 0.490 0.223
FBNELL 0.700 0.564 0.726 0.605 0.833 0.737
Inductive $e$
FB-v1 0.646 0.523 0.705 0.599 0.814 0.723
FB-v2 0.695 0.570 0.713 0.590 0.847 0.761
FB-v3 0.679 0.553 0.742 0.644 0.860 0.780
FB-v4 0.638 0.488 0.766 0.665 0.873 0.799
WN-v1 0.836 0.740 0.792 0.613 0.924 0.858
WN-v2 0.853 0.790 0.764 0.572 0.924 0.863
WN-v3 0.707 0.577 0.741 0.568 0.937 0.888
WN-v4 0.860 0.803 0.764 0.570 0.937 0.886
NL-v1 0.636 0.358 0.657 0.453 0.862 0.731
NL-v2 0.742 0.652 0.780 0.696 0.893 0.855
NL-v3 0.669 0.544 0.725 0.612 0.815 0.731
NL-v4 0.606 0.489 0.794 0.691 0.868 0.807
ILPC Small 0.905 0.843 0.919 0.872 0.955 0.921
ILPC Large 0.875 0.799 0.894 0.829 0.948 0.908
HM 1k 0.626 0.447 0.663 0.414 0.687 0.500
HM 3k 0.592 0.439 0.664 0.418 0.714 0.549
HM 5k 0.605 0.452 0.672 0.428 0.746 0.593
HM Indigo 0.681 0.559 0.852 0.765 0.956 0.921
Table 13: Zero-shot transductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Transductive
NELL995 0.583 0.437 0.578 0.457 0.684 0.555
NELL23k 0.669 0.548 0.756 0.657 0.831 0.762
WDsinger 0.668 0.546 0.720 0.621 0.823 0.738
ConceptNet100k 0.181 0.083 0.650 0.469 0.795 0.658
CoDExSmall 0.900 0.820 0.961 0.935 0.982 0.970
CoDExLarge 0.892 0.824 0.902 0.837 0.973 0.950
YAGO310 0.646 0.403 0.783 0.598 0.971 0.943
AristoV4 0.254 0.201 0.389 0.265 0.597 0.496
DBpedia100k 0.650 0.509 0.717 0.582 0.919 0.861
Hetionet 0.634 0.524 0.809 0.707 0.940 0.890
FB15k-237(10) 0.688 0.550 0.795 0.711 0.918 0.876
FB15k-237(20) 0.695 0.558 0.834 0.758 0.952 0.923
FB15k-237(50) 0.717 0.591 0.876 0.812 0.968 0.946
Table 14: Finetuned inductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Inductive $e,r$
FB-25 0.684 0.563 0.805 0.724 0.909 0.857
FB-50 0.696 0.575 0.780 0.699 0.881 0.820
FB-75 0.754 0.638 0.822 0.699 0.911 0.854
FB-100 0.851 0.769 0.921 0.880 0.965 0.939
WK-25 0.897 0.834 0.905 0.860 0.968 0.954
WK-50 0.865 0.793 0.881 0.840 0.925 0.876
WK-75 0.911 0.875 0.937 0.910 0.965 0.948
WK-100 0.924 0.879 0.916 0.885 0.970 0.946
NL-0 0.632 0.502 0.655 0.518 0.731 0.602
NL-25 0.737 0.622 0.709 0.606 0.757 0.634
NL-50 0.808 0.704 0.774 0.683 0.814 0.721
NL-75 0.795 0.678 0.790 0.671 0.848 0.774
NL-100 0.803 0.678 0.885 0.793 0.937 0.887
MT1 tax 0.990 0.984 0.995 0.990 0.999 0.998
MT1 health 0.929 0.867 0.973 0.949 0.994 0.988
MT2 org 0.981 0.963 0.987 0.978 0.994 0.988
MT2 sci 0.977 0.961 0.990 0.984 0.995 0.992
MT3 art 0.907 0.851 0.887 0.828 0.950 0.916
MT3 infra 0.966 0.947 0.970 0.952 0.996 0.993
MT4 sci 0.954 0.929 0.972 0.952 0.983 0.968
MT4 health 0.951 0.919 0.986 0.979 0.995 0.991
Metafam 0.368 0.112 0.265 0.024 0.997 0.995
FBNELL 0.720 0.576 0.766 0.639 0.879 0.801
Inductive $e$
FB-v1 0.650 0.513 0.705 0.599 0.855 0.766
FB-v2 0.675 0.547 0.713 0.590 0.887 0.812
FB-v3 0.677 0.556 0.742 0.644 0.879 0.810
FB-v4 0.690 0.560 0.766 0.665 0.884 0.807
WN-v1 0.844 0.754 0.776 0.591 0.926 0.879
WN-v2 0.834 0.766 0.765 0.574 0.927 0.869
WN-v3 0.707 0.577 0.756 0.594 0.950 0.911
WN-v4 0.861 0.795 0.804 0.651 0.943 0.898
NL-v1 0.719 0.504 0.590 0.341 0.883 0.766
NL-v2 0.668 0.549 0.811 0.740 0.911 0.870
NL-v3 0.646 0.484 0.757 0.643 0.868 0.795
NL-v4 0.570 0.412 0.822 0.735 0.906 0.849
ILPC Small 0.922 0.876 0.919 0.872 0.953 0.918
ILPC Large 0.875 0.799 0.894 0.829 0.953 0.915
HM 1k 0.626 0.447 0.663 0.414 0.756 0.561
HM 3k 0.592 0.439 0.664 0.418 0.790 0.623
HM 5k 0.605 0.452 0.672 0.428 0.744 0.591
HM Indigo 0.726 0.614 0.835 0.746 0.946 0.903
Table 15: Finetuned transductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Pretraining
FB15k-237 0.795 0.709 0.924 0.870 0.976 0.957
WN18RR 0.914 0.871 0.783 0.634 0.982 0.968
CoDExMedium 0.919 0.870 0.931 0.886 0.974 0.952
Transductive
NELL995 0.630 0.513 0.578 0.457 0.713 0.584
NELL23k 0.688 0.571 0.755 0.658 0.869 0.805
WDsinger 0.730 0.603 0.721 0.627 0.885 0.815
ConceptNet100k 0.612 0.488 0.712 0.551 0.885 0.813
CoDExSmall 0.942 0.900 0.964 0.943 0.981 0.967
CoDExLarge 0.907 0.850 0.908 0.845 0.973 0.950
YAGO310 0.930 0.891 0.826 0.666 0.970 0.942
AristoV4 0.254 0.201 0.498 0.381 0.651 0.547
DBpedia100k 0.650 0.509 0.780 0.665 0.923 0.869
Hetionet 0.737 0.646 0.922 0.862 0.942 0.897
FB15k-237(10) 0.688 0.550 0.795 0.711 0.940 0.905
FB15k-237(20) 0.695 0.558 0.846 0.778 0.958 0.931
FB15k-237(50) 0.728 0.618 0.903 0.858 0.970 0.948
Datasets.
This section provides the full details for all experiments described in the main text. For pretraining, we fit the Flock model on three standard transductive knowledge graph completion benchmarks, following Galkin et al. (2024): FB15k-237 (Toutanova & Chen, 2015), WN18RR (Dettmers et al., 2018), and CoDEx Medium (Safavi & Koutra, 2020). Then, we evaluate zero-shot transfer of both entity prediction and relation prediction, as well as the finetuning performance on multiple datasets grouped as follows:
  • Inductive $e,r$. Link prediction tasks involving previously unseen nodes and relation types. This includes the 13 datasets from Ingram (Lee et al., 2023): FB-25, FB-50, FB-75, FB-100, WK-25, WK-50, WK-75, WK-100, NL-0, NL-25, NL-50, NL-75, NL-100, as well as 10 datasets from MTDEA (Zhou et al., 2023): MT1 tax, MT1 health, MT2 org, MT2 sci, MT3 art, MT3 infra, MT4 sci, MT4 health, Metafam, and FBNELL.
  • Inductive $e$. Link prediction tasks involving novel nodes but fixed relation types. This category comprises 12 GraIL datasets (Teru et al., 2020) (WN-v1 through WN-v4, FB-v1 through FB-v4, NL-v1 through NL-v4), 4 INDIGO benchmarks (Liu et al., 2021) (HM 1k, HM 3k, HM 5k, HM Indigo), and 2 NodePiece datasets (Galkin et al., 2022): ILPC Small and ILPC Large.
  • Transductive. Link prediction tasks where both entities and relations are observed during training. These include CoDEx Small, CoDEx Large (Safavi & Koutra, 2020), NELL-995 (Xiong et al., 2017), YAGO 310 (Mahdisoltani et al., 2015), WDsinger, NELL23k, FB15k-237(10), FB15k-237(20), FB15k-237(50) (Lv et al., 2020), AristoV4 (Chen et al., 2021), DBpedia100k (Ding et al., 2018), ConceptNet100k (Malaviya et al., 2020), and Hetionet (Himmelstein et al., 2017).
Full results of Section 5.2.
Full tables of zero-shot inductive and transductive entity prediction results for Flock are presented in Table 8 and Table 9, and full tables of finetuned performance are given in Table 10 and Table 11. We further provide the complete zero-shot and finetuned relation prediction results in Table 14, Table 13, and Table 15. Table 19 presents the pretraining graph mix shown in Section 5.3. Finally, detailed hyperparameter settings can be found in Table 20, while the number of epochs used for finetuning is summarized in Table 22.
Training.
Following conventions in the literature (Zhu et al., 2021; Huang et al., 2023), for each triple $(h,r,t)$, we add the corresponding inverse triple $(h,r^{-1},t)$, where $r^{-1}$ is a fresh relation symbol. All Flock instances and its variants are optimized to minimize the negative log-likelihood over positive and negative facts under the partial completeness assumption (Galárraga et al., 2013), where negatives are generated by randomly corrupting either the head or the tail entity (for entity prediction) or by corrupting the relation (for relation prediction). To reduce overfitting, we remove edges that directly connect the queried endpoints. The best checkpoint is selected by validation performance. For entity prediction, we take the embedding for potential target $t$ and relations $r$, and obtain the score $p(h,r,t)$ by passing into a 2-layer MLP. For relation prediction, we concatenate the embedding for source $h$, target $t$, and potential relation $r$ to obtain the score $p(h,r,t)$.
Let $(h,r,t)$ be a positive triple and let $k$ denote the number of negatives sampled per positive, where $(h_{i},r,t_{i})$ is the $i$-th negative samples for entity prediction, and $h,r_{i},t_{i}$ is the $i$-th negative samples for relation prediction. Following Sun et al. (2019), we also consider a self-adversarial variant where negatives are reweighted according to their current difficulty. With adversarial temperature $\alpha>0$, the weights for entity and relation prediction, respectively, are
$w^{\mathrm{ent}}_{i,\alpha}=\operatorname{Softmax}\!\left(\frac{\log\!\big(1-p(h_{i}^{\prime},r,t_{i}^{\prime})\big)}{\alpha}\right),\qquad w^{\mathrm{rel}}_{i,\alpha}=\operatorname{Softmax}\!\left(\frac{\log\!\big(1-p(h,r_{i}^{\prime},t)\big)}{\alpha}\right).$
The corresponding losses become
$\mathcal{L}_{\mathrm{ent}}^{\mathrm{adv}}$ $=-\log p(h,r,t)\;-\;\sum_{i=1}^{k}w^{\mathrm{ent}}_{i,\alpha}\,\log\!\big(1-p(h_{i}^{\prime},r,t_{i}^{\prime})\big),$
$\mathcal{L}_{\mathrm{rel}}^{\mathrm{adv}}$ $=-\log p(h,r,t)\;-\;\sum_{i=1}^{k}w^{\mathrm{rel}}_{i,\alpha}\,\log\!\big(1-p(h,r_{i}^{\prime},t)\big).$
Table 16: Dataset statistics for inductive-$e,r$ link prediction datasets. Triples are the number of edges given at training, validation, or test graphs, respectively, whereas Valid and Test denote triples to be predicted in the validation and test graphs.
Dataset Training Graph Validation Graph Test Graph
Entities Rels Triples Entities Rels Triples Valid Entities Rels Triples Test
FB-25 5190 163 91571 4097 216 17147 5716 4097 216 17147 5716
FB-50 5190 153 85375 4445 205 11636 3879 4445 205 11636 3879
FB-75 4659 134 62809 2792 186 9316 3106 2792 186 9316 3106
FB-100 4659 134 62809 2624 77 6987 2329 2624 77 6987 2329
WK-25 12659 47 41873 3228 74 3391 1130 3228 74 3391 1131
WK-50 12022 72 82481 9328 93 9672 3224 9328 93 9672 3225
WK-75 6853 52 28741 2722 65 3430 1143 2722 65 3430 1144
WK-100 9784 67 49875 12136 37 13487 4496 12136 37 13487 4496
NL-0 1814 134 7796 2026 112 2287 763 2026 112 2287 763
NL-25 4396 106 17578 2146 120 2230 743 2146 120 2230 744
NL-50 4396 106 17578 2335 119 2576 859 2335 119 2576 859
NL-75 2607 96 11058 1578 116 $18$ 606 1578 116 $18$ 607
NL-100 1258 55 7832 1709 53 2378 793 1709 53 2378 793
Metafam 1316 28 13821 1316 28 13821 590 656 28 7257 184
FBNELL 4636 100 10275 4636 100 10275 1055 4752 183 10685 597
Wiki MT1 tax 10000 10 17178 10000 10 17178 1908 10000 9 16526 1834
Wiki MT1 health 10000 7 14371 10000 7 14371 1596 10000 7 14110 1566
Wiki MT2 org 10000 10 23233 10000 10 23233 2581 10000 11 21976 2441
Wiki MT2 sci 10000 16 16471 10000 16 16471 1830 10000 16 14852 1650
Wiki MT3 art 10000 45 27262 10000 45 27262 3026 10000 45 28023 3113
Wiki MT3 infra 10000 24 21990 10000 24 21990 2443 10000 27 21646 2405
Wiki MT4 sci 10000 42 12576 10000 42 12576 1397 10000 42 12516 1388
Wiki MT4 health 10000 21 15539 10000 21 15539 1725 10000 20 15337 1703
Table 17: Dataset statistics for inductive-$e$ link prediction datasets. Triples are the number of edges given at training, validation, or test graphs, respectively, whereas Valid and Test denote triples to be predicted in the validation and test graphs.
Dataset Rels Training Graph Validation Graph Test Graph
Entities Triples Entities Triples Valid Entities Triples Test
FB-v1 180 1594 4245 1594 4245 489 1093 1993 411
FB-v2 200 2608 9739 2608 9739 1166 1660 4145 947
FB-v3 215 3668 17986 3668 17986 2194 2501 7406 1731
FB-v4 219 4707 27203 4707 27203 3352 3051 11714 2840
WN-v1 9 2746 5410 2746 5410 630 922 1618 373
WN-v2 10 6954 15262 6954 15262 1838 2757 4011 852
WN-v3 11 12078 25901 12078 25901 3097 5084 6327 1143
WN-v4 9 3861 7940 3861 7940 934 7084 12334 2823
NL-v1 14 3103 4687 3103 4687 414 225 833 201
NL-v2 88 2564 8219 2564 8219 922 2086 4586 935
NL-v3 142 4647 16393 4647 16393 1851 3566 8048 1620
NL-v4 76 2092 7546 2092 7546 876 2795 7073 1447
ILPC Small 48 10230 78616 6653 20960 2908 6653 20960 2902
ILPC Large 65 46626 202446 29246 77044 10179 29246 77044 10184
HM 1k 11 36237 93364 36311 93364 1771 9899 18638 476
HM 3k 11 32118 71097 32250 71097 1201 19218 38285 1349
HM 5k 11 28601 57601 28744 57601 900 23792 48425 2124
HM Indigo 229 12721 121601 12797 121601 14121 14775 250195 14904
Table 18: Dataset statistics for transductive link prediction datasets. Entity task denotes the entity-prediction task: $h/t$ is predicting both heads and tails, and $t$ is predicting only tails.
Dataset Entities Rels Train Valid Test Entity Task
FB15k-237 14541 237 272115 17535 20466 $h/t$
WN18RR 40943 11 86835 3034 3134 $h/t$
CoDEx Small 2034 42 32888 1827 1828 $h/t$
CoDEx Medium 17050 51 185584 10310 10311 $h/t$
CoDEx Large 77951 69 551193 30622 30622 $h/t$
NELL995 74536 200 149678 543 2818 $h/t$
YAGO310 123182 37 1079040 5000 5000 $h/t$
WDsinger 10282 135 16142 2163 2203 $h/t$
NELL23k 22925 200 25445 4961 4952 $h/t$
AristoV4 44949 1605 242567 20000 20000 $h/t$
DBpedia100k 99604 470 597572 50000 50000 $h/t$
ConceptNet100k 78334 34 100000 1200 1200 $h/t$
FB15k-237(10) 11512 237 27211 15624 18150 $t$
FB15k-237(20) 13166 237 54423 16963 19776 $t$
FB15k-237(50) 14149 237 136057 17449 20324 $t$
Hetionet 45158 24 2025177 112510 112510 $h/t$
Table 19: Different graph pretraining mix shown in Section 5.3.
1 2 3 4 5 6 8
FB15k-237 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
WN18RR $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CoDEx Medium $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NELL995 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
YAGO 310 $\checkmark$ $\checkmark$ $\checkmark$
ConceptNet100k $\checkmark$ $\checkmark$
DBpedia100k $\checkmark$
AristoV4 $\checkmark$
Table 20: Hyperparameter for Flock in pretraining and finetuning setup, for entity prediction and relation prediction.
Hyperparameter Entity prediction Relation prediction
Random walk Walk length $\ell$ 128 128
# Pretraining base walk $n_{\text{train}}$ 128 128
# Test-time or finetuning base walk $n$ 16–512 16–512
Sequence processor # Layers 1 1
Hidden dimension 64 64
Consensus protocol # Heads $h$ 4 4
Head dimension $d_{h}$ 16 16
Update # Update step II 6 6
Ensemble # Maximum ensembled passes $P$ 16 16
Pretraining Optimizer AdamW AdamW
Learning rate 0.0005 0.0005
Training steps 400,000 40,000
Adversarial temperature 1 1
# Negatives 512 512
Batch size 8 8
Weight decay 0.01 0.00
Finetuning Optimizer AdamW AdamW
Learning rate 0.0005 0.0005
Adversarial temperature 1 1
# Negatives 256 256
Batch size 4–32 4–8
Table 21: Detailed finetuning and inference hyperparameters for Flock in entity prediction. For each dataset, we report the finetuning epochs, batches per epoch, batch size, and the inference settings for both zero-shot and finetuned modes: test-time ensemble size $P$, base walk count $n$. For Hetionet finetuning we used $(P,n)=(1,1024)$, instead of $(2,512)$ as in zero-shot.
Dataset Epoch # Batch/Epoch Batch Size # Ensembled Passes $P$ # Base Walk $n$
FB15k-237 1 full 8 16 128
WN18RR 1 full 8 16 128
CoDEx Small 1 full 32 16 16
CoDEx Medium 1 full 8 16 128
CoDEx Large 1 2000 4 2 512
NELL-995 1 full 8 16 128
YAGO310 1 2000 4 8 512
WDsinger 1 full 8 16 16
NELL23k 3 full 8 16 32
FB15k-237(10) 1 full 8 16 32
FB15k-237(20) 1 full 8 16 64
FB15k-237(50) 1 full 8 16 64
Hetionet 1 4000 8 2 512
DBpedia100k 1 1000 4 2 512
AristoV4 1 full 8 4 256
ConceptNet100k 1 full 8 16 128
FB v1–v4 1 full 8 16 16
WN v1–v4 1 full 8 16 16
NL v1–v4 3 full 8 16 16
ILPC Small 1 full 8 16 16
ILPC Large 1 full 8 16 64
FB 25–100 3 full 8 16 16
WK 25–100 3 full 8 16 16
NL 0–100 3 full 8 16 16
Wiki MT1 tax 3 full 8 16 16
Wiki MT1 health 3 full 8 16 16
Wiki MT2 org 3 full 16 16 32
Wiki MT2 sci 3 full 8 16 16
Wiki MT3 art 3 full 16 16 32
Wiki MT3 infra 3 full 16 16 32
Wiki MT4 sci 3 full 8 16 16
Wiki MT4 health 3 full 8 16 16
Metafam 3 full 8 16 16
FBNELL 3 full 8 16 16
HM 1k 1 full 8 16 16
HM 3k 1 full 16 16 32
HM 5k 1 full 8 16 64
HM Indigo 1 full 8 16 128
Table 22: Detailed finetuning and inference hyperparameters for Flock in relation prediction. For each dataset, we report the finetuning epochs, batches per epoch, batch size, and the inference settings for both zero-shot and finetuned modes: test-time ensemble size $P$ and base walk count $n$.
Dataset Epoch # Batch/Epoch Batch Size # Ensembled Passes $P$ # Base Walk $n$
FB15k-237 1 1000 8 16 128
WN18RR 1 1000 8 16 128
CoDEx Small 1 1000 8 16 16
CoDEx Medium 1 1000 8 16 128
CoDEx Large 1 1000 4 2 512
NELL-995 1 1000 8 16 128
YAGO310 1 1000 8 4 512
WDsinger 1 1000 8 16 16
NELL23k 1 1000 8 16 32
FB15k-237(10) 1 1000 8 16 32
FB15k-237(20) 1 1000 8 16 64
FB15k-237(50) 1 1000 8 16 64
Hetionet 1 1000 4 2 512
DBpedia100k 1 1000 4 2 512
AristoV4 1 1000 8 4 256
ConceptNet100k 1 1000 8 16 128
FB v1–v4 1 1000 8 16 16
WN v1–v4 1 1000 8 16 16
NL v1–v4 1 1000 8 16 16
ILPC Small 1 1000 8 16 16
ILPC Large 1 1000 8 16 64
FB 25–100 1 1000 8 16 16
WK 25–100 1 1000 8 16 16
NL 0–100 1 1000 8 16 16
Wiki MT1 tax 1 1000 8 16 16
Wiki MT1 health 1 1000 8 16 16
Wiki MT2 org 1 1000 8 16 32
Wiki MT2 sci 1 1000 8 16 16
Wiki MT3 art 1 1000 8 16 32
Wiki MT3 infra 1 1000 8 16 32
Wiki MT4 sci 1 1000 8 16 16
Wiki MT4 health 1 1000 8 16 16
Metafam 1 1000 8 16 16
FBNELL 1 1000 8 16 16
HM 1k 1 1000 8 16 16
HM 3k 1 1000 8 16 32
HM 5k 1 1000 8 16 64
HM Indigo 1 1000 8 16 128

### v2

@@
 Abstract
-We study the problem of zero-shot link prediction on knowledge graphs (KGs), which requires models to generalize over novel entities and novel relations.
-Knowledge graph foundation models (KGFMs) address this task by enforcing equivariance over both nodes and relations, learning from structural properties of nodes and relations, which are then transferable to novel graphs with similar structural properties.
-However, the conventional notion of deterministic equivariance imposes inherent limits on the expressive power of KGFMs, preventing them from distinguishing structurally similar but semantically distinct relations.
-To overcome this limitation, we introduce probabilistic node-relation equivariance, which preserves equivariance in distribution while incorporating a principled randomization to break symmetries during inference.
-Building on this principle, we present Flock, a KGFM that iteratively samples random walks, encodes them into sequences via a recording protocol, embeds them with a sequence model, and aggregates representations of nodes and relations via learned pooling.
-Crucially, Flock respects probabilistic node-relation equivariance and is a universal approximator for isomorphism-invariant link-level functions over KGs.
-Empirically, Flock perfectly solves our new diagnostic dataset Petals where current KGFMs fail, and achieves state-of-the-art performances on entity- and relation prediction tasks on 54 KGs from diverse domains.
+We study the problem of zero-shot link prediction on knowledge graphs (KGs), which requires models to generalize to novel entities and novel relations.
+Knowledge graph foundation models (KGFMs) address this task by enforcing equivariance over both nodes and relations, which enables them to learn structural properties of nodes and relations that transfer to novel KGs with similar structure.
+However, the conventional notion of deterministic equivariance inherently limits the expressive power of KGFMs, as it prevents them from distinguishing relations that are structurally similar but semantically distinct.
+To overcome this limitation, we propose to leverage probabilistic node-relation equivariance, which preserves equivariance in distribution while using structured randomness to break symmetries at inference time.
+Building on this principle, we present Flock, a KGFM that iteratively samples random walks, encodes them into sequences, embeds them with a sequence model, and aggregates node and relation representations through learned pooling.
+Flock respects probabilistic node-relation equivariance and, crucially, is a universal approximator for isomorphism-invariant link-level functions over KGs.
+Empirically, Flock perfectly solves our new diagnostic dataset Petals on which current KGFMs fail, and achieves state-of-the-art performance on entity and relation prediction tasks across 54 KGs from diverse domains.
 1 Introduction
-Knowledge graph foundation models (KGFMs) (Lee et al., 2023; Geng et al., 2023; Galkin et al., 2024; Zhang et al., 2024; Cui et al., 2024; Huang et al., 2025) aim to infer missing links over novel knowledge graphs (KGs) that are not part of the training graphs or domains.
+Knowledge graph foundation models (KGFMs) (Galkin et al., 2024; Zhang et al., 2024; Huang et al., 2025) aim to infer missing links over novel knowledge graphs (KGs) that are not part of the training data or domains.
 This task requires generalization to both unseen nodes and unseen relation types.
-To achieve this, KGFMs rely on learning node and relation invariants: structural properties of nodes and relations that are transferable across KGs even when their relational vocabularies differ.
-This inductive bias is formalized by Gao et al.
-(2023) as double-equivariance — equivariance under permutations of both entities and relations — and used as a core principle in the design of KGFMs in the literature.
+To achieve this, KGFMs learn node and relation invariants: structural properties of nodes and relations that are transferable across KGs even when their relational vocabularies differ.
+This inductive bias is formalized as double equivariance (Gao et al., 2023)—equivariance under permutations of both entities and relations—and used as a core design principle of current KGFMs.
 $\mathsf{Luke}$$\mathsf{Yoda}$$\mathsf{HanSolo}$$\mathsf{Emperor}$$\mathsf{Jabba}$$\mathsf{DarthVader}$$\mathsf{Chewbacca}$$\mathsf{Leia}$
-Figure 1: A KG representing characters’ relationships in Star Wars movies.
-Blue arrows indicate $\mathsf{like}$, red arrows – $\mathsf{dislike}$, and green arrows indicate relation ($\mathsf{friendWith}$).
+Figure 1: A KG representing characters’ relationships in Star Wars movies. Blue arrows indicate $\mathsf{like}$, red – $\mathsf{dislike}$, and green arrows indicate $\mathsf{friendWith}$.
 Problem statement.
-In this work, we challenge the fundamental assumption of existing KGFMs dictated by strict equivariance: structural isomorphism of relations implies semantic equivalence.
-Consider, for example the KG from Figure 1, where the relations like and dislike are structurally isomorphic, and yet they represent semantically opposite relations.
-In this motivating example, any KGFM that computes relation invariants is forced to assign the same representation to both like and dislike — losing the ability to distinguish between two entities with opposite relationships.
-This expressiveness limitation is an architectural one and cannot be resolved through finetuning, which further limits the downstream use of existing KGFMs.
-This raises a central question: How to design KGFMs that are both expressive and have the right inductive bias for generalization?
-Our approach.
-We propose a new approach for KGFMs, which relies on probabilistic node-relation equivariance as inductive bias.
-Instead of enforcing deterministic equivariance over nodes and relations, these KGFMs respect probabilistic node-relation equivariance.
-This relaxes the hard constraint that “structurally isomorphic relations must have identical representations”, and requires only that “the representations of structurally isomorphic relations need to be equivalent in distribution” over a model’s stochastic processes.
-This way, the model retains crucial inductive bias needed for generalizing across different KGs, while the stochasticity of each forward pass ensures that structurally identical but semantically distinct relations are assigned different representations, allowing the model to distinguish between them (Srinivasan & Ribeiro, 2020; Abboud et al., 2021).
+In this work, we challenge a fundamental assumption dictated by strict equivariance in existing KGFMs: structural isomorphism of relations implies semantic equivalence.
+Consider, for example, the KG in Figure˜1, where the relations $\mathsf{like}$ and $\mathsf{dislike}$ are structurally isomorphic yet semantically opposite.
+Any KGFM that computes relation invariants is forced to assign the same representation to both $\mathsf{like}$ and $\mathsf{dislike}$—losing the ability to distinguish entities with opposite relationships.
+This expressiveness limitation is an architectural one and cannot be resolved through finetuning, further limiting the downstream use of existing KGFMs.
+This raises a central question: how can we design KGFMs that are both expressive and have the right inductive bias for generalization?
+Approach.
+We propose a new approach for KGFMs that relies on probabilistic node-relation equivariance as inductive bias, instead of enforcing deterministic equivariance over nodes and relations.
+Probabilistic equivariance relaxes the hard constraint that “structurally isomorphic relations must have identical representations”, and requires only that the representations of structurally isomorphic relations need to be equivalent in distribution over a model’s stochastic predictive processes (Srinivasan & Ribeiro, 2020; Abboud et al., 2021).
+This way, the model retains the inductive bias needed for generalizing across different KGs, while the stochasticity of each forward pass ensures that structurally identical but semantically distinct relations are assigned different representations.
 Inspired by the success of models that learn probabilistic invariants via random walks (Perozzi et al., 2014; Grover & Leskovec, 2016; Nikolentzos & Vazirgiannis, 2020; Kim et al., 2025), we introduce Flock, a KGFM that inherently computes probabilistic node-relation invariants.
-Given a (potentially unseen) KG, and a query, in each iteration, Flock first samples a set of random walks over KG based on the query, noting down both encountered nodes and relations with a recording protocol.
+Given a potentially unseen KG and a query, Flock iteratively samples random walks over the KG based on the query, encoding both encountered nodes and relations with a recording protocol.
 To ensure the model can generalize to unseen entities and relation types, we anonymize all nodes and relations, enforcing that Flock only learn from their structural roles.
 These anonymized sequences are then fed into a sequence processor, and the representations for each node and relation are aggregated via a consensus protocol.
-Finally, we construct per-query (triple) features from the aggregated entity and relation embeddings and input them into a binary classifier for link prediction.
+Finally, we construct per-query features from the aggregated entity and relation embeddings and input them into a binary classifier for link prediction.
 Key findings and contributions.
 The design of Flock offers several key advantages over existing KGFMs.
-First, it entirely abandons the conventional two-stage process of encoding relations and node representations via two separate networks, and does not rely on message-passing at all, thereby avoiding the well-known expressivity limitations of MPNNs on KGs (Barceló et al., 2022; Huang et al., 2023; 2025).
-Second, Flock is a universal approximator (see Proposition 4.1), and thus can approximate every link-level function defined on KGs of any bounded size.
-Finally, Flock architecture inherently respects the principle of probabilistic node-relation equivariance, which enables a strong generalization capacity.
-Our experimental results over both entity prediction and relation prediction validate the strength of this approach, demonstrating that Flock consistently outperforms state-of-the-art KGFMs on existing benchmarks.
-Our contributions can be summarized as follows:
-  • We highlight a limitation in existing KGFMs: their over-reliance on deterministic node–relation equivariance prevents them from distinguishing between structurally similar but semantically different relations, limiting their expressivity.
-  • We introduce probabilistic node-relation equivariance, a property for KGFMs that ensures invariance only in distribution, thus balancing the model expressivity and generalization.
+First, it entirely abandons the conventional two-stage process of encoding relations and node representations via two separate networks, and does not rely on message passing at all, thereby avoiding the well-known expressivity limitations of MPNNs on KGs (Barceló et al., 2022; Huang et al., 2023; 2025).
+Second, Flock is a universal approximator (Proposition 4.1), capable of approximating every link-level function on KGs of any bounded size.
+Finally, Flock’s architecture inherently respects probabilistic node-relation equivariance, enabling strong generalization.
+Our experiments on both entity and relation prediction validate this approach, demonstrating that Flock consistently outperforms state-of-the-art KGFMs on existing benchmarks.
+Our contributions are:
+  • We highlight a limitation of existing KGFMs: their reliance on deterministic node–relation equivariance prevents them from distinguishing between structurally similar but semantically different relations, limiting their expressivity.
+  • We propose to leverage probabilistic node-relation equivariance, a property for KGFMs that ensures equivariance only in distribution, as an effective solution balancing the model expressivity and generalization.
   • We propose Flock, a KGFM that respects probabilistic node-relation equivariance.
-Flock replaces the conventional two-stage, message-passing paradigm with a direct sequence encoding approach based on random walks, and acts as a universal approximator of link-level functions.
-  • We validate our approach on both entity and relation prediction tasks across 54 diverse KGs, where Flock consistently achieves state-of-the-art performance over existing KGFMs.
-We further design a synthetic dataset Petals to confirm our theoretical results empirically.
-All proofs of the theoretical results are provided in Appendix C. The code is available at https://github.com/jw9730/flock-pytorch.
+Flock replaces the conventional two-stage, message-passing paradigm with a direct sequence encoding approach based on random walks, and is a universal approximator of link-level functions.
+  • We validate our approach on entity and relation prediction tasks across 54 diverse KGs, where Flock consistently achieves state-of-the-art performance.
+We further construct a synthetic dataset Petals to confirm our theoretical results empirically.
+All proofs are in Appendix˜B. The code is available at https://github.com/jw9730/flock.
 2 Related work
 Link prediction and KGFMs.
-Early methods for inferring missing links in KGs  (Bordes et al., 2013; Sun et al., 2019; Balazevic et al., 2019; Abboud et al., 2020; Schlichtkrull et al., 2018; Vashishth et al., 2020) rely on learned embeddings, hence operating in the transductive setting, incapable of generalizing to unseen entities or relation types.
-Later GNN-based approaches based on the labeling trick (Teru et al., 2020; Zhang et al., 2021) or conditional message passing (Zhu et al., 2021; 2023; Zhang & Yao, 2022; Zhang et al., 2023b; Huang et al., 2023), unlocked the node inductive scenario, while remaining restricted to a fixed relational vocabulary.
-KGFMs eliminate this restriction and enable node-relation inductive link prediction over both unseen nodes and relation types through the use of a two-stage process by first encoding relations and then nodes.
-The first examples of this paradigm are InGram (Lee et al., 2023) and ULTRA (Galkin et al., 2024).
-Their ideas were extended by TRIX (Zhang et al., 2024) to build a more expressive framework.
-KG-ICL (Cui et al., 2024) achieved full inductivity by combining in-context learning with node-relation tokenization.
-ISDEA (Gao et al., 2023) and MTDEA (Zhou et al., 2023) highlighted the benefits of equivariance over both nodes and relations.
-MOTIF (Huang et al., 2025) was proposed as a general KGFM framework, supported by a theoretical analysis of the expressive power of KGFMs.
-Our work further advances the field with the first stochastic KGFM, which is invariant in probability and provably more expressive than all the existing methods.
-Notably, Flock achieves universality without any form of message passing, instead relying on random walks and sequence models to encode both nodes and relations anonymously to ensure generalization.
+Early methods for inferring missing links in KGs relied on learned embeddings (Bordes et al., 2013; Sun et al., 2019; Balazevic et al., 2019; Abboud et al., 2020; Schlichtkrull et al., 2018; Vashishth et al., 2020), operating in the transductive setting and incapable of generalizing to unseen entities or relation types.
+Later MPNN-based methods based on the labeling trick (Teru et al., 2020; Zhang et al., 2021) or conditional message passing (Zhu et al., 2021; 2023; Zhang & Yao, 2022; Zhang et al., 2023b; Huang et al., 2023) unlocked the node-inductive scenario, while remaining restricted to a fixed relational vocabulary.
+KGFMs eliminate this restriction and enable node-relation inductive link prediction over both unseen nodes and relations (Geng et al., 2023), typically by first encoding relations and then nodes.
+Early examples are InGram (Lee et al., 2023) and ULTRA (Galkin et al., 2024), later extended by TRIX (Zhang et al., 2024) into a more expressive framework.
+KG-ICL (Cui et al., 2024) achieved full inductiveness by combining in-context learning with node-relation tokenization.
+ISDEA (Gao et al., 2023) and MTDEA (Zhou et al., 2023) highlighted the benefits of node-relation equivariance, while MOTIF (Huang et al., 2025) proposed a general KGFM framework with a theoretical analysis of their expressive power.
+Our work advances the field with a stochastic KGFM that achieves invariance in probability to ensure generalization while being provably more expressive than the existing methods.
+Notably, Flock achieves universality without any form of message passing, instead relying on random walks and sequence models to encode both nodes and relations anonymously.
+This is distinct from prior stochastic KGFMs that rely on random initialization of message passing (Lee et al., 2023; Gao et al., 2023).
 Random walks for graph representations.
-Random walks have attracted a lot of attention in graph learning, due to their simplicity and ability to gather context from neighborhoods.
-DeepWalk (Perozzi et al., 2014) and node2vec (Grover & Leskovec, 2016) were among the first to explore the potential of random walks for producing graph embeddings, treating walks as analogues of sentences in natural language and processing them using skip-gram models.
-Nikolentzos & Vazirgiannis (2020) generated graph-level task predictions by executing joint random walks on direct products of graphs with their extracted subgraphs.
-CRaWL (Tönshoff et al., 2021) represents the input graph as a collection of random walks and processes them with a 1-dimensional convolutional NN.
-WalkLM (Tan et al., 2023) samples random walks from graphs with textual features, passing them to a language model for embedding generation.
-RWNN (Kim et al., 2025) and RUM (Wang & Cho, 2024) anonymize the extracted walks and process them with sequence models and RNNs, respectively.
-NeuralWalker (Chen et al., 2025) aggregates embeddings derived by encoding random walks into message passing layers.
+Random walks have been widely adopted in graph learning due to their simplicity and ability to gather context from neighborhoods.
+DeepWalk (Perozzi et al., 2014) and node2vec (Grover & Leskovec, 2016) were among the first approaches, treating walks as analogues of natural language sentences and processing them with skip-gram models.
+Subsequent work has developed diverse neural architectures based on random walks: Nikolentzos & Vazirgiannis (2020) generates graph-level predictions using joint walks on direct products of graphs and their subgraphs, CRaWL (Tönshoff et al., 2021) represents a graph as a collection of walks and processes them with a convolutional network, WalkLM (Tan et al., 2023) samples walks from graphs with textual features and embeds them using a language model, RWNN (Kim et al., 2025) and RUM (Wang & Cho, 2024) anonymize walks and process them with sequence networks, and NeuralWalker (Chen et al., 2025) integrates walk encodings into message passing.
 Probabilistic invariance.
@@
 Despite their stochasticity, such methods can remain probabilistically invariant, ensuring that equivalent inputs yield identical expected outputs, or even identical output distributions.
-We extend the notion of probabilistic invariance to KGs and prove that Flock satisfies invariance in distribution.
+In line with a prior work (Gao et al., 2023), we extend the notion of probabilistic invariance to KGs and prove that Flock satisfies invariance in distribution, and further clarify its usefulness in KG learning.
 3 Preliminary
@@
 We denote by $R^{-1}$ the set of inverses of relations $R$, defined as $\{r^{-1}\mid r\in R\}$, and mean $r$ when writing $(r^{-1})^{-1}$.
+Further, let $\mathbb{K}_{n,m}$ be the space of knowledge graphs with $n$ vertices and $m$ relation types.
 Isomorphism.
@@
 Probabilistic invariance.
-Let $\mathbb{K}_{n,m}$ be the space of knowledge graphs with $n$ vertices and $m$ relation types.
-A stochastic KG model $\varphi$ can be viewed as a function that takes a KG $G$ as the input and returns a random variable $\varphi(G)$.
+A stochastic KG model $\varphi$ can be viewed as a function that takes a KG and returns a random variable $\varphi(G)$.
 Following Kim et al.
 (2025), we call $\varphi$ invariant in probability if
-$\forall G,G^{\prime}\in\mathbb{K}_{n,m}:\qquad G\simeq G^{\prime}\implies\varphi(G)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\varphi(G^{\prime})$
-i.e. the distributions of $\varphi(G)$ and $\varphi(G^{\prime})$ are equal. In particular, this implies $\mathbb{E}[\varphi(G)]=\mathbb{E}[\varphi(G^{\prime})]$.
+$\forall G,G^{\prime}\in\mathbb{K}_{n,m}:\qquad G\simeq G^{\prime}\implies\varphi(G)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\varphi(G^{\prime}),$
+i.e., the distributions of $\varphi(G)$ and $\varphi(G^{\prime})$ are equal. In particular, this implies $\mathbb{E}[\varphi(G)]=\mathbb{E}[\varphi(G^{\prime})]$.
 4 Methodology
 […]
-Figure 2: Overall pipeline of Flock.
-In each updating step, Flock samples random walks on the KG, anonymizes the encountered nodes and relations via a recording protocol, and feeds the sequences in a sequence processor to compute node and relation representations.
-A consensus protocol then pools them back to the original KG’s nodes and relations.
+Figure 2: An overview.
+In each updating step, Flock (1) samples random walks on the KG (two walks indicated by red and teal, respectively), (2) anonymizes the encountered nodes and relations via a recording protocol (for each walk, nodes are anonymized as $1,2,$...
+and relations as $\alpha,\beta,$...), and (3) feeds the sequences in a sequence processor to compute node and relation representations.
+(4) A consensus protocol then pools them back to the original KG’s nodes and relations.
 We present Flock, a KGFM respecting probabilistic node-relation invariance.
@@
 For relation prediction, it outputs $\hat{{\bf y}}:R\to[0,1]$ such that a link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(r)$.
-In practice, we often ensemble repeated predictions $\hat{\bf y}_{1},...,\hat{\bf y}_{P}\sim X_{\theta}(G,q),P\in\mathbb{N}$ via averaging.
-We describe the architecture of Flock in Section 4.1 focused on four main components, and then analyze its theoretical properties in Section 4.2, showing universality and probabilistic equivariance.
+At test time, we average multiple ($P$) independent stochastic predictions to produce the final output. This improves performance and reduces variance through an ensembling effect.
+We describe the architecture of Flock in Section˜4.1 focused on four main components, and then analyze its theoretical properties in Section˜4.2, showing universality and probabilistic equivariance.
+An expansion on the model details can be found in Appendix˜A.
 4.1 Flock
@@
 Each update is done residually using a randomized function $U_{\theta_{i}}$:
-${\bf v}^{(i+1)}\coloneqq{\bf v}^{(i)}+\Delta{\bf v},\quad{\bf r}^{(i+1)}\coloneqq{\bf r}^{(i)}+\Delta{\bf r},\quad(\Delta{\bf v},\Delta{\bf r})\sim{\rm update}_{\theta_{i}}({\bf v}^{(i)},{\bf r}^{(i)}).$ (1)
+${\bf v}^{(i+1)}\coloneqq{\bf v}^{(i)}+\Delta{\bf v},\qquad{\bf r}^{(i+1)}\coloneqq{\bf r}^{(i)}+\Delta{\bf r},\qquad(\Delta{\bf v},\Delta{\bf r})\sim{\rm update}_{\theta_{i}}({\bf v}^{(i)},{\bf r}^{(i)}).$
 The final hidden states ${\bf v}^{(I)}:V\to\mathbb{R}^{d}$ and ${\bf r}^{(I)}:R\to\mathbb{R}^{d}$ are then processed by a binary classifier ${\rm head}:\mathbb{R}^{d}\to[0,1]$ to produce the output $\hat{\bf y}$ which is $V\to[0,1]$ or $R\to[0,1]$ depending on task.
@@
   1. 1.
-    Random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $l$ on the input KG.
+    Random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $\ell$ on the input KG.
   2. 2.
@@
 For the ease of exposition, we explain for entity prediction tasks $q=(h,r,?)$, but relation prediction is similar.
-Random walks. In Flock, random walks are central in two ways: they rewrite the connectivity of nodes and relations as sequences, and support generalization via probabilistic equivariance.
-Formally, the random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $l$ on KG $G$. Each random walk $\eta$ is a chain of random variables, written as:n as:
-$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l}}v_{l},\quad v_{s}\in V,r_{s}\in R,(v_{s-1},r_{s},v_{s})\in E,$ (2)
-where the underlying transition mechanism and $l$ are hyperparameters.
+Random walks.
+In Flock, random walks are central in two ways: they rewrite the connectivity of nodes and relations as sequences, and support generalization via probabilistic equivariance.
+Formally, the random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $\ell$ on KG $G$. Each random walk $\eta$ is a chain of random variables, written as:n as:
+$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}v_{\ell},\qquad v_{s}\in V,r_{s}\in R,(v_{s-1},r_{s},v_{s})\in E,$
+where the underlying transition mechanism and $\ell$ are hyperparameters.
 To support probabilistic equivariance, we ask the walk algorithm to be invariant in probability.
 We say $\eta$ is invariant in probability if for any $G\simeq H$ in $\mathbb{K}_{n,m}$ with isomorphism $(\pi,\phi)$ from $G$ to $H$:
-$\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\cdots\xrightarrow{\phi(r_{l})}\pi(v_{l})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\cdots\xrightarrow{s_{l}}u_{l}$ (3)
-where $v_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{l}}v_{l}$ and $u_{0}\xrightarrow{s_{1}}\cdots\xrightarrow{s_{l}}u_{l}$ follow the distributions of $\eta(G,l)$ and $\eta(H,l)$, respectively.
+$\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\cdots\xrightarrow{\phi(r_{\ell})}\pi(v_{\ell})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\cdots\xrightarrow{s_{\ell}}u_{\ell},$
+where $v_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}v_{\ell}$ and $u_{0}\xrightarrow{s_{1}}\cdots\xrightarrow{s_{\ell}}u_{\ell}$ follow the distributions of $\eta(G,\ell)$ and $\eta(H,\ell)$, respectively.
 In such case, the isomorphism $(\pi,\phi)$ yields a natural translation from walks in $G$ to $H$.
-In Flock, we use a simple random walk algorithm, which we show to be invariant in probability, while working robustly in practice.
-Specifically, given a base walk count $n$, for entity prediction queries $(h,r,?)$, we use $3n$ walks and choose three types of start locations.
-We start $n$ walks at query node $h$ by fixing $v_{0}=h$, start $n$ walks by choosing a random relation $s$, then a random edge $(v,s,u)\in E$, and then fixing the first step as $v\xrightarrow{s}u$, and start the rest of $n$ walks at random nodes in $V$.
-For relation prediction queries $(h,?,t)$, we additionally start $n$ walks at the tail node $t$ by fixing $v_{0}=t$, sampling a total of $4n$ walks.
-For the transition mechanism, we use uniform walks with non-backtracking, with minor modifications to handle directed multi-edges of KGs.
-Despite the simplicity, we find that this choice works well in practice, consistent with findings of prior work (Tönshoff et al., 2021; Kim et al., 2025).
+In Flock, we use a simple random walk algorithm which we show to be invariant in probability.
+Specifically, we use uniform walks with non-backtracking, with minor modifications to handle directed multi-edges of KGs.
+Despite the simplicity, we find that this choice works well in practice, consistent with findings of prior works (Tönshoff et al., 2021; Kim et al., 2025).
+Under this choice, we diversify the starting locations of walks such that local context around the query $q$ and broad regions of the nodes and relations in a KG are both well-captured.
+Our diversification strategy is as follows: given a base walk count $n$, for entity prediction queries $(h,r,?)$, we use $3n$ walks with three types of start locations.
+The first $n$ walks start at query node $h$, capturing local context around the query; the second $n$ walks start by traversing a random edge $(v,s,u)$ where $s$ is a uniformly chosen relation, broadly capturing the relations of the KG including $r$; the last $n$ walks start at random nodes, broadly capturing various regions of the KG.
+For relation prediction queries $(h,?,t)$, we additionally start $n$ walks at the tail node $t$, sampling a total of $4n$ walks.
 We lastly discuss how to choose the base walk count $n$.
@@
 We thus propose test-time adaptation of walk counts, and use:
-$n=n_{\rm train}\times{\rm harmonic}\ {\rm mean}\left(\frac{|V|}{|V|_{\rm train}},\frac{|E|}{|E|_{\rm train}}\right)$ (4)
+$n=n_{\rm train}\times{\rm harmonic}\ {\rm mean}\left(\frac{|V|}{|V|_{\rm train}},\frac{|E|}{|E|_{\rm train}}\right),$ (1)
 where $|V|_{\rm train},|E|_{\rm train}$ are average numbers of nodes and edges in pretraining KGs, respectively.
 Intuitively, this scales $n$ proportionally to the sizes of test KGs relative to pretraining.
-In practice, we clamp $n$ to the nearest power of $2$ in an interval to keep GPU memory usage in a range.
+In practice, we clamp $n$ to the nearest power of $2$ and limit its value in an interval to avoid out-of-memory errors.
 Recording protocol.
@@
 for relations:
-$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}v_{2}\xrightarrow{r_{1}^{-1}}v_{0}\qquad\mapsto\qquad 1\xrightarrow{\alpha}2\xrightarrow{\beta}3\xrightarrow{\alpha^{-1}}1,$ (5)
+$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}v_{2}\xrightarrow{r_{1}^{-1}}v_{0}\qquad\mapsto\qquad 1\xrightarrow{\alpha}2\xrightarrow{\beta}3\xrightarrow{\alpha^{-1}}1,$
 where $(\cdot)^{-1}$ marks direction of a relation.
@@
 Specifically, we use bidirectional GRU (Cho et al., 2014) equipped with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU feedforward network (Shazeer, 2020), which provided robust results.
+To convert anonymizations into input feature vectors to the GRU, we use trainable embedding tables.
 Given that $f_{\theta}$ is a sequence network, it is convenient to interpret its output $\bf h$} as positionally aligned with each step of the walk $\eta$ or record $\bf z$}.
 Specifically, for the example in Equation 2, we obtain:
-$f_{\theta}:{\bf z}\mapsto{\bf h}=(\Delta{\bf v}_{0},a_{0})\xrightarrow{\Delta{\bf r}_{1},b_{1}}(\Delta{\bf v}_{1},a_{1})\xrightarrow{\Delta{\bf r}_{2},b_{2}}\cdots.$ (7)
+$f_{\theta}:{\bf z}\mapsto{\bf h}=(\Delta{\bf v}_{0},a_{0})\xrightarrow{\Delta{\bf r}_{1},b_{1}}(\Delta{\bf v}_{1},a_{1})\xrightarrow{\Delta{\bf r}_{2},b_{2}}\cdots.$
 where $\Delta{\bf v}_{s},\Delta{\bf r}_{s}\in\mathbb{R}^{h\times d_{h}}$ and $a_{s},b_{s}\in\mathbb{R}^{h}$ are the decoded outputs at each position using linear projections.
@@
 Consensus protocol.
-After sequence processing, we are left with a handful of state update proposals ${\bf h}_{1:N}$ from $f_{\theta}$, that are positionally aligned with random walks $\eta_{1:N}$ on KG $G=(V,E,R)$.
+After sequence processing, we are left with a handful of state update proposals ${\bf h}_{1:N}$ from $f_{\theta}$, that are positionally aligned with random walks $\eta_{1:N}$ on KG.
 The consensus protocol $c$ uses the information to decide final state updates $\Delta{\bf v}:V\to\mathbb{R}^{d}$ and $\Delta{\bf r}:R\to\mathbb{R}^{d}$.
@@
 Again, we can show that the consensus protocol does not operate in a way specific to particular KGs, and hence retains node-relation equivariance.
-4.2 Theoretical properties of Flock
+4.2 Theoretical analysis
 Expressivity.
 Following the notion of probabilistic expressivity introduced by Abboud et al.
-(2021), we say that a Flock model $X_{\theta}$ is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ if for any link invariant $\varphi:\mathbb{K}_{n,m}\rightarrow(V\times R\times V\rightarrow[0,1])$ and any $\epsilon,\delta>0$, there exists a choice of the network parameters $\theta$ and the length of the sampled random walks $l$, such that:
+(2021), we say that a Flock model $X_{\theta}$ is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ if for any link invariant $\varphi:\mathbb{K}_{n,m}\rightarrow(V\times R\times V\rightarrow[0,1])$ and any $\epsilon,\delta>0$, there exists a choice of the network parameters $\theta$ and the length of the sampled random walks $\ell$, such that:
 ${\mathbb{P}}(|\varphi(G)((h,r,t))-X_{\theta}(G,(h,r,?))(t)|<\epsilon)>1-\delta$
@@
 With a powerful enough sequence processor $f_{\theta}$, the Flock framework described above is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
+All proofs are in Appendix˜B. To offer an intuition behind the result, we provide a proof sketch.
+Proof sketch.
+A sufficiently long random walk will cover all edges of the graph with high probability.
+Then, from its anonymized version, assigning unique positional identifiers to every node and relation, one can reconstruct the input graph, up to isomorphism.
+Thus, with a sufficiently expressive sequence processor, Flock can approximate any link-invariant function.
 Invariance. Despite the stochastic nature of our framework, beyond randomized node embeddings (Abboud et al., 2021), Flock can be provably guaranteed to satisfy probabilistic invariance:
@@
 Then, regardless of the choice of the deterministic sequence processor $f_{\theta}$, the corresponding Flock model is invariant in probability.
+Proof sketch. Since each of these components is invariant (in probability), and invariance of individual component is preserved under composition, we have that Flock is invariant.
 Moreover, the designs of Flock’s components provided earlier in this section satisfy the conditions of Proposition 4.2. Therefore, the suggested pipeline is indeed invariant in probability:
 Proposition 4.3.
-Any Flock model with components as outlined in this section, and detailed in Appendix B is invariant in probability.
+Any Flock model with components as outlined in this section, and detailed in Appendix˜A is invariant in probability.
 5 Experiments
-We evaluate Flock over a wide range of KGs with both entity and relation prediction tasks, aiming to answer the following questions:
+We test Flock over a wide range of KGs for both entity and relation predictions, aiming to answer:
   • Q1.
@@
   • Q3.
-    How does performance scale with the pretraining graph mix and the size of test-time ensemble?
+    How does performance scale with the sizes of pretraining graph mix and test-time ensemble?
   • Q4.
-    What is the impact of test-time adaptation of walk counts?
-  • Q5.
-    How do current KGFMs perform under noise injection, and how do they compare with Flock? (see Appendix F)
-In addition to these experiments, we report the computational complexity of Flock in Appendix D and conduct a detailed scalability analysis in Appendix E.
-We report ensembled predictions by averaging $P$ independent stochastic forward passes to reduce variance.
-Please refer to the detailed per-dataset choices in Table 22.
-5.1 Synthetic datasets experiments
+    What is the impact of choices of each component on the behavior and performance of Flock?
+We further provide detailed scalability analysis in Appendix˜D and Appendix˜E, and comparisons against current KGFMs augmented with noise injection in Appendix˜F.
+Experiment details and hyperparameters are in Appendix˜I.
+5.1 Synthetic dataset
 […]
-Figure 3: Example KG from Petals. KGFMs with relational invariants must equate blue $r_{1}$ and red $r_{2}$, thus predicting the same scores for both dashed queries with $r_{0}$.
+Figure 3: A KG in Petals. KGFMs with relational invariants must equate $r_{1}$ and $r_{2}$, thus predicting the same scores for dashed queries with $r_{0}$.
 Setup.
-We construct a synthetic benchmark Petals to validate the limitations of KGFMs following the node-relation equivariance (Q1).
-The Petals benchmark contains $220$ instances, with each instance including the following elements: 1) a KG $G=(V,E,R)$, which consists of a ‘central’ node $s\in V$, a ‘stem’ $T\subset V$ with query relation $r_{0}\in R$, and multiple cyclic ‘petals’, each ‘colored’ with a different pair of relations in $R\setminus\{r_{0}\}$, 2) an entity prediction query $(h,r_{0},?)$ with $h\in\{s\}\cup T$, and 3) two candidate targets $t_{1}$ and $t_{2}$ from the same ‘petal’, located at the same distance from $s$.
-An example is shown in Figure 3.
-See Appendix A for more details.
-Petals is designed such that each instance always admits non-trivial automorphisms, meaning that swapping relations occurring in the same ‘petal’ will result in an isomorphic KG.
+To validate the limitations of KGFMs with node-relation equivariance (Q1), we construct a synthetic benchmark Petals.
+It contains $220$ instances, each including: (1) a KG $G=(V,E,R)$ consisting of a ‘central’ node $s$, a ‘stem’ $T\subset V$ with query relation $r_{0}$, and multiple cyclic ‘petals’, each ‘colored’ with a different pair of relations in $R\setminus\{r_{0}\}$, (2) an entity prediction query $(h,r_{0},?)$ with $h\in\{s\}\cup T$, and (3) two candidate targets $t_{1}$ and $t_{2}$ from the same ‘petal’, located at the same distance from $s$.
+An example is in Figure˜3.
+See Appendix˜C for more details.
+Petals is designed such that each instance always admits non-trivial automorphisms, meaning that swapping relations occurring in the same ‘petal’ results in an isomorphic KG.
 Consequently, any model computing relation invariants will not be able to distinguish between potential links $(s,r_{0},t_{1})$ and $(s,r_{0},t_{2})$.
 However, the samples are constructed so that these links are not isomorphic from the graph perspective, making them distinguishable for general link-invariant functions.
-We say a model successfully solves an instance if it can classify $(s,r_{0},t_{1})$ as true and $(s,r_{0},t_{2})$ as false, and we measure the accuracy in the experiment.
-Table 1: Accuracy of KGFMs of our Petals benchmark.
+We say a model solves an instance if it can classify $(s,r_{0},t_{1})$ as true and $(s,r_{0},t_{2})$ as false.
+For empirical validation, we train Ultra (Galkin et al., 2024), $\textsc{Motif}({\mathcal{F}}_{\text{Path}}^{3})$ (Huang et al., 2025), Trix (Zhang et al., 2024), and Flock from scratch and validate them on the training instances.
+Table 1: Petals accuracies.
 Model Petals
@@
 Flock 100%
-We trained Ultra (Galkin et al., 2024), $\textsc{Motif}({\mathcal{F}}_{\text{Path}}^{3})$ (Huang et al., 2025), Trix (Zhang et al., 2024), and Flock from scratch and validated them on the training instances.
 Results.
-We present the results in Table 1.
+The results are provided in Table˜1.
 As expected, all existing KGFMs relying on learning deterministic relational invariants fail to distinguish between the candidate target triplets completely, achieving $50\%$ accuracy due to random guesses.
@@
 5.2 Entity and relation prediction over knowledge graphs
-Setup.
-We follow the protocol of Galkin et al.
-(2024); Zhang et al.
-(2024) and pretrain Flock on FB15k-237 (Toutanova & Chen, 2015), WN18RR (Dettmers et al., 2018), and CoDEx Medium (Safavi & Koutra, 2020).
-We then evaluate its zero-shot and finetuned inference performance with the test set of 54 KG datasets (see Table 15 for details).
-These KG datasets are extracted from diverse domains across three settings: inductive on nodes and relations (Inductive $e,r$), inductive on nodes (Inductive $e$), and transductive.
-Note that these settings differ only during finetuning setup; in zero-shot setup, all entities and relations are unseen.
-We choose two state-of-the-art KGFMs Ultra (Galkin et al., 2024) and Trix (Zhang et al., 2024) as baselines, since they are pretrained on the same pretraining graph mix to ensure a fair comparison.
-Following conventions in the literature (Zhu et al., 2021; Huang et al., 2023), for each triple $(h,r,t)$, we add the corresponding inverse triple $(t,r^{-1},h)$, where $r^{-1}$ is a fresh relation symbol.
-For entity prediction, we report both head and tail results for each triple on all datasets, except for three from Lv et al.
-(2020) where only tail results are available, following Zhang et al.
-(2024).
-For evaluation, we use the filtered ranking protocol (Bordes et al., 2013), reporting Mean Reciprocal Rank (MRR) and Hits@10 for entity prediction, and Hits@1 for relation prediction, as some datasets have fewer than $10$ relations.
-Detailed per-dataset results are shown in Appendix G.
-Entity prediction results.
-We present the average entity prediction results in Table 2 (Q2).
-In the zero-shot setting, Flock consistently outperforms Ultra and Trix on all metrics, demonstrating its strong generalization on KGs over diverse domains.
-Notably, on Metafam (Zhou et al., 2023), a dataset designed to challenge models with conflicting and compositional relational patterns, Flock roughly doubles MRR over Ultra and achieves about a 40% MRR gain over Trix in zero-shot performance.
-These gains align with our hypothesis that probabilistic node–relation equivariance improves expressivity without sacrificing generalization.
-In the finetuning setting, we observe a similar pattern: Flock maintains a consistent improvement over all datasets except transductive splits, where the underlying KGs are generally larger.
-We hypothesize that this gap stems from random walk coverages.
-Unlike Ultra and Trix whose message passing guarantees a full $k$-hop neighborhood coverage over the queried node, Flock relies on sampling random walks, which may not fully cover the target nodes of interest.
 Table 2: Average entity prediction MRR and Hits@10 over 54 KGs from distinct domains.
@@
 Flock (finetuned) 0.929 0.889 0.887 0.808 0.897 0.844 0.907 0.851 0.977 0.959
-Relation prediction results.
-Table 3 presents the average MRR and Hits@1 for relation prediction results (Q2).
-We observe that Flock substantially outperforms all existing KGFMs across all categories in the zero-shot setting, achieving an $11.2\%$ relative improvement in MRR compared to the best baseline Trix.
+Setup.
+To answer Q2, we follow the protocol of Galkin et al.
+(2024); Zhang et al.
+(2024) and pretrain Flock on FB15k-237 (Toutanova & Chen, 2015), WN18RR (Dettmers et al., 2018), and CoDEx Medium (Safavi & Koutra, 2020).
+We then evaluate its zero-shot and finetuned inference performance with the test set of 54 KGs (see Appendix˜I for details).
+These KGs are extracted from diverse domains across three settings: inductive on nodes and relations (Inductive $e,r$), inductive on nodes (Inductive $e$), and transductive.
+Note that these settings differ only during finetuning setup; in zero-shot setup, all entities and relations are unseen.
+We choose state-of-the-art KGFMs Ultra (Galkin et al., 2024) and Trix (Zhang et al., 2024) as baselines, as they are pretrained on the same KGs, to ensure a fair comparison.
+For evaluation, we use the filtered ranking protocol (Bordes et al., 2013), reporting mean reciprocal rank (MRR) and Hits@10 for entity prediction, and Hits@1 for relation prediction, as some KGs have fewer than 10 relations.
+Per-dataset results are in Appendix˜I.
+Entity prediction.
+Table˜2 shows the entity prediction results.
+In the zero-shot setting, Flock consistently outperforms Ultra and Trix, demonstrating strong generalization across diverse domains.
+Notably, on Metafam (Zhou et al., 2023), a dataset designed to challenge models with conflicting and compositional relational patterns, Flock roughly doubles MRR over Ultra and achieves $\approx$ 40% gain in MRR over Trix in the zero-shot setting.
+We find that Flock distinguishes structurally similar but semantically conflicting relations while Ultra fails (Appendix˜G), explaining the gain.
+These findings align with our hypothesis that probabilistic equivariance improves expressivity without sacrificing generalization.
+In the finetuning setting, we observe a similar pattern: Flock maintains a consistent improvement over all datasets except transductive ones, where KGs are generally larger.
+We hypothesize that the gap stems from walk coverages.
+Unlike Ultra and Trix, whose message passing guarantees a full neighborhood coverage around the queried node, Flock relies on sampled walks that may not fully cover target nodes of interest.
+We find that Flock favors sparse KGs (Appendix˜H), consistent with this hypothesis as random walks cover sparse graphs faster.
+Relation prediction.
+Table˜3 presents the relation prediction results.
+Flock substantially outperforms all existing KGFMs across all categories in the zero-shot setting, achieving an $11.2\%$ relative improvement in MRR compared to the best baseline Trix.
 Flock shows a further performance boost of $12.8\%$ in the finetuned setting.
-We attribute this huge performance gain to Flock’s joint encoding of entities and relations during the updating step using the sequence encoder, while existing KGFMs, Ultra and Trix, require separate updating steps for entities and relations.
-This joint updating mechanism yields more holistic representations of both entities and relations with minimal information loss.
-5.3 Scaling behavior of Flock
-Pretraining graph mixing scaling.
+We attribute this huge gain to Flock’s joint encoding of entities and relations during the updating step via the sequence encoder, while existing KGFMs, Ultra and Trix, have separate update steps for entities and relations.
+This joint update mechanism yields more holistic representations of both entities and relations with minimal information loss.
+5.3 Scaling analysis
+Size of pretraining graph mix.
 To assess whether Flock benefits from more pretraining graph and data (Q3), we follow the setup of Galkin et al.
@@
 We then evaluate them on all $41$ inductive benchmarks for a fair comparison.
-We present the detailed pretraining graph mix in Table 19.
-As shown in Figure 4(a), Flock’s generalization improves consistently as the number of pretraining KGs increases, exhibiting clear scaling behavior, which is a core characteristic of being a foundation model.
+We present the detailed pretraining graph mix in Table˜16.
+As shown in Figure˜4(a), Flock’s generalization improves consistently as the number of pretraining KGs increases, exhibiting clear scaling behavior, which is a core characteristic of being a foundation model.
 (a) Zero-shot MRR vs. #pretraining graphs.
@@
 To assess how test-time ensemble size $P$ affects performance (Q3), we take the pretrained Flock and run zero-shot entity prediction on 41 inductive KGs by increasing the number of ensembled passes.
-As shown in Figure 4(b), performance improves from 1 to 8 passes and then begins to plateau beyond 12.
+As shown in Figure˜4(b), performance improves from 1 to 8 passes and then begins to plateau beyond 12.
 This indicates a clear scaling behavior: larger ensembles provide a more accurate estimate of the underlying node and relation distributions.
-5.4 Impact of Test-Time Adaptation of Walk Counts
-Table 4: Zero-shot entity prediction of Flock with and without adaptive test-time walks. We show the average number of entities $|V|$, triples $|E|$, base walks $n$, MRR, and Hits@10.
-Dataset split Statistics Flock Flock w/o Adap.
+5.4 Ablation studies
+Table 4: Ablation study of adaptive test-time walks with zero-shot entity prediction task. We show the average number of entities $|V|$, triples $|E|$, base walks $n$, MRR, and Hits@10.
+Dataset split Statistics Flock Flock w/o Adap. Equation˜1
 $|V|$ $|E|$ $n$ MRR Hits@10 $n$ MRR Hits@10
@@
 Transductive 47,810 387,491 214.15 0.340 0.509 128 0.334 0.493
-Recall that we employ test-time adaptation of walk counts, which adaptively selects the base walk count $n$ based on the graph size, computed via the harmonic-mean rule shown in Equation 4 during inference.
-To answer Q4, we conduct an ablation study over this adaptive mechanism.
-Table 4 compares this adaptive setting with a fixed setting that uses $128$ base walks per sample for all datasets, matching the pretraining setup ($n_{\text{train}}=128$).
+Table 5: Detailed ablation study with zero-shot entity prediction task. For the transductive split, considering resource limits, we test NELL995, NELL23k, WDsinger, ConceptNet100k, and YAGO310.
+Inductive $e,r$ Inductive $e$ Transductive Total Avg Model (23 graphs) (18 graphs) (5 graphs) (46 graphs) MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 Flock ($\ell=128$) 0.369 0.554 0.456 0.604 0.360 0.542 0.395 0.567 w/o non-backtracking 0.370 0.549 0.456 0.605 0.334 0.499 0.386 0.551 $\ell=64$ 0.372 0.556 0.459 0.606 0.351 0.534 0.394 0.565 $\ell=256$ 0.360 0.548 0.458 0.605 0.338 0.508 0.385 0.553 w/o diverse starts 0.360 0.539 0.448 0.596 0.319 0.488 0.385 0.553 transformer $f_{\theta}$ 0.356 0.542 0.410 0.591 0.312 0.477 0.359 0.537 w/o weighted consensus 0.351 0.526 0.448 0.593 0.361 0.515 0.387 0.545
+Setup.
+To understand the impact of each design choice on the performance and behavior of Flock (Q4), we conduct a series of ablation studies spanning random walks, sequence processor and the consensus protocol, in the entity prediction task in the zero-shot setting.
+Adaptive test-time walks.
+Recall that we employ test-time adaptation of walk counts, which adaptively selects the base walk count $n$ based on the graph size, computed via the harmonic-mean rule shown in Equation˜1 during inference.
+Table˜4 compares this adaptive setting with a fixed setting that uses $128$ base walks per sample for all datasets, matching the pretraining setup ($n_{\text{train}}=128$).
 As expected, the average selected base count $n$ is smaller on both inductive splits and larger on the transductive split, yet the adaptive mechanism improves performance across all settings.
 This is consistent with the intuition that adaptive $n$ scales up walks on larger KGs to improve coverage while allocating fewer walks on smaller KGs to reduce redundant visits; Flock maintains comparable visiting rates and coverage to those seen during pretraining, thereby producing representations closer to the pretraining distribution and resulting in consistent performance gains.
+Non-backtracking walks.
+Flock employs non-backtracking uniform random walk, which has an effect of faster exploration and coverage of distant regions (Alon et al., 2007).
+In Table˜5, we compare this with uniform walk that may backtrack and hence is slower in global exploration.
+While non-backtracking does not alter results much on inductive splits, it significantly improves performance on the transductive split.
+This is consistent with the idea that improving coverage especially benefits performances on large KGs, which Flock achieves via non-backtracking.
+Walk lengths.
+Flock uses random walks of length $\ell=128$, a choice made by finding the lowest $\ell$ reliably visiting target node and relation on various KGs.
+Table˜5 compares this with shorter and longer walks by a factor of two.
+As expected from coverage, shorter walks show degraded results on the transductive split with large KGs.
+Longer walks are overall worse, which is explained by higher learning complexity of the sequence processor that has a small hidden dimension (64) for scalability.
+Flock finds a balance of coverage and learnability, achieving robust results on diverse KGs.
+Diverse starting locations of walks.
+We recall that Flock uses a diversification strategy of starting locations of walks, with $n$ walks from the query node, $n$ walks from random relation, and $n$ walks from random node, adding up to $3n$ walks capturing both local context near query and global information of KG.
+Table˜5 compares this against all $3n$ walks starting at the query node.
+As expected, this causes degradations on all splits, showing the benefit of using both local and global information.
+Sequence processor.
+Flock uses a sequence processor $f_{\theta}$ with bidirectional GRU.
+In Table˜5, we compare this against a transformer $f_{\theta}$ with a similar SwiGLU-RMSNorm architecture (Dubey et al., 2024) and parameter count.
+This alternative does not deliver good results, which is explained by the restrictions on model scales that are enforced to scale to large KGs.
+Flock benefits from reasoning efficiency of GRU in limited parameter regime, gaining good performance and scalability together.
+Consensus protocol.
+Flock uses softmax-weighted averaging to pool sequence processor outputs into state updates for nodes and relations, under the intuition that this can suppress uninformative proposals from the sequence processor better than simple, unweighted averaging.
+Table˜5 provides a comparison, showing that weighted consensus outperforms the unweighted counterpart.
+This verifies our intuition on how the design of the consensus protocol strengthens Flock.
 6 Conclusions
-We introduced Flock, as a knowledge graph foundation model that respects probabilistic node-relation equivariance.
-Flock iteratively samples query-conditioned random walks from the input KG, records encountered nodes and relations via a recording protocol, and relies on a sequence processor and consensus protocol to obtain node and relation representations.
-We empirically evaluate Flock over 54 KGs across different domains for both entity and relation prediction tasks, demonstrating its superior zero-shot and finetuned performances.
+We introduced Flock, a knowledge graph foundation model that respects probabilistic node-relation equivariance.
+Flock iteratively samples query-conditioned random walks, records encountered nodes and relations via a recording protocol, and relies on a sequence processor and consensus protocol to obtain node and relation representations.
+We evaluate Flock on 54 KGs across different domains for both entity and relation prediction, demonstrating superior zero-shot and finetuned performance.
 We further construct a synthetic dataset petals to validate our theoretical findings.
-One limitation is scalability (discussed in Appendix E): ensuring coverage of the sampled random walk in a large KG requires an extensive number of longer walks, which can quickly become computationally infeasible.
-A future direction is to develop approximation strategies (Chamberlain et al., 2023) that reduce the cost of random walk sampling while retaining Flock’s downstream performance.
+One limitation is scalability (Appendix˜E): ensuring coverage of the sampled random walk in large KGs requires many long walks, which can quickly become computationally infeasible.
+A future direction is to develop approximation strategies (Chamberlain et al., 2023; Łącki et al., 2020) that reduce the cost of random walk sampling while retaining Flock’s downstream performance.
+Another avenue for future work is studying the families of approximable functions when the walk lengths are restricted, for example based on connections to subgraph-based reconstructions (Cotta et al., 2021).
 Acknowledgment
-This work is supported in part by the National Research Foundation of Korea grant (No.RS-2024-00351212) and “HPC support” project funded by the Korea government (MSIT and NIPA).
+This work was in part supported by the National Research Foundation of Korea (RS-2024-00351212 and RS-2024-00436165) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) (RS-2024-00509279, RS-2022-II220926, RS-2022-II220959, and RS-2019-II190075) funded by the Korean government (MSIT).
+Computational resources were in part provided by the “HPC support” project funded by MSIT and NIPA.
 MB is supported by EPSRC Turing AI World-Leading Research Fellowship No.
@@
 EP/Y028872/1.
-Appendix A Details of the Petals benchmark
-$b_{0}$$a^{(1)}_{2}$$a^{(1)}_{4}$$a^{(1)}_{1}$$a^{(1)}_{3}$$a^{(1)}_{5}$$a^{(2)}_{1}$$a^{(2)}_{3}$$a^{(2)}_{2}$$a^{(2)}_{4}$$a^{(2)}_{5}$$a^{(3)}_{2}$$a^{(3)}_{4}$$a^{(3)}_{1}$$a^{(3)}_{3}$$a^{(3)}_{5}$$a^{(4)}_{1}$$a^{(4)}_{3}$$a^{(4)}_{2}$$a^{(4)}_{4}$$a^{(4)}_{5}$$b_{1}$$b_{2}$$b_{3}$
-Figure 5: An example of a graph from Petals with $c=4$, $l=2$ and $t=3$, and the associated link prediction instances (dashed).
-The relation types ‘red’, ‘blue’, ‘pink’ and ‘yellow’ are structurally isomorphic, hence become equated in the eyes of the existing KGFMs.
-State-of-the-art knowledge graph foundation models (KGFMs) typically impose relational invariance.
-Formally, given two knowledge graphs $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$, if there exists an isomorphism $(\pi,\phi)$ from $G$ to $H$, then for any $r\in R$, the model enforces identical representations for $r$ and its image $\phi(r)\in R^{\prime}$.
-This design promotes generalization across different graphs, as it aligns analogous relations, but reduces expressivity within a single graph ($G=H$), where relations related by automorphisms are forced to be indistinguishable.
-Concretely, if an automorphism $(\pi,\phi)$ of $G$ maps $r_{1}$ to $r_{2}$, then the model must treat $r_{1}$ and $r_{2}$ as identical during inference.
-While some approaches mitigate this limitation via the labeling trick, assigning distinct embeddings to query-specific nodes and relations, this only isolates the queried relation type and does not resolve the underlying issue in general.
-Motivated by this limitation, we introduce the Petals benchmark.
-Petals comprises 220 graphs, each paired with a link prediction query $(h,r,?)$ and a target set $\{t_{1},t_{2}\}$.
-While $t_{1}$ and $t_{2}$ are non-isomorphic, KGFMs enforcing relational invariance are unable to distinguish them, producing identical predictions.
-We empirically validate this property by evaluating the classification accuracy of marking $t_{1}$ as true and $t_{2}$ as false, reported in Table 1.
-A.1 Structure of the studied KGs
-Knowledge graphs in Petals follow a flower-like structure, parametrized by the number $c$ of ‘petals’, their length $l$ and the length $t$ of the ‘stem’ (see Figure 5 for visualization).
-Vertices.
-Each ‘petal’ is a set $A^{(i)}$ of $2l+1$ vertices $A^{(i)}=\left\{a_{1}^{(i)},a_{2}^{(i)},\dots,a_{2l+1}^{(i)}\right\}$, while the stem $B$ consists of $t+1$ nodes $B=\{b_{0},b_{1},\dots,b_{t}\}$.
-The full set of entities is then:
-$V=B\cup\bigcup_{i=1}^{c}A(i)=\{b_{0},b_{1},\dots,b_{t}\}\cup\left\{a_{j}^{(i)}\mid 1\leq i\leq c,1\leq j\leq 2l+1\right\}$
-We call $b_{0}$ the ‘central’ node, as it is connected to every petal, as described below.
-Edges.
-The nodes of the stem are connected in a consecutive manner by the same relation type $r_{0}$.
-Precisely, for each $i\in{1,\cdots,t}$, there exists an edge $(b_{i-1},r_{0},b_{i})$.
-Each petal $A^{(i)}$ is associated with two edge types $r^{(i)}_{1},r^{(i)}_{2}$, and is connected to the central node $b_{0}$ with links $\left(b_{0},r^{(i)}_{1},a_{1}^{(i)}\right)$ and $\left(b_{0},r^{(i)}_{2},a_{2}^{(i)}\right)$.
-The rest of the petal is connected with edges of type $r_{1}^{(i)}$ only, going from $a^{(i)}_{2j-1}$ to $a^{(i)}_{2j+1}$, and from $a^{(i)}_{2j}$ to $a^{(i)}_{2j+2}$.
-Finally, there are also edges linking $a^{(i)}_{2l-1}$ and $a^{(i)}_{2l}$ to $a^{(i)}_{2l+1}$.
-Therefore, the full set of edges can be characterized as:
-$E=\left(\{(b_{i-1},r_{0},b_{i})\mid 1\leq i\leq t\}\right)$ $\cup\left(\bigcup_{i=1}^{c}\left\{(b_{0},r^{(i)}_{1},a^{(i)}_{1}),(b_{0},r^{(i)}_{2},a^{(i)}_{2})\right\}\right)$
-$\cup\left(\bigcup_{i=1}^{c}\bigcup_{j=1}^{l-1}\left\{\!\left(a^{(i)}_{2j-1},r^{(i)}_{1},a^{(i)}_{2j+1}\right),\!\left(a^{(i)}_{2j},r^{(i)}_{1},a^{(i)}_{2j+2}\right)\right\}\right)$
-$\cup\left(\bigcup_{i=1}^{c}\left\{\!\left(a^{(i)}_{2l-1},r^{(i)}_{1},a^{(i)}_{2l+1}\right),\!\left(a^{(i)}_{2l},r^{(i)}_{1},a^{(i)}_{2l+1}\right)\right\}\right)$
-We select each of the types $r^{(i)}_{1}$ and $r^{(i)}_{2}$ from the set of considered relations $R=\{r_{1},\dots,r_{|R|}\}$ so that any relation-invariant model will equate all petals (i.e.
-so that for each pair of petals, there is an automorphism taking one to another).
-For instance, Figure 5 displays a cyclic pattern, in which $r^{(i)}_{2}=r^{(i+1)}_{1}$.
-Such symmetry causes all petals to be isomorphic, and leads to the inability of KGFMs to distinguish between the relations inside them.
-Link prediction instances.
-Although the petals are isomorphic to each other, given the asymmetry of edge types from $b_{0}$ to $a^{(i)}_{1}$ and $a^{(i)}_{2}$, the nodes within a single petal generally can be distinguished.
-Therefore, for each graph with the structure as described above, we randomly sample one of the stem nodes $b_{s}$, and ask the link prediction query $(b_{s},r_{0},?)$.
-For the target nodes, we randomly select petal index $i$ and distance $j$ from the central node $b_{0}$, and consider the predictions for $a^{(i)}_{2j-1}$ and $a^{(i)}_{2j}$.
-For example, Figure 5 shows the case when $b_{s}=b_{0}$, $i=1$ and $j=1$, where the query is $(b_{0},r_{0},?)$ and we are interested in the scores for $a^{(1)}_{1}$ and $a^{(1)}_{2}$.
-A.2 Parameters and generation
-We construct Petals by manually designing 11 relation-assignment schemes that guarantee isomorphism across all petals.
-For each such selection, which already determines the number $c$ of petals, we generate 20 graphs corresponding to all combinations of $t\in\{1,2,3,4\}$ and $l\in\{1,2,3,4,5\}$.
-Each graph is paired with a link prediction query and two target nodes, sampled as described above.
-This yields $11\cdot 20=220$ instances that constitute the Petals benchmark.
-Appendix B Methodology - details
-In this section, we expand on the descriptions of individual components of Flock summarized in Section 4: the random walk algorithm, the recording protocol, the sequence processor, and the consensus protocol.
-B.1 Uniform random walk
-Let $G=(V,E,R)$ be a knowledge graph, and let $l$ be the length of random walks. For each node $v\in V$, we will denote by $\mathcal{N}(v)$ the set of neighbors of $v$:
+Ethics statement
+This work introduces a probabilistic framework for knowledge graph foundation models, aiming to improve the generalization in zero-shot link prediction.
+Our contributions are methodological and theoretical, with evaluations performed on publicly available benchmarks and a synthetic dataset designed to validate our theoretical results.
+We do not anticipate any direct ethical risks associated with this approach.
+We acknowledge and adhere to the ICLR Code of Ethics.
+Reproducibility statement
+We make every effort to ensure the reproducibility of the experiments in our paper.
+We release a codebase at https://github.com/jw9730/flock with training and evaluation scripts for Flock, including pretraining scripts and checkpoints on FB15k-237, WN18RR, and CoDEx Medium, evaluation over 54 KGs, and the synthetic dataset Petals generator.
+All architectural details needed to reimplement the method, including the random-walk sampler, recording protocol, sequence processor, and consensus protocol, are specified in Appendix˜A, and our theoretical claims are supported with complete proofs in Appendix˜B.
+We additionally include further experimental details in Appendix˜I.
+Appendix A Methodology - details
+In this section, we expand on the descriptions of individual components of Flock summarized in Section˜4: the random walk algorithm, the recording protocol, the sequence processor, and the consensus protocol.
+A.1 Uniform random walk
+Let $G=(V,E,R)$ be a knowledge graph, and let $\ell$ be the length of random walks. For each node $v\in V$, we will denote by $\mathcal{N}(v)$ the set of neighbors of $v$:
 $\mathcal{N}(v)=\{w\in V:\exists r\in R.(v,r,w)\in E\lor(w,r,v)\in E\}$
@@
 $\cup\{(v,r^{-1},w)\in R^{-1}\times\{v\}\times\{w\}:(w,r,v)\in E\}$
-where $R^{-1}$ is the set symbolizing the inverses of relation types in $R$. The uniform random walk with no backtracking $\eta(G,l)$ of length $l$ over $G$, represented as:
-$V_{0}\xrightarrow{R_{1}}V_{1}\xrightarrow{R_{2}}\cdots\xrightarrow{R_{l}}V_{l}$
+where $R^{-1}$ is the set symbolizing the inverses of relation types in $R$. The uniform random walk with no backtracking $\eta(G,\ell)$ of length $\ell$ over $G$, represented as:
+$V_{0}\xrightarrow{R_{1}}V_{1}\xrightarrow{R_{2}}\cdots\xrightarrow{R_{\ell}}V_{\ell}$
 is a second-order Markov process that follows the rules:
@@
 For that problem, each scenario is chosen with probability $\frac{1}{4}$.
-B.2 Recording function
-Given a KG $G=(V,E,R)$, a query $q=(h_{q},r_{q},?)$, a walk $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}$ of length $l$ over $G$, and a set of embeddings $\mathbf{v}$ of nodes $V$ and $\mathbf{r}$ of relations $R$, our recording function $w$ first splits the walk into a sequence of $l+1$ steps:
-$(r_{0},v_{0}),(r_{1},v_{1}),\dots(r_{l},v_{l})$
+A.2 Recording function
+Given a KG $G=(V,E,R)$, a query $q=(h_{q},r_{q},?)$, a walk $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}$ of length $\ell$ over $G$, and a set of embeddings $\mathbf{v}$ of nodes $V$ and $\mathbf{r}$ of relations $R$, our recording function $w$ first splits the walk into a sequence of $\ell+1$ steps:
+$(r_{0},v_{0}),(r_{1},v_{1}),\dots(r_{\ell},v_{\ell})$
 with $r_{0}=r_{\varnothing}$ being a special marker for no relation. Each step $(r_{i},v_{i})$ is transformed into a 7-tuple:
@@
 The output of $w$ for $\bar{\eta}$ given $G,q,\mathbf{v},\mathbf{r}$ is then:
-$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{l})$
-B.3 Sequence processor
+$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{\ell})$
+A.3 Sequence processor
 Once the sampled walks are anonymized by the recording protocol $w$, the output for each walk $\bar{\eta}_{i}$:
-$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{l})$
+$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{\ell})$
 is passed through the sequence processor $f_{\theta}$, parametrized by the following modules:
-  • $\mathbf{A}_{v},\mathbf{A}_{r}\in\mathbb{R}^{(l+1)\times d}$: embedding tables for anonymized vertices and relations, respectively,
+  • $\mathbf{A}_{v},\mathbf{A}_{r}\in\mathbb{R}^{(\ell+1)\times d}$: embedding tables for anonymized vertices and relations, respectively,
   • $\mathbf{D}\in\mathbb{R}^{2\times d}$: look-up table for the direction embedding,
@@
 These are then passed to the GRU cell $\mathbf{\Omega}$, which fuses the features across the whole walk and produces multi-head embeddings of vertices and relations, as well as the associated weights:
-$\left(\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i},\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\right)=\mathbf{\Omega}([\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{l}])$
-where $\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i}\in\mathbb{R}^{(l+1)\times h\times d_{h}}$ and $\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\in\mathbb{R}^{(l+1)\times h}$.
+$\left(\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i},\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\right)=\mathbf{\Omega}([\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{\ell}])$
+where $\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i}\in\mathbb{R}^{(\ell+1)\times h\times d_{h}}$ and $\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\in\mathbb{R}^{(\ell+1)\times h}$.
 Stacking all $N$ of them gives us the final output of the sequence processor.
-B.4 Consensus protocol
+A.4 Consensus protocol
 Given walks $\bar{\eta}_{1:N}$ over $G=(V,E,R)$ and the outputs $\mathbf{s}^{(V)},\mathbf{s}^{(R)},\mathbf{a}^{(V)},\mathbf{a}^{(R)}$ of the sequence processor, the consensus protocol $c$ aggregates the signal for each node by evaluating a weighted sum over the appearances of this node across the walks.
@@
 $\Delta\mathbf{r}(r)$ $=\Delta\mathbf{r}^{\prime}(\phi(r))\quad$ $\forall r\in R$
-Appendix C Proofs
-C.1 Expressivity
+Appendix B Proofs
+B.1 Expressivity
 The main proposition of this section formalizes the fact that Flock can approximate any link-invariant function over fixed-size knowledge graphs in probability.
-Intuitively, when the length of the sampled walks $l$ becomes higher, the probability of a single walk witnessing all the edges grows to 1.
+Intuitively, when the length of the sampled walks $\ell$ becomes higher, the probability of a single walk witnessing all the edges grows to 1.
 Once a walk visits all the edges, a sufficiently powerful sequence processor can derive the whole graph structure from its anonymized representation, recreating the graph in its entirety, up to isomorphism.
@@
 We start by showing that the edge cover time $C_{E}(\cdot)$ of graphs in $K_{n,m}$ is bounded:
-Lemma C.1.
-Let $G\in\mathbb{K}_{n,m}$ for some $n,m$. The edge cover time $C_{E}(G)$ of $G$, using the algorithm from Section B.1, is finite.
+Lemma B.1.
+Let $G\in\mathbb{K}_{n,m}$ for some $n,m$. The edge cover time $C_{E}(G)$ of $G$, using the algorithm from Section˜A.1, is finite.
 Proof.
 Let $G=(V,E,R)\in\mathbb{K}_{n,m}$ be a graph.
-For any edge $e\in E$ and any vertex $v\in V$, let $H_{v}(e)$ denote the expected number of steps of the random walk algorithm $\eta$ described in Section B.1.
+For any edge $e\in E$ and any vertex $v\in V$, let $H_{v}(e)$ denote the expected number of steps of the random walk algorithm $\eta$ described in Section˜A.1.
 Then, the edge cover time $C_{E}(G)$ of $G$ with $\eta$, i.e.
@@
 By the definition of the used random walk algorithm, which only looks at the previously taken edge, we can deduce that the events $A_{i}$ are all mutually independent.
-Moreover, let $v_{i(d+2)}=u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\dots\xrightarrow{s_{l}}u_{l}\in\{h_{e},t_{e}\}$ be a shortest path from $v_{i(d+2)}$ to one of $h_{e},t_{e}$.
+Moreover, let $v_{i(d+2)}=u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\dots\xrightarrow{s_{\ell}}u_{\ell}\in\{h_{e},t_{e}\}$ be a shortest path from $v_{i(d+2)}$ to one of $h_{e},t_{e}$.
 Note that by minimality, there cannot be any backtracking while following this path.
 Therefore, the probability of the next visited node is dependent only on the value of the previous one, and we can bound the probability $P(A_{i})$ of $A_{i}$ from below by:
-${\mathbb{P}}(A_{i})\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{l-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
+${\mathbb{P}}(A_{i})\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{\ell-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
 The first term on the right hand side is the probability of selecting $e$ while being at $h_{e}$ or $t_{e}$, which is the probability of first selecting the other endpoint (out of at most $\Delta$ neighbors) and then picking $e$ over other edges between $h_{e}$ and $t_{e}$ (of which there is at most $\rho$).
@@
 Combining these observations, we can derive a bound for ${\mathbb{P}}(A_{i})$ in terms of $\Delta,\rho$ and $d$:
-${\mathbb{P}}(A_{i})$ $\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{l-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
-$\geq\frac{1}{\Delta\cdot\rho}\cdot\prod_{j=0}^{l-1}\frac{1}{\Delta}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{l}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{d}\quad=\quad\frac{1}{\rho\Delta^{d+1}}$
+${\mathbb{P}}(A_{i})$ $\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{\ell-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
+$\geq\frac{1}{\Delta\cdot\rho}\cdot\prod_{j=0}^{\ell-1}\frac{1}{\Delta}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{\ell}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{d}\quad=\quad\frac{1}{\rho\Delta^{d+1}}$
 Finally, note that if $A_{i}$ is true, then the first index $t$ such that $v_{t-1}\xrightarrow{r_{t}}v_{t}$ traverses $e$ is at most $(i+1)(d+2)$.
@@
 Since $\rho\leq m,d+2\leq n$ and $\Delta\leq n$, we have $H_{v}(e)\leq m(n+2)n^{n}$, which completes the proof. ∎
-Remark C.2.
-The bound obtained in the proof of Lemma C.1 is very crude.
+Remark B.2.
+The bound obtained in the proof of Lemma B.1 is very crude.
 In fact, we could transform the given knowledge graph into a simple graph (undirected, with no multi-edges) by substituting each edge $u\xrightarrow{r}v$ with two undirected edges $u\leftrightarrow v_{(u,r,v)}\leftrightarrow v$.
@@
 Let us now prove a fact about the number of distinct, up to isomorphism, graphs in $\mathbb{K}_{n,m}$.
-Lemma C.3.
+Lemma B.3.
 For any $n,m$, the number of isomorphism classes in $\mathbb{K}_{n,m}$ is finite.
@@
 ∎
-Lemma C.4.
-For each pair $(n,m)$, there exists a number $C_{n,m}$ such that the edge cover time, using the algorithm from Section B.1, of any knowledge graph in $\mathbb{K}_{n,m}$ is at most $C_{n,m}$.
+Lemma B.4.
+For each pair $(n,m)$, there exists a number $C_{n,m}$ such that the edge cover time, using the algorithm from Section˜A.1, of any knowledge graph in $\mathbb{K}_{n,m}$ is at most $C_{n,m}$.
 Proof.
-The result follows from Lemmas C.1 and C.3.
+The result follows from Lemmas B.1 and B.3.
 As two isomorphic graphs have identical cover time, we can set $C_{n,m}$ to be the maximum of cover times of representatives of all isomorphic classes, which, by finiteness of both, is well-defined.
 ∎
-Lemma C.5.
+Lemma B.5.
 Let $G\in\mathbb{K}_{n,m}$ be a graph, $q=(h_{q},r_{q},?)$ be a link query over $G$, and $\bar{\eta}$ be a walk over $G$.
-If $\bar{\eta}$ traverses all edges of $G$, then using only the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording function $w$ detailed in Section B.2, we can construct a graph-query pair $(H,q^{\prime})$ isomorphic to $(G,q)$.
+If $\bar{\eta}$ traverses all edges of $G$, then using only the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording function $w$ detailed in Section˜A.2, we can construct a graph-query pair $(H,q^{\prime})$ isomorphic to $(G,q)$.
 Proof.
-Suppose that $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}$ visits all edges of $G=(V,E,R)$ and let $l$ be its length.
-Recall the anonymization functions $\operatorname{id}_{V}(\cdot;\bar{\eta})$ and $\operatorname{id}_{R}(\cdot;\bar{\eta})$ as defined in Section B.2.
-The output $w(\bar{\eta};G,q,\cdot,\cdot)$ (the embedding functions provided as the last two arguments are irrelevant) is a sequence of tuples $S_{0},S_{1},\dots,S_{l}$ with each $S_{i}$ equal to:
+Suppose that $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}$ visits all edges of $G=(V,E,R)$ and let $\ell$ be its length.
+Recall the anonymization functions $\operatorname{id}_{V}(\cdot;\bar{\eta})$ and $\operatorname{id}_{R}(\cdot;\bar{\eta})$ as defined in Section˜A.2.
+The output $w(\bar{\eta};G,q,\cdot,\cdot)$ (the embedding functions provided as the last two arguments are irrelevant) is a sequence of tuples $S_{0},S_{1},\dots,S_{\ell}$ with each $S_{i}$ equal to:
 $S_{i}=\left(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{i};\bar{\eta}),\operatorname{dir}_{i},\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}},\cdot,\cdot\right)$
@@
 and a query $q^{\prime}=(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{j},\bar{\eta}),?)$ for $i,j$ such that $\delta_{v_{i}=h_{q}}=1$ and $\delta_{r_{j}=r_{q}}=1$.
-Then by the definition of $w$ (Section B.2), it is straightforward to check that the pair $(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$ defines an isomorphism from $(G,q)$ to $(H,q^{\prime})$.
+Then by the definition of $w$ (Section˜A.2), it is straightforward to check that the pair $(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$ defines an isomorphism from $(G,q)$ to $(H,q^{\prime})$.
 Indeed, both these functions are injective by construction, and as $\bar{\eta}$ witnesses all nodes and relations, they are well-defined bijections.
@@
 We are now ready to prove the main result regarding the universality of Flock as an approximation of link invariant functions.
-The outline of the proof is as follows: 1) Using the upper-bound on the edge cover time of graphs in $\mathbb{K}_{n,m}$ derived in Lemma C.4, we can bound the probability of sampling a walk that visits all edges, 2) Once such a walk is sampled, we can recover the graph and query, up to isomorphism, from its anonymized form (Lemma C.5), 3) Lastly, we can return the value of the approximated function for the derived isomorphic instance.
+The outline of the proof is as follows: 1) Using the upper-bound on the edge cover time of graphs in $\mathbb{K}_{n,m}$ derived in Lemma B.4, we can bound the probability of sampling a walk that visits all edges, 2) Once such a walk is sampled, we can recover the graph and query, up to isomorphism, from its anonymized form (Lemma B.5), 3) Lastly, we can return the value of the approximated function for the derived isomorphic instance.
 Since the approximated function is link invariant, if the reconstructed graph matches the original one, we return precisely the correct value.
 Proposition 4.1.
-With a powerful enough sequence processor $f_{\theta}$, the Flock framework described in Section 4 is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
+With a powerful enough sequence processor $f_{\theta}$, the Flock framework described in Section˜4 is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
 Proof.
@@
 ${\mathbb{P}}(|\varphi(G)((h,r,t))-X_{\theta}(G,(h,r,?))(t)|<\epsilon)>1-\delta$ (4)
-For simplicity, let us consider a situation where only a single walk $\bar{\eta}$ of length $l$ is sampled by the model (otherwise, omit additional walks).
+For simplicity, let us consider a situation where only a single walk $\bar{\eta}$ of length $\ell$ is sampled by the model (otherwise, omit additional walks).
 We will also restrict the argument to a single refinement case – the result can be extended to multiple refinement steps by returning $\Delta\mathbf{v},\Delta\mathbf{r}=0$ during all additional iterations.
-Consider a sequence processor $f_{\theta}$ that given the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording protocol, creates a graph-query pair $(H,q^{\prime})$ with $q^{\prime}=(h_{q^{\prime}},r_{q^{\prime}},?)$ using the strategy described in the proof of Lemma C.5, and returns a vector $\mathbf{h}\in\mathbb{R}^{l+1}$ whose $i^{\text{th}}$ entry is equal $\mathbf{h}_{i}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(v_{i};\bar{\eta}))$ where $v_{i}$ is the $i^{\text{th}}$ node visited by $\bar{\eta}$.
+Consider a sequence processor $f_{\theta}$ that given the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording protocol, creates a graph-query pair $(H,q^{\prime})$ with $q^{\prime}=(h_{q^{\prime}},r_{q^{\prime}},?)$ using the strategy described in the proof of Lemma B.5, and returns a vector $\mathbf{h}\in\mathbb{R}^{l+1}$ whose $i^{\text{th}}$ entry is equal $\mathbf{h}_{i}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(v_{i};\bar{\eta}))$ where $v_{i}$ is the $i^{\text{th}}$ node visited by $\bar{\eta}$.
 The consensus protocol $c$, provided $t$ was visited by $\bar{\eta}$, can then identify $t$ as one of the $v_{j}$ and pull the corresponding embedding $\mathbf{h}_{j}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$, returning it as the output $\mathbf{v}(t)=\mathbf{h}_{j}$ (note that no matter which specific value of $j$ is chosen, this value will be the same).
@@
 $\varphi(G)((h,r,t))=X_{\theta}(G,(h,r,?))(t)$
-By Lemma C.5, in such case, the reconstructed pair $(H,q^{\prime})$ is isomorphic to $(G,q)$ by the isomorphism $id=(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$.
+By Lemma B.5, in such case, the reconstructed pair $(H,q^{\prime})$ is isomorphic to $(G,q)$ by the isomorphism $id=(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$.
 Since $\varphi$ is link invariant, we can write:
@@
 $\varphi(G)((h,r,t))=X_{\theta}(G,(h,r,?))(t)$
-Hence, to show (9), it suffices to prove that we can uniformly choose the length $l$ of the random walk so that the probability of $\bar{\eta}$ covering all the edges is greater than $1-\delta$.
+Hence, to show (4), it suffices to prove that we can uniformly choose the length $\ell$ of the random walk so that the probability of $\bar{\eta}$ covering all the edges is greater than $1-\delta$.
 By Markov’s inequality:
@@
 […]
-$=\frac{C_{E}(G)}{l}$
-But by Lemma C.4, $C_{E}(G)\leq C_{n,m}$ for some constant $C_{n,m}$. Hence, taking $l>\frac{C_{n,m}}{\delta}$, we get:
-${\mathbb{P}}(\bar{\eta}\text{ does not cover all edges})\leq\frac{C_{E}(G)}{l}\leq\frac{C_{n,m}}{l}<\delta$
-This means that for such a choice of $l$:
+$=\frac{C_{E}(G)}{\ell}$
+But by Lemma B.4, $C_{E}(G)\leq C_{n,m}$ for some constant $C_{n,m}$. Hence, taking $\ell>\frac{C_{n,m}}{\delta}$, we get:
+${\mathbb{P}}(\bar{\eta}\text{ does not cover all edges})\leq\frac{C_{E}(G)}{\ell}\leq\frac{C_{n,m}}{\ell}<\delta$
+This means that for such a choice of $\ell$:
 ${\mathbb{P}}(\bar{\eta}\text{ witnesses all edges of }G)>1-\delta$
-which leads to the conclusion that for $l>\frac{C_{n,m}}{\delta}$, the proposed Flock framework satisfies:
+which leads to the conclusion that for $\ell>\frac{C_{n,m}}{\delta}$, the proposed Flock framework satisfies:
 ${\mathbb{P}}(|\varphi(G)((h,r,t))-X_{\theta}(G,(h,r,?))(t)|<\epsilon)>1-\delta$
 for any choice of $G=(V,E,R)\in\mathbb{K}_{n,m}$ and $(h,r,t)\in V\times R\times V$. ∎
-C.2 Invariance
+B.2 Invariance
 First, let us recall the definition of invariance for the context of knowledge graphs and the associated notion of invariance in probability.
@@
 $\mu(q)=\mu((h,r,?))=(\pi(h),\phi(r),?)$
-Similarly, let $\eta=v_{0}\xrightarrow{r_{1}}\dots\xrightarrow{r_{l}}v_{l}$ be a walk of length $l$ in $G$. The view of $\eta$ with $\mu$ is defined as:
-$\mu\left(v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\right)=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{l})}\pi(v_{l})$
+Similarly, let $\eta=v_{0}\xrightarrow{r_{1}}\dots\xrightarrow{r_{\ell}}v_{\ell}$ be a walk of length $\ell$ in $G$. The view of $\eta$ with $\mu$ is defined as:
+$\mu\left(v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}\right)=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{\ell})}\pi(v_{\ell})$
 Let $f$ be a function taking inputs drawn from KGs.
@@
 $\forall G\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$\mu$}}}}}}{{\simeq}}H,q,\mathbf{v}\in\mathbb{R}^{d}:\quad\varphi(G,q,\mathbf{v})=\varphi(\mu(G),\mu(q),\mathbf{v})$
-Following the definition of invariance in probability, provided in Section 3, we extend all the definitions above to the stochastic case, replacing equality (==) with equality in distribution ($\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}$).
-We can now prove the main propositions stated in Section 4.2. Let’s begin with the more general:
+Following the definition of invariance in probability, provided in Section˜3, we extend all the definitions above to the stochastic case, replacing equality (==) with equality in distribution ($\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}$).
+We can now prove the main propositions stated in Section˜4.2. Let’s begin with the more general:
 Proposition 4.2.
@@
 which completes the induction step, and hence the proof. ∎
-We can use the conclusion from Proposition 4.2 to prove the probabilistic invariance of the architecture proposed in Section 4.
+We can use the conclusion from Proposition 4.2 to prove the probabilistic invariance of the architecture proposed in Section˜4.
 To be able to apply it, we first need to verify the invariance of all used components, which we formalize in the following lemmas.
-Lemma C.6.
-The choice of the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ of the uniform random walk algorithm described in Section B.1 is invariant.
+Lemma B.6.
+The choice of the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ of the uniform random walk algorithm described in Section˜A.1 is invariant.
 Proof.
@@
 Consider a link prediction query $q=(h,r,?)$ over $G$, and its identification $q^{\prime}=\mu(q)=(\pi(h),\phi(r),?)$.
-The goal is to show that when using $\eta$ described in Section B.1 for $(G,q)$ and $(H,q^{\prime})$, the first steps:
+The goal is to show that when using $\eta$ described in Section˜A.1 for $(G,q)$ and $(H,q^{\prime})$, the first steps:
 $V_{0}\xrightarrow{R_{1}}V_{1}\qquad\text{and}\qquad U_{0}\xrightarrow{S_{1}}U_{1}$
@@
 Either way, we find that the selection process of the first step of $\eta$ over $G$ translates naturally via $\mu$ to the choice of the first step over $H$, proving the desired statement. ∎
-Lemma C.7.
+Lemma B.7.
 Suppose that the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ is chosen in an invariant manner. Then, the uniform random walk with no backtracking algorithm $\eta$ is invariant in probability.
 Proof.
-Let $G=(V,E,R)$ be a knowledge graph, and let $l$ be the length of random walks. Let $H$ be a KG isomorphic to $G$ via the isomorphism $\mu=(\pi,\phi)$. We aim to show that:
-$\mu(\eta(G,l))=\pi(V_{0})\xrightarrow{\phi(R_{1})}\pi(V_{1})\xrightarrow{\phi(R_{2})}\dots\xrightarrow{\phi(R_{l})}\pi(V_{l})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}U_{0}\xrightarrow{S_{1}}U_{1}\xrightarrow{S_{2}}\dots\xrightarrow{S_{l}}U_{l}=\eta(H,l)$
-Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\in\mathcal{W}(G)$ be a walk of length $l$ over $G$.
+Let $G=(V,E,R)$ be a knowledge graph, and let $\ell$ be the length of random walks. Let $H$ be a KG isomorphic to $G$ via the isomorphism $\mu=(\pi,\phi)$. We aim to show that:
+$\mu(\eta(G,\ell))=\pi(V_{0})\xrightarrow{\phi(R_{1})}\pi(V_{1})\xrightarrow{\phi(R_{2})}\dots\xrightarrow{\phi(R_{\ell})}\pi(V_{\ell})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}U_{0}\xrightarrow{S_{1}}U_{1}\xrightarrow{S_{2}}\dots\xrightarrow{S_{\ell}}U_{\ell}=\eta(H,\ell)$
+Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}\in\mathcal{W}(G)$ be a walk of length $\ell$ over $G$.
 It suffices to show that the probability of sampling $\bar{\eta}$ from $G$ is identical to the probability of sampling $\mu(\bar{\eta})$ from $H$:
-${\mathbb{P}}(\eta(G,l)=\bar{\eta})={\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))$
-To see this, let us expand the definitions of ${\mathbb{P}}(\eta(G,l)=\bar{\eta})$:
-${\mathbb{P}}(\eta(G,l)=\bar{\eta})=$ ${\mathbb{P}}(V_{0}=v_{0})$ (15)
+${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})={\mathbb{P}}(\eta(H,\ell)=\mu(\bar{\eta}))$
+To see this, let us expand the definitions of ${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})$:
+${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})=$ ${\mathbb{P}}(V_{0}=v_{0})$ (10)
 $\cdot{\mathbb{P}}(V_{1}=v_{1}\mid V_{0}=v_{0})$
-$\cdot\prod^{l-2}_{i=0}{\mathbb{P}}(V_{i+2}=v_{i+2}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
-$\cdot\prod^{l-1}_{i=0}{\mathbb{P}}(R_{i+1}=r_{i+1}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
-and $P(\eta(H,l)=\mu(\bar{\eta}))$:
-${\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))=$ ${\mathbb{P}}(U_{0}=\pi(v_{0}))$ (16)
+$\cdot\prod^{\ell-2}_{i=0}{\mathbb{P}}(V_{i+2}=v_{i+2}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
+$\cdot\prod^{\ell-1}_{i=0}{\mathbb{P}}(R_{i+1}=r_{i+1}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
+and $P(\eta(H,\ell)=\mu(\bar{\eta}))$:
+${\mathbb{P}}(\eta(H,\ell)=\mu(\bar{\eta}))=$ ${\mathbb{P}}(U_{0}=\pi(v_{0}))$ (11)
 $\cdot{\mathbb{P}}(U_{1}=\pi(v_{1})\mid U_{0}=\pi(v_{0}))$
-$\cdot\prod^{l-2}_{i=0}{\mathbb{P}}(U_{i+2}=\pi(v_{i+2})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
-$\cdot\prod^{l-1}_{i=0}{\mathbb{P}}(S_{i+1}=\phi(r_{i+1})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
+$\cdot\prod^{\ell-2}_{i=0}{\mathbb{P}}(U_{i+2}=\pi(v_{i+2})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
+$\cdot\prod^{\ell-1}_{i=0}{\mathbb{P}}(S_{i+1}=\phi(r_{i+1})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
 Given that the graph isomorphism preserves the number of neighbors for each node and is a bijection, we can easily verify using the definitions from (3) that the following indeed hold:
@@
 ${\mathbb{P}}(R_{j+1}=r_{j+1}\mid V_{j+1}=v_{j+1},V_{j}=v_{j})$ $={\mathbb{P}}(S_{j+1}=\phi(r_{j+1})\mid U_{j+1}=\pi(v_{j+1}),U_{j}=\pi(v_{j}))$
-for all $i\in\{0,1,\dots,l-2\},j\in\{1,\dots,l-1\}$. Moreover, by the assumption that the first step $V_{0}\xrightarrow{R_{1}}V_{1}$ is invariant, we have:
+for all $i\in\{0,1,\dots,\ell-2\},j\in\{1,\dots,\ell-1\}$. Moreover, by the assumption that the first step $V_{0}\xrightarrow{R_{1}}V_{1}$ is invariant, we have:
 ${\mathbb{P}}((V_{0},R_{1},V_{1})=(v_{0},r_{1},v_{1}))={\mathbb{P}}((U_{0},S_{1},U_{1})=(\pi(v_{0}),\phi(r_{1}),\pi(v_{1})))$ (13)
@@
 $={\mathbb{P}}(S_{1}=\phi(r_{1})\mid U_{0}=\pi(v_{0}),U_{1}=\pi(v_{1}))\cdot{\mathbb{P}}(U_{1}=\pi(v_{0})\mid U_{0}=\pi(v_{0}))\cdot{\mathbb{P}}(U_{0}=\pi(v_{0}))$
-Substituting these equalities into (18) and multiplying both sides by the equalities from (17) for all choices of $i\in\{0,1,\dots,l-2\},j\in\{1,\dots,l-1\}$, we get precisely the equality of the right sides of equations (15) and (16).
+Substituting these equalities into (13) and multiplying both sides by the equalities from (12) for all choices of $i\in\{0,1,\dots,\ell-2\},j\in\{1,\dots,\ell-1\}$, we get precisely the equality of the right sides of Equation˜10 and Equation˜11.
 Hence,
-${\mathbb{P}}(\eta(G,l)=\bar{\eta})={\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))$
-and we can conclude that $\mu(\eta(G,l))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\eta(H,l)$, and the algorithm $\eta$ is invariant in probability. ∎
-Corollary C.8.
-The random walk algorithm presented in Section B.1 is invariant in probability.
-Lemma C.9.
-The recording protocol $w$, as described in Section B.2, is invariant, provided that the embedding functions $\mathbf{v}$ and $\mathbf{r}$ are invariant.
+${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})={\mathbb{P}}(\eta(H,\ell)=\mu(\bar{\eta}))$
+and we can conclude that $\mu(\eta(G,\ell))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\eta(H,\ell)$, and the algorithm $\eta$ is invariant in probability. ∎
+Corollary B.8.
+The random walk algorithm presented in Section˜A.1 is invariant in probability.
+Lemma B.9.
+The recording protocol $w$, as described in Section˜A.2, is invariant, provided that the embedding functions $\mathbf{v}$ and $\mathbf{r}$ are invariant.
 Proof.
@@
 Let $q=(h_{q},r_{q},?)$ be a link prediction query over $G$, and $\mu(q)=(\pi(h_{q}),\phi(r_{q}),?)$ be its identification in $H$.
-Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\in\mathcal{W}(G)$ be a walk over $G$, and $\bar{\eta}^{\prime}=\mu(\bar{\eta})=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{l})}\pi(v_{l})$ be the analogous walk over $H$.
-To prove that the recording protocol $w$ outlined in Section B.2 is invariant, it suffices to show that the encoding of each step:
+Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}\in\mathcal{W}(G)$ be a walk over $G$, and $\bar{\eta}^{\prime}=\mu(\bar{\eta})=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{\ell})}\pi(v_{\ell})$ be the analogous walk over $H$.
+To prove that the recording protocol $w$ outlined in Section˜A.2 is invariant, it suffices to show that the encoding of each step:
 $S_{i}=\left(\operatorname{id}_{V}(v_{i};\bar{\eta}),\operatorname{id}_{R}(r_{i};\bar{\eta}),\operatorname{dir}_{i},\delta_{v_{i}=h_{q}},\delta_{r_{i}=r_{q}},\mathbf{v}(v_{i}),\mathbf{r}(r_{i})\right)$
@@
 Combining all these observations, we can conclude that $w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=w(\mu(\bar{\eta});H,\mu(q),\mathbf{v},\mathbf{r})$ and $w$ is indeed invariant. ∎
-Lemma C.10.
-The consensus protocol $c$, as described in Section B.4, is invariant.
+Lemma B.10.
+The consensus protocol $c$, as described in Section˜A.4, is invariant.
 Proof.
@@
 Proposition 4.3.
-Flock with components as described in Section 4 is invariant in probability.
+Flock with components as described in Section˜4 is invariant in probability.
 Proof.
-The result follows naturally from aggregating the results of Corollary C.8 and Lemmas C.9 and C.10, followed by applying Proposition 4.2. ∎
+The result follows naturally from aggregating the results of Corollary B.8 and Lemmas B.9 and B.10, followed by applying Proposition 4.2. ∎
+Appendix C Details of the Petals benchmark
+$b_{0}$$a^{(1)}_{2}$$a^{(1)}_{4}$$a^{(1)}_{1}$$a^{(1)}_{3}$$a^{(1)}_{5}$$a^{(2)}_{1}$$a^{(2)}_{3}$$a^{(2)}_{2}$$a^{(2)}_{4}$$a^{(2)}_{5}$$a^{(3)}_{2}$$a^{(3)}_{4}$$a^{(3)}_{1}$$a^{(3)}_{3}$$a^{(3)}_{5}$$a^{(4)}_{1}$$a^{(4)}_{3}$$a^{(4)}_{2}$$a^{(4)}_{4}$$a^{(4)}_{5}$$b_{1}$$b_{2}$$b_{3}$
+Figure 5: An example of a graph from Petals with $c=4$, $l=2$ and $t=3$, and the associated link prediction instances (dashed).
+The relation types ‘red’, ‘blue’, ‘pink’ and ‘yellow’ are structurally isomorphic, hence become equated in the eyes of the existing KGFMs.
+State-of-the-art knowledge graph foundation models (KGFMs) typically impose relational invariance.
+Formally, given two knowledge graphs $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$, if there exists an isomorphism $(\pi,\phi)$ from $G$ to $H$, then for any $r\in R$, the model enforces identical representations for $r$ and its image $\phi(r)\in R^{\prime}$.
+This design promotes generalization across different graphs, as it aligns analogous relations, but reduces expressivity within a single graph ($G=H$), where relations related by automorphisms are forced to be indistinguishable.
+Concretely, if an automorphism $(\pi,\phi)$ of $G$ maps $r_{1}$ to $r_{2}$, then the model must treat $r_{1}$ and $r_{2}$ as identical during inference.
+While some approaches mitigate this limitation via the labeling trick, assigning distinct embeddings to query-specific nodes and relations, this only isolates the queried relation type and does not resolve the underlying issue in general.
+Motivated by this limitation, we introduce the Petals benchmark.
+Petals comprises 220 graphs, each paired with a link prediction query $(h,r,?)$ and a target set $\{t_{1},t_{2}\}$.
+While $t_{1}$ and $t_{2}$ are non-isomorphic, KGFMs enforcing relational invariance are unable to distinguish them, producing identical predictions.
+We empirically validate this property by evaluating the classification accuracy of marking $t_{1}$ as true and $t_{2}$ as false, reported in Table˜1.
+C.1 Structure of the studied KGs
+Knowledge graphs in Petals follow a flower-like structure, parametrized by the number $c$ of ‘petals’, their length $l$ and the length $t$ of the ‘stem’ (see Figure˜5 for visualization).
+Vertices.
+Each ‘petal’ is a set $A^{(i)}$ of $2l+1$ vertices $A^{(i)}=\left\{a_{1}^{(i)},a_{2}^{(i)},\dots,a_{2l+1}^{(i)}\right\}$, while the stem $B$ consists of $t+1$ nodes $B=\{b_{0},b_{1},\dots,b_{t}\}$.
+The full set of entities is then:
+$V=B\cup\bigcup_{i=1}^{c}A(i)=\{b_{0},b_{1},\dots,b_{t}\}\cup\left\{a_{j}^{(i)}\mid 1\leq i\leq c,1\leq j\leq 2l+1\right\}$
+We call $b_{0}$ the ‘central’ node, as it is connected to every petal, as described below.
+Edges.
+The nodes of the stem are connected in a consecutive manner by the same relation type $r_{0}$.
+Precisely, for each $i\in{1,\cdots,t}$, there exists an edge $(b_{i-1},r_{0},b_{i})$.
+Each petal $A^{(i)}$ is associated with two edge types $r^{(i)}_{1},r^{(i)}_{2}$, and is connected to the central node $b_{0}$ with links $\left(b_{0},r^{(i)}_{1},a_{1}^{(i)}\right)$ and $\left(b_{0},r^{(i)}_{2},a_{2}^{(i)}\right)$.
+The rest of the petal is connected with edges of type $r_{1}^{(i)}$ only, going from $a^{(i)}_{2j-1}$ to $a^{(i)}_{2j+1}$, and from $a^{(i)}_{2j}$ to $a^{(i)}_{2j+2}$.
+Finally, there are also edges linking $a^{(i)}_{2\ell-1}$ and $a^{(i)}_{2l}$ to $a^{(i)}_{2l+1}$.
+Therefore, the full set of edges can be characterized as:
+$E=\left(\{(b_{i-1},r_{0},b_{i})\mid 1\leq i\leq t\}\right)$ $\cup\left(\bigcup_{i=1}^{c}\left\{(b_{0},r^{(i)}_{1},a^{(i)}_{1}),(b_{0},r^{(i)}_{2},a^{(i)}_{2})\right\}\right)$
+$\cup\left(\bigcup_{i=1}^{c}\bigcup_{j=1}^{\ell-1}\left\{\!\left(a^{(i)}_{2j-1},r^{(i)}_{1},a^{(i)}_{2j+1}\right),\!\left(a^{(i)}_{2j},r^{(i)}_{1},a^{(i)}_{2j+2}\right)\right\}\right)$
+$\cup\left(\bigcup_{i=1}^{c}\left\{\!\left(a^{(i)}_{2\ell-1},r^{(i)}_{1},a^{(i)}_{2l+1}\right),\!\left(a^{(i)}_{2l},r^{(i)}_{1},a^{(i)}_{2l+1}\right)\right\}\right)$
+We select each of the types $r^{(i)}_{1}$ and $r^{(i)}_{2}$ from the set of considered relations $R=\{r_{1},\dots,r_{|R|}\}$ so that any relation-invariant model will equate all petals (i.e.
+so that for each pair of petals, there is an automorphism taking one to another).
+For instance, Figure˜5 displays a cyclic pattern, in which $r^{(i)}_{2}=r^{(i+1)}_{1}$.
+Such symmetry causes all petals to be isomorphic, and leads to the inability of KGFMs to distinguish between the relations inside them.
+Link prediction instances.
+Although the petals are isomorphic to each other, given the asymmetry of edge types from $b_{0}$ to $a^{(i)}_{1}$ and $a^{(i)}_{2}$, the nodes within a single petal generally can be distinguished.
+Therefore, for each graph with the structure as described above, we randomly sample one of the stem nodes $b_{s}$, and ask the link prediction query $(b_{s},r_{0},?)$.
+For the target nodes, we randomly select petal index $i$ and distance $j$ from the central node $b_{0}$, and consider the predictions for $a^{(i)}_{2j-1}$ and $a^{(i)}_{2j}$.
+For example, Figure˜5 shows the case when $b_{s}=b_{0}$, $i=1$ and $j=1$, where the query is $(b_{0},r_{0},?)$ and we are interested in the scores for $a^{(1)}_{1}$ and $a^{(1)}_{2}$.
+C.2 Parameters and generation
+We construct Petals by manually designing 11 relation-assignment schemes that guarantee isomorphism across all petals.
+For each such selection, which already determines the number $c$ of petals, we generate 20 graphs corresponding to all combinations of $t\in\{1,2,3,4\}$ and $l\in\{1,2,3,4,5\}$.
+Each graph is paired with a link prediction query and two target nodes, sampled as described above.
+This yields $11\cdot 20=220$ instances that constitute the Petals benchmark.
 Appendix D Computational Complexity
@@
 In total, the time complexity is $O\big(PIn\ell Ld^{2}\big)$, which scales linearly with the number of (base) walks $n$, the length of walks $\ell$, and the number of ensembled predictions $P$.
-We empirically verified this in Appendix E.
+We empirically verified this in Appendix˜E.
 Compared with message-passing KGFMs like Ultra and Trix, Flock’s complexity is independent of the graph size and average degree; empirically, however, using more walks (increasing $n$) and longer walks (increasing $\ell$) improves coverage and yields more fine-grained representation.
@@
 Flock 801,969 1.30 27.89
-Table 6: Inference scalability on a single NVIDIA RTX A6000 (48 GB); batch size =8=8. Left columns specify base walks $n$ and ensembled passes $P$. Dashes indicate not applicable.
+Table 7: Inference scalability on a single NVIDIA RTX A6000 (48 GB) with batch size =8=8. Left columns specify base walks $n$ and ensembled passes $P$. Dashes indicate not applicable.
 Model # Base Walks $n$ Ensemble $P$ Time /batch (s) GPU memory (GB)
@@
 128 8 10.05 43.68
-To investigate the scalability of the proposed method Flock, we report the training and inference time per batch and peak GPU memory for Ultra, Trix, and variants of Flock on a single RTX A6000 (48 GB) in Tables 5 and 6.
+To investigate the scalability of the proposed method Flock, we report the training and inference time per batch and peak GPU memory for Ultra, Trix, and variants of Flock on a single RTX A6000 (48 GB) in Tables˜6 and 7.
 Training.
@@
 As a result, pretraining typically takes about three days.
-One avenue for future work is to develop similarly highly optimized kernels for random-walk sampling to speed up the process.
+One avenue for future work is to develop similarly highly optimized kernels for random-walk sampling.
 Inference.
-Additionally, we report the inference results in Table 6, where we vary the number of walks $n$ and ensembled passes $P$.
+Additionally, we report the inference results in Table˜7, where we vary the number of walks $n$ and ensembled passes $P$.
 We observe near-linear growth of latency and VRAM with $n$, reflecting the dominant costs of walk sampling and sequence processing.
@@
 Appendix F Noise injection over existing KGFMs
-Setup.
-To answer Q5, we apply noise injection over the best performing KGFMs baselines Trix.
-Specifically, in each forward pass, we subtract an element-wise noise sampled from a uniform distribution $\epsilon\sim{\mathcal{U}}[0,0.5]$ to all relation and entity embeddings after the initialization stage.
-Note that the addition of noise will technically break deterministic node-relation equivariance, but the resulting model (Trix + noise) still respects probabilistic node-relation equivariance.
-We then pretrain Trix using the same experimental setup shown in Section 5.2, and compare with Trix without noise injection and Flock.
-To minimize the variance induced by injected noise and to ensure a fair comparison, we report ensembled prediction results with 16 samples for both Trix + noise and Flock.
-(a) Zero-shot performance for entity prediction.
+Table 8: Noise injection over the best performing KGFM baseline Trix.
+(a) Zero-shot entity prediction.
 MRR Hits@10
@@
 Flock 0.391 0.560
-(b) Zero-shot performance for relation prediction.
+(b) Zero-shot relation prediction.
 MRR Hits@1
@@
 Flock 0.881 0.817
+(c) Accuracy on Petals.
+Accuracy
+Trix 50%
++ noise 52%
+Flock 100%
+Setup.
+Since noise injection is a possible way to build a probabilistic equivariant KGFM in a different way from our approach (Gao et al., 2023), it is natural to ask how such KGFMs would perform compared to Flock.
+To answer this question, we apply noise injection over the best performing KGFMs baselines Trix.
+Specifically, in each forward pass, we add element-wise noise sampled from a uniform distribution $\epsilon\sim{\mathcal{U}}[-0.5,0]$ to all relation and entity embeddings after the initialization stage.
+Note that the addition of noise technically breaks deterministic node-relation equivariance, but the resulting model (Trix + noise) still respects probabilistic node-relation equivariance.
+We then pretrain Trix using the same experimental setup shown in Section˜5.2, and compare with Trix without noise injection and Flock.
+To minimize the variance induced by injected noise and to ensure a fair comparison, we report ensembled prediction results with 16 samples for both Trix + noise and Flock.
+This is a strong baseline implementing the ideas of prior work on noise injection and test-time ensembling for message passing networks on KGs (Lee et al., 2023; Gao et al., 2023).
 Results.
-We report the average zero-shot performance for entity prediction and relation prediction over 54 KGs in Tables 7(a) and 7(b), respectively.
+We report the average zero-shot performance for entity prediction and relation prediction over 54 KGs in Tables˜8(a) and 8(b), respectively, as well as trained performance for Petals in Table˜8(c).
 Across all tasks, Trix with naive noise injection fails to close the gap between Flock.
@@
 Together, these findings suggest that simply adding structure-agonistic noise is insufficient; performance gains only arise when stochasticity is topology-aware and is induced from the graph structure in a principled way.
-Appendix G Further experimental details
-Table 8: Zero-shot inductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
-Dataset Ultra Trix Flock
-MRR Hits@10 MRR Hits@10 MRR Hits@10
-Inductive $e,r$
-FB-25 0.388 0.640 0.393 0.650 0.404 0.664
-FB-50 0.338 0.543 0.334 0.547 0.352 0.566
-FB-75 0.403 0.604 0.401 0.611 0.418 0.622
-FB-100 0.449 0.642 0.436 0.635 0.452 0.663
-WK-25 0.316 0.532 0.305 0.496 0.280 0.491
-WK-50 0.166 0.324 0.166 0.313 0.136 0.278
-WK-75 0.365 0.537 0.368 0.513 0.382 0.538
-WK-100 0.164 0.286 0.188 0.299 0.187 0.304
-NL-0 0.342 0.523 0.385 0.549 0.381 0.606
-NL-25 0.395 0.569 0.377 0.589 0.345 0.590
-NL-50 0.407 0.570 0.404 0.548 0.366 0.565
-NL-75 0.368 0.547 0.351 0.525 0.311 0.524
-NL-100 0.471 0.651 0.486 0.676 0.452 0.692
-MT1 tax 0.224 0.305 0.358 0.452 0.282 0.383
-MT1 health 0.298 0.374 0.376 0.457 0.385 0.481
-MT2 org 0.095 0.159 0.091 0.156 0.100 0.163
-MT2 sci 0.258 0.354 0.323 0.465 0.318 0.458
-MT3 art 0.259 0.402 0.284 0.441 0.301 0.466
-MT3 infra 0.619 0.755 0.655 0.797 0.684 0.821
-MT4 sci 0.274 0.449 0.290 0.460 0.301 0.463
-MT4 health 0.624 0.737 0.677 0.775 0.680 0.780
-Metafam 0.238 0.644 0.341 0.815 0.476 0.935
-FBNELL 0.485 0.652 0.473 0.660 0.502 0.700
-Inductive $e$
-FB-v1 0.498 0.656 0.515 0.682 0.500 0.697
-FB-v2 0.512 0.700 0.525 0.730 0.535 0.737
-FB-v3 0.491 0.654 0.501 0.669 0.511 0.685
-FB-v4 0.486 0.677 0.493 0.687 0.505 0.702
-WN-v1 0.648 0.768 0.699 0.791 0.698 0.803
-WN-v2 0.663 0.765 0.678 0.781 0.696 0.790
-WN-v3 0.376 0.476 0.418 0.541 0.467 0.608
-WN-v4 0.611 0.705 0.648 0.723 0.653 0.729
-NL-v1 0.785 0.913 0.806 0.898 0.658 0.863
-NL-v2 0.526 0.707 0.569 0.768 0.588 0.797
-NL-v3 0.515 0.702 0.558 0.743 0.590 0.783
-NL-v4 0.479 0.712 0.538 0.765 0.555 0.786
-HM 1k 0.059 0.092 0.072 0.128 0.069 0.119
-HM 3k 0.037 0.077 0.069 0.119 0.067 0.118
-HM 5k 0.034 0.071 0.062 0.110 0.064 0.116
-HM Indigo 0.440 0.648 0.436 0.645 0.423 0.638
-ILPC Small 0.302 0.443 0.303 0.455 0.309 0.459
-ILPC Large 0.290 0.424 0.307 0.428 0.318 0.438
-Table 9: Zero-shot transductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
-Dataset Ultra Trix Flock
-MRR Hits@10 MRR Hits@10 MRR Hits@10
-Transductive
-NELL995 0.406 0.543 0.472 0.629 0.494 0.655
-NELL23k 0.239 0.408 0.290 0.497 0.233 0.398
-WDsinger 0.382 0.498 0.511 0.609 0.410 0.528
-ConceptNet100k 0.082 0.162 0.193 0.345 0.248 0.453
-CoDEx Small 0.472 0.667 0.472 0.670 0.441 0.644
-CoDEx Large 0.338 0.469 0.335 0.469 0.342 0.464
-YAGO310 0.451 0.615 0.409 0.627 0.414 0.674
-AristoV4 0.182 0.282 0.181 0.286 0.308 0.443
-DBpedia100k 0.398 0.576 0.426 0.603 0.450 0.627
-Hetionet 0.257 0.379 0.279 0.420 0.246 0.371
-FB15k-237(10) 0.248 0.398 0.246 0.393 0.246 0.402
-FB15k-237(20) 0.272 0.436 0.269 0.430 0.273 0.444
-FB15k-237(50) 0.324 0.526 0.321 0.521 0.319 0.518
-Table 10: Finetuned inductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
-Dataset Ultra Trix Flock
-MRR Hits@10 MRR Hits@10 MRR Hits@10
-Inductive $e,r$
-FB-25 0.383 0.635 0.393 0.650 0.405 0.666
-FB-50 0.334 0.538 0.334 0.547 0.357 0.570
-FB-75 0.400 0.598 0.401 0.611 0.425 0.630
-FB-100 0.444 0.643 0.436 0.633 0.460 0.668
-WK-25 0.321 0.535 0.300 0.493 0.298 0.506
-WK-50 0.140 0.280 0.166 0.313 0.127 0.260
-WK-75 0.380 0.530 0.368 0.513 0.405 0.556
-WK-100 0.168 0.286 0.188 0.299 0.187 0.306
-NL-0 0.329 0.551 0.385 0.549 0.418 0.619
-NL-25 0.407 0.596 0.377 0.589 0.405 0.626
-NL-50 0.418 0.595 0.405 0.555 0.391 0.562
-NL-75 0.374 0.570 0.351 0.525 0.344 0.544
-NL-100 0.458 0.684 0.482 0.691 0.486 0.714
-MT1 tax 0.330 0.459 0.397 0.508 0.413 0.497
-MT1 health 0.380 0.467 0.376 0.457 0.394 0.493
-MT2 org 0.104 0.170 0.098 0.162 0.107 0.174
-MT2 sci 0.311 0.451 0.331 0.526 0.366 0.525
-MT3 art 0.306 0.473 0.289 0.441 0.330 0.483
-MT3 infra 0.657 0.807 0.672 0.810 0.709 0.838
-MT4 sci 0.303 0.478 0.305 0.482 0.324 0.509
-MT4 health 0.704 0.785 0.702 0.785 0.711 0.790
-Metafam 0.997 1.000 0.997 1.000 0.992 1.000
-FBNELL 0.481 0.661 0.478 0.655 0.531 0.714
-Inductive $e$
-FB-v1 0.509 0.670 0.515 0.682 0.549 0.721
-FB-v2 0.524 0.710 0.525 0.730 0.553 0.754
-FB-v3 0.504 0.663 0.501 0.669 0.528 0.696
-FB-v4 0.496 0.684 0.493 0.687 0.510 0.702
-WN-v1 0.685 0.793 0.705 0.798 0.715 0.811
-WN-v2 0.679 0.779 0.682 0.780 0.702 0.795
-WN-v3 0.411 0.546 0.425 0.543 0.494 0.627
-WN-v4 0.614 0.720 0.650 0.722 0.665 0.741
-NL-v1 0.757 0.878 0.804 0.899 0.762 0.928
-NL-v2 0.575 0.761 0.571 0.764 0.612 0.806
-NL-v3 0.563 0.755 0.571 0.759 0.606 0.803
-NL-v4 0.469 0.733 0.551 0.772 0.572 0.801
-HM 1k 0.042 0.100 0.072 0.128 0.071 0.153
-HM 3k 0.030 0.090 0.069 0.119 0.067 0.153
-HM 5k 0.025 0.068 0.074 0.118 0.061 0.130
-HM Indigo 0.432 0.639 0.436 0.645 0.418 0.633
-ILPC Small 0.303 0.453 0.303 0.455 0.305 0.454
-ILPC Large 0.308 0.431 0.310 0.431 0.320 0.441
-Table 11: Finetune transductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
-Dataset Ultra Trix Flock
-MRR Hits@10 MRR Hits@10 MRR Hits@10
-Pretrained
-FB15k-237 0.368 0.564 0.366 0.559 0.343 0.532
-WN18RR 0.480 0.614 0.514 0.611 0.550 0.656
-CoDEx Medium 0.372 0.525 0.365 0.521 0.351 0.496
-Transductive
-NELL995 0.509 0.660 0.506 0.648 0.531 0.665
-NELL23k 0.268 0.450 0.306 0.536 0.280 0.465
-WDsinger 0.417 0.526 0.502 0.620 0.435 0.543
-ConceptNet100k 0.310 0.529 0.340 0.564 0.352 0.580
-CoDEx Small 0.490 0.686 0.484 0.676 0.463 0.648
-CoDEx Large 0.343 0.478 0.348 0.481 0.342 0.467
-YAGO310 0.557 0.710 0.541 0.702 0.552 0.700
-AristoV4 0.343 0.496 0.345 0.499 0.383 0.523
-DBpedia100k 0.436 0.603 0.457 0.619 0.470 0.623
-Hetionet 0.399 0.538 0.394 0.534 0.314 0.465
-FB15k-237(10) 0.254 0.411 0.253 0.408 0.260 0.420
-FB15k-237(20) 0.274 0.445 0.273 0.441 0.284 0.459
-FB15k-237(50) 0.325 0.528 0.322 0.522 0.317 0.517
-Table 12: Zero-shot inductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock. Bold indicates the best score per row.
-Dataset Ultra Trix Flock
-MRR Hits@1 MRR Hits@1 MRR Hits@1
-Inductive $e,r$
-FB-25 0.687 0.565 0.805 0.724 0.895 0.839
-FB-50 0.696 0.575 0.780 0.699 0.880 0.820
-FB-75 0.698 0.555 0.822 0.747 0.903 0.844
-FB-100 0.830 0.728 0.921 0.880 0.962 0.938
-WK-25 0.857 0.760 0.881 0.823 0.952 0.929
-WK-50 0.865 0.793 0.868 0.818 0.921 0.882
-WK-75 0.911 0.875 0.916 0.883 0.962 0.944
-WK-100 0.887 0.812 0.907 0.869 0.963 0.937
-NL-0 0.632 0.502 0.658 0.519 0.714 0.574
-NL-25 0.688 0.562 0.742 0.614 0.729 0.632
-NL-50 0.680 0.569 0.755 0.636 0.813 0.728
-NL-75 0.795 0.692 0.788 0.699 0.833 0.756
-NL-100 0.743 0.564 0.884 0.796 0.939 0.889
-MT1 tax 0.985 0.976 0.975 0.958 0.998 0.997
-MT1 health 0.721 0.561 0.973 0.949 0.991 0.983
-MT2 org 0.974 0.951 0.986 0.973 0.991 0.984
-MT2 sci 0.976 0.961 0.964 0.941 0.995 0.992
-MT3 art 0.881 0.798 0.885 0.825 0.944 0.907
-MT3 infra 0.962 0.935 0.940 0.905 0.989 0.980
-MT4 sci 0.933 0.891 0.966 0.944 0.974 0.957
-MT4 health 0.826 0.719 0.937 0.898 0.990 0.983
-Metafam 0.124 0.000 0.291 0.011 0.490 0.223
-FBNELL 0.700 0.564 0.726 0.605 0.833 0.737
-Inductive $e$
-FB-v1 0.646 0.523 0.705 0.599 0.814 0.723
-FB-v2 0.695 0.570 0.713 0.590 0.847 0.761
-FB-v3 0.679 0.553 0.742 0.644 0.860 0.780
-FB-v4 0.638 0.488 0.766 0.665 0.873 0.799
-WN-v1 0.836 0.740 0.792 0.613 0.924 0.858
-WN-v2 0.853 0.790 0.764 0.572 0.924 0.863
-WN-v3 0.707 0.577 0.741 0.568 0.937 0.888
-WN-v4 0.860 0.803 0.764 0.570 0.937 0.886
-NL-v1 0.636 0.358 0.657 0.453 0.862 0.731
-NL-v2 0.742 0.652 0.780 0.696 0.893 0.855
-NL-v3 0.669 0.544 0.725 0.612 0.815 0.731
-NL-v4 0.606 0.489 0.794 0.691 0.868 0.807
-ILPC Small 0.905 0.843 0.919 0.872 0.955 0.921
-ILPC Large 0.875 0.799 0.894 0.829 0.948 0.908
-HM 1k 0.626 0.447 0.663 0.414 0.687 0.500
-HM 3k 0.592 0.439 0.664 0.418 0.714 0.549
-HM 5k 0.605 0.452 0.672 0.428 0.746 0.593
-HM Indigo 0.681 0.559 0.852 0.765 0.956 0.921
-Table 13: Zero-shot transductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock.
-Dataset Ultra Trix Flock
-MRR Hits@1 MRR Hits@1 MRR Hits@1
-Transductive
-NELL995 0.583 0.437 0.578 0.457 0.684 0.555
-NELL23k 0.669 0.548 0.756 0.657 0.831 0.762
-WDsinger 0.668 0.546 0.720 0.621 0.823 0.738
-ConceptNet100k 0.181 0.083 0.650 0.469 0.795 0.658
-CoDExSmall 0.900 0.820 0.961 0.935 0.982 0.970
-CoDExLarge 0.892 0.824 0.902 0.837 0.973 0.950
-YAGO310 0.646 0.403 0.783 0.598 0.971 0.943
-AristoV4 0.254 0.201 0.389 0.265 0.597 0.496
-DBpedia100k 0.650 0.509 0.717 0.582 0.919 0.861
-Hetionet 0.634 0.524 0.809 0.707 0.940 0.890
-FB15k-237(10) 0.688 0.550 0.795 0.711 0.918 0.876
-FB15k-237(20) 0.695 0.558 0.834 0.758 0.952 0.923
-FB15k-237(50) 0.717 0.591 0.876 0.812 0.968 0.946
-Table 14: Finetuned inductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock. Bold indicates the best score per row.
-Dataset Ultra Trix Flock
-MRR Hits@1 MRR Hits@1 MRR Hits@1
-Inductive $e,r$
-FB-25 0.684 0.563 0.805 0.724 0.909 0.857
-FB-50 0.696 0.575 0.780 0.699 0.881 0.820
-FB-75 0.754 0.638 0.822 0.699 0.911 0.854
-FB-100 0.851 0.769 0.921 0.880 0.965 0.939
-WK-25 0.897 0.834 0.905 0.860 0.968 0.954
-WK-50 0.865 0.793 0.881 0.840 0.925 0.876
-WK-75 0.911 0.875 0.937 0.910 0.965 0.948
-WK-100 0.924 0.879 0.916 0.885 0.970 0.946
-NL-0 0.632 0.502 0.655 0.518 0.731 0.602
-NL-25 0.737 0.622 0.709 0.606 0.757 0.634
-NL-50 0.808 0.704 0.774 0.683 0.814 0.721
-NL-75 0.795 0.678 0.790 0.671 0.848 0.774
-NL-100 0.803 0.678 0.885 0.793 0.937 0.887
-MT1 tax 0.990 0.984 0.995 0.990 0.999 0.998
-MT1 health 0.929 0.867 0.973 0.949 0.994 0.988
-MT2 org 0.981 0.963 0.987 0.978 0.994 0.988
-MT2 sci 0.977 0.961 0.990 0.984 0.995 0.992
-MT3 art 0.907 0.851 0.887 0.828 0.950 0.916
-MT3 infra 0.966 0.947 0.970 0.952 0.996 0.993
-MT4 sci 0.954 0.929 0.972 0.952 0.983 0.968
-MT4 health 0.951 0.919 0.986 0.979 0.995 0.991
-Metafam 0.368 0.112 0.265 0.024 0.997 0.995
-FBNELL 0.720 0.576 0.766 0.639 0.879 0.801
-Inductive $e$
-FB-v1 0.650 0.513 0.705 0.599 0.855 0.766
-FB-v2 0.675 0.547 0.713 0.590 0.887 0.812
-FB-v3 0.677 0.556 0.742 0.644 0.879 0.810
-FB-v4 0.690 0.560 0.766 0.665 0.884 0.807
-WN-v1 0.844 0.754 0.776 0.591 0.926 0.879
-WN-v2 0.834 0.766 0.765 0.574 0.927 0.869
-WN-v3 0.707 0.577 0.756 0.594 0.950 0.911
-WN-v4 0.861 0.795 0.804 0.651 0.943 0.898
-NL-v1 0.719 0.504 0.590 0.341 0.883 0.766
-NL-v2 0.668 0.549 0.811 0.740 0.911 0.870
-NL-v3 0.646 0.484 0.757 0.643 0.868 0.795
-NL-v4 0.570 0.412 0.822 0.735 0.906 0.849
-ILPC Small 0.922 0.876 0.919 0.872 0.953 0.918
-ILPC Large 0.875 0.799 0.894 0.829 0.953 0.915
-HM 1k 0.626 0.447 0.663 0.414 0.756 0.561
-HM 3k 0.592 0.439 0.664 0.418 0.790 0.623
-HM 5k 0.605 0.452 0.672 0.428 0.744 0.591
-HM Indigo 0.726 0.614 0.835 0.746 0.946 0.903
-Table 15: Finetuned transductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock.
-Dataset Ultra Trix Flock
-MRR Hits@1 MRR Hits@1 MRR Hits@1
-Pretraining
-FB15k-237 0.795 0.709 0.924 0.870 0.976 0.957
-WN18RR 0.914 0.871 0.783 0.634 0.982 0.968
-CoDExMedium 0.919 0.870 0.931 0.886 0.974 0.952
-Transductive
-NELL995 0.630 0.513 0.578 0.457 0.713 0.584
-NELL23k 0.688 0.571 0.755 0.658 0.869 0.805
-WDsinger 0.730 0.603 0.721 0.627 0.885 0.815
-ConceptNet100k 0.612 0.488 0.712 0.551 0.885 0.813
-CoDExSmall 0.942 0.900 0.964 0.943 0.981 0.967
-CoDExLarge 0.907 0.850 0.908 0.845 0.973 0.950
-YAGO310 0.930 0.891 0.826 0.666 0.970 0.942
-AristoV4 0.254 0.201 0.498 0.381 0.651 0.547
-DBpedia100k 0.650 0.509 0.780 0.665 0.923 0.869
-Hetionet 0.737 0.646 0.922 0.862 0.942 0.897
-FB15k-237(10) 0.688 0.550 0.795 0.711 0.940 0.905
-FB15k-237(20) 0.695 0.558 0.846 0.778 0.958 0.931
-FB15k-237(50) 0.728 0.618 0.903 0.858 0.970 0.948
+Appendix G Case study: relation embedding on Metafam
+Figure 6: PCA of relation embeddings on Metafam.
+Ultra (Left) maps several inverse pairs (e.g., fatherOf vs.
+sonOf) to almost similar embeddings, where Flock (Right) yields clearly separated embeddings, indicating that its probabilistic equivariance allow Flock to distinguish between these semantically different relations, explaining its strong zero-shot performances.
+Setup.
+To further showcase why expressivity matters for zero-shot generalization, we present a case study on the Metafam dataset (Zhou et al., 2023).
+Metafam is built from a fixed family-relations ontology: during training, models observe edges with relations motherOf, fatherOf, daughterOf, sonOf, while the test queries only involve motherOf and fatherOf.
+Up to gender symmetries, this reduces to two effective predictive patterns (parent_of vs.
+child_of), and the test set focuses on a single one (parent_of).
+In the zero-shot setting, KGFMs cannot adapt to this ontology and must rely on their pretrained relation representations.
+Notice that here, Metafam is challenging: many relations are structurally similar (e.g., fatherOf, sonOf, sisterOf, nieceOf) yet encode opposite predictive patterns.
+Result.
+Figure˜6 shows that Ultra’s relation embeddings largely collapse these families, placing fatherOf and sonOf in almost identical positions in the PCA plane.
+This collapse makes it difficult to distinguish who is the parent and who is the child, leading to poor zero-shot performance.
+Flock, in contrast, can distinguish between these relations even if they are structurally similar, thanks to its random-walk sampling which introduce probabilistic equivariance on nodes and relations.
+As a results, Flock can produce distinct embeddings to fatherOf and sonOf and achieves much stronger zero-shot performance on Metafam.
+Appendix H Analysis of KG sparsity and performance
+Figure 7: The zero-shot entity prediction performance of Flock relative to Ultra, plotted against the densities of the 53 KGs.
+Performance of Flock and log-density of KGs have a Pearson correlation coefficient of -0.53 with p-value 5.0e-5, showing a statistically significant negative correlation.
+Setup.
+While Appendix˜G explains Flock’s high performance for Metafam, understanding its performances for other general KGs would be beneficial.
+Thus, we present an additional analysis on the 53 remaining KGs by identifying a structural property that is indicative of the performance of Flock.
+For the performance measure, we use the relative gain of Flock’s zero-shot entity prediction MRR compared to Ultra.
+For the structural property, we focus on density of KGs defined by $\frac{|E|}{|V|(|V|-1)}$ which affects the speed of random walks traversing all edges of a KG (intuitively because more edges means more time needed to traverse all of them), and hence is relevant in the context of our theory in Section˜4.2.
+We make the argument more grounded below.
+Result.
+Figure˜7 shows that Flock tends to perform well on sparse KGs, while less so on dense KGs, and the tendency is statistically significant.
+Interestingly, this agrees with our theoretical analysis in Section˜4.2, in which a necessary condition for universality is that the random walk covers all edges of a KG with a high probability (Proposition 4.1).
+The time taken until covering all edges is called the edge cover time, and it is known to be e.g., $O(|V||E|)$ for uniform walks (Zuckerman, 1991), which is proportional to the density of a graph.
+This suggests the performance of Flock is associated with the easiness to visit as many edges as possible rapidly, which is more challenging for dense KGs.
+This analysis is consistent with the observations of recent works on graph learning based on random walks, e.g., Wang & Cho (2024, Section 6) and Kim et al.
+(2025).
+Appendix I Further experimental details
 Datasets.
@@
 These include CoDEx Small, CoDEx Large (Safavi & Koutra, 2020), NELL-995 (Xiong et al., 2017), YAGO 310 (Mahdisoltani et al., 2015), WDsinger, NELL23k, FB15k-237(10), FB15k-237(20), FB15k-237(50) (Lv et al., 2020), AristoV4 (Chen et al., 2021), DBpedia100k (Ding et al., 2018), ConceptNet100k (Malaviya et al., 2020), and Hetionet (Himmelstein et al., 2017).
-Full results of Section 5.2.
-Full tables of zero-shot inductive and transductive entity prediction results for Flock are presented in Table 8 and Table 9, and full tables of finetuned performance are given in Table 10 and Table 11.
-We further provide the complete zero-shot and finetuned relation prediction results in Table 14, Table 13, and Table 15.
-Table 19 presents the pretraining graph mix shown in Section 5.3.
-Finally, detailed hyperparameter settings can be found in Table 20, while the number of epochs used for finetuning is summarized in Table 22.
+Full results of Section˜5.2.
+Full tables of zero-shot entity prediction results are presented in Table˜9, and full tables of finetuned performance are given in Table˜10.
+We further provide the complete zero-shot and finetuned relation prediction results in Table˜11 and Table˜12.
+The dataset statistics are in Table˜13, Table˜14 and Table˜15.
+Table˜16 presents the pretraining graph mix shown in Section˜5.3.
+Finally, detailed hyperparameter settings can be found in Table˜17, Table˜18 and Table˜19.
 Training.
@@
 $\mathcal{L}_{\mathrm{rel}}^{\mathrm{adv}}$ $=-\log p(h,r,t)\;-\;\sum_{i=1}^{k}w^{\mathrm{rel}}_{i,\alpha}\,\log\!\big(1-p(h,r_{i}^{\prime},t)\big).$
+Table 9: Zero-shot entity prediction results. Bold indicates the best score per row.
+Dataset Ultra Trix Flock MRR Hits@10 MRR Hits@10 MRR Hits@10 Inductive $e,r$ FB-25 0.388 0.640 0.393 0.650 0.404 0.664 FB-50 0.338 0.543 0.334 0.547 0.352 0.566 FB-75 0.403 0.604 0.401 0.611 0.418 0.622 FB-100 0.449 0.642 0.436 0.635 0.452 0.663 WK-25 0.316 0.532 0.305 0.496 0.280 0.491 WK-50 0.166 0.324 0.166 0.313 0.136 0.278 WK-75 0.365 0.537 0.368 0.513 0.382 0.538 WK-100 0.164 0.286 0.188 0.299 0.187 0.304 NL-0 0.342 0.523 0.385 0.549 0.381 0.606 NL-25 0.395 0.569 0.377 0.589 0.345 0.590 NL-50 0.407 0.570 0.404 0.548 0.366 0.565 NL-75 0.368 0.547 0.351 0.525 0.311 0.524 NL-100 0.471 0.651 0.486 0.676 0.452 0.692 MT1 tax 0.224 0.305 0.358 0.452 0.282 0.383 MT1 health 0.298 0.374 0.376 0.457 0.385 0.481 MT2 org 0.095 0.159 0.091 0.156 0.100 0.163 MT2 sci 0.258 0.354 0.323 0.465 0.318 0.458 MT3 art 0.259 0.402 0.284 0.441 0.301 0.466 MT3 infra 0.619 0.755 0.655 0.797 0.684 0.821 MT4 sci 0.274 0.449 0.290 0.460 0.301 0.463 MT4 health 0.624 0.737 0.677 0.775 0.680 0.780 Metafam 0.238 0.644 0.341 0.815 0.476 0.935 FBNELL 0.485 0.652 0.473 0.660 0.502 0.700 Inductive $e$ FB-v1 0.498 0.656 0.515 0.682 0.500 0.697 FB-v2 0.512 0.700 0.525 0.730 0.535 0.737 FB-v3 0.491 0.654 0.501 0.669 0.511 0.685 FB-v4 0.486 0.677 0.493 0.687 0.505 0.702 WN-v1 0.648 0.768 0.699 0.791 0.698 0.803 WN-v2 0.663 0.765 0.678 0.781 0.696 0.790 WN-v3 0.376 0.476 0.418 0.541 0.467 0.608 WN-v4 0.611 0.705 0.648 0.723 0.653 0.729 NL-v1 0.785 0.913 0.806 0.898 0.658 0.863 NL-v2 0.526 0.707 0.569 0.768 0.588 0.797 NL-v3 0.515 0.702 0.558 0.743 0.590 0.783 NL-v4 0.479 0.712 0.538 0.765 0.555 0.786 HM 1k 0.059 0.092 0.072 0.128 0.069 0.119 HM 3k 0.037 0.077 0.069 0.119 0.067 0.118 HM 5k 0.034 0.071 0.062 0.110 0.064 0.116 HM Indigo 0.440 0.648 0.436 0.645 0.423 0.638 ILPC Small 0.302 0.443 0.303 0.455 0.309 0.459 ILPC Large 0.290 0.424 0.307 0.428 0.318 0.438 Transductive NELL995 0.406 0.543 0.472 0.629 0.494 0.655 NELL23k 0.239 0.408 0.290 0.497 0.233 0.398 WDsinger 0.382 0.498 0.511 0.609 0.410 0.528 ConceptNet100k 0.082 0.162 0.193 0.345 0.248 0.453 CoDEx Small 0.472 0.667 0.472 0.670 0.441 0.644 CoDEx Large 0.338 0.469 0.335 0.469 0.342 0.464 YAGO310 0.451 0.615 0.409 0.627 0.414 0.674 AristoV4 0.182 0.282 0.181 0.286 0.308 0.443 DBpedia100k 0.398 0.576 0.426 0.603 0.450 0.627 Hetionet 0.257 0.379 0.279 0.420 0.246 0.371 FB15k-237(10) 0.248 0.398 0.246 0.393 0.246 0.402 FB15k-237(20) 0.272 0.436 0.269 0.430 0.273 0.444 FB15k-237(50) 0.324 0.526 0.321 0.521 0.319 0.518
+Table 10: Finetuned entity prediction results. Bold indicates the best score per row.
+Dataset Ultra Trix Flock MRR Hits@10 MRR Hits@10 MRR Hits@10 Inductive $e,r$ FB-25 0.383 0.635 0.393 0.650 0.405 0.666 FB-50 0.334 0.538 0.334 0.547 0.357 0.570 FB-75 0.400 0.598 0.401 0.611 0.425 0.630 FB-100 0.444 0.643 0.436 0.633 0.460 0.668 WK-25 0.321 0.535 0.300 0.493 0.298 0.506 WK-50 0.140 0.280 0.166 0.313 0.127 0.260 WK-75 0.380 0.530 0.368 0.513 0.405 0.556 WK-100 0.168 0.286 0.188 0.299 0.187 0.306 NL-0 0.329 0.551 0.385 0.549 0.418 0.619 NL-25 0.407 0.596 0.377 0.589 0.405 0.626 NL-50 0.418 0.595 0.405 0.555 0.391 0.562 NL-75 0.374 0.570 0.351 0.525 0.344 0.544 NL-100 0.458 0.684 0.482 0.691 0.486 0.714 MT1 tax 0.330 0.459 0.397 0.508 0.413 0.497 MT1 health 0.380 0.467 0.376 0.457 0.394 0.493 MT2 org 0.104 0.170 0.098 0.162 0.107 0.174 MT2 sci 0.311 0.451 0.331 0.526 0.366 0.525 MT3 art 0.306 0.473 0.289 0.441 0.330 0.483 MT3 infra 0.657 0.807 0.672 0.810 0.709 0.838 MT4 sci 0.303 0.478 0.305 0.482 0.324 0.509 MT4 health 0.704 0.785 0.702 0.785 0.711 0.790 Metafam 0.997 1.000 0.997 1.000 0.992 1.000 FBNELL 0.481 0.661 0.478 0.655 0.531 0.714 Inductive $e$ FB-v1 0.509 0.670 0.515 0.682 0.549 0.721 FB-v2 0.524 0.710 0.525 0.730 0.553 0.754 FB-v3 0.504 0.663 0.501 0.669 0.528 0.696 FB-v4 0.496 0.684 0.493 0.687 0.510 0.702 WN-v1 0.685 0.793 0.705 0.798 0.715 0.811 WN-v2 0.679 0.779 0.682 0.780 0.702 0.795 WN-v3 0.411 0.546 0.425 0.543 0.494 0.627 WN-v4 0.614 0.720 0.650 0.722 0.665 0.741 NL-v1 0.757 0.878 0.804 0.899 0.762 0.928 NL-v2 0.575 0.761 0.571 0.764 0.612 0.806 NL-v3 0.563 0.755 0.571 0.759 0.606 0.803 NL-v4 0.469 0.733 0.551 0.772 0.572 0.801 HM 1k 0.042 0.100 0.072 0.128 0.071 0.153 HM 3k 0.030 0.090 0.069 0.119 0.067 0.153 HM 5k 0.025 0.068 0.074 0.118 0.061 0.130 HM Indigo 0.432 0.639 0.436 0.645 0.418 0.633 ILPC Small 0.303 0.453 0.303 0.455 0.305 0.454 ILPC Large 0.308 0.431 0.310 0.431 0.320 0.441 Transductive NELL995 0.509 0.660 0.506 0.648 0.531 0.665 NELL23k 0.268 0.450 0.306 0.536 0.280 0.465 WDsinger 0.417 0.526 0.502 0.620 0.435 0.543 ConceptNet100k 0.310 0.529 0.340 0.564 0.352 0.580 CoDEx Small 0.490 0.686 0.484 0.676 0.463 0.648 CoDEx Large 0.343 0.478 0.348 0.481 0.342 0.467 YAGO310 0.557 0.710 0.541 0.702 0.552 0.700 AristoV4 0.343 0.496 0.345 0.499 0.383 0.523 DBpedia100k 0.436 0.603 0.457 0.619 0.470 0.623 Hetionet 0.399 0.538 0.394 0.534 0.314 0.465 FB15k-237(10) 0.254 0.411 0.253 0.408 0.260 0.420 FB15k-237(20) 0.274 0.445 0.273 0.441 0.284 0.459 FB15k-237(50) 0.325 0.528 0.322 0.522 0.317 0.517 Pretrained FB15k-237 0.368 0.564 0.366 0.559 0.343 0.532 WN18RR 0.480 0.614 0.514 0.611 0.550 0.656 CoDEx Medium 0.372 0.525 0.365 0.521 0.351 0.496
+Table 11: Zero-shot relation prediction results. Bold indicates the best score per row.
+Dataset Ultra Trix Flock MRR Hits@1 MRR Hits@1 MRR Hits@1 Inductive $e,r$ FB-25 0.687 0.565 0.805 0.724 0.895 0.839 FB-50 0.696 0.575 0.780 0.699 0.880 0.820 FB-75 0.698 0.555 0.822 0.747 0.903 0.844 FB-100 0.830 0.728 0.921 0.880 0.962 0.938 WK-25 0.857 0.760 0.881 0.823 0.952 0.929 WK-50 0.865 0.793 0.868 0.818 0.921 0.882 WK-75 0.911 0.875 0.916 0.883 0.962 0.944 WK-100 0.887 0.812 0.907 0.869 0.963 0.937 NL-0 0.632 0.502 0.658 0.519 0.714 0.574 NL-25 0.688 0.562 0.742 0.614 0.729 0.632 NL-50 0.680 0.569 0.755 0.636 0.813 0.728 NL-75 0.795 0.692 0.788 0.699 0.833 0.756 NL-100 0.743 0.564 0.884 0.796 0.939 0.889 MT1 tax 0.985 0.976 0.975 0.958 0.998 0.997 MT1 health 0.721 0.561 0.973 0.949 0.991 0.983 MT2 org 0.974 0.951 0.986 0.973 0.991 0.984 MT2 sci 0.976 0.961 0.964 0.941 0.995 0.992 MT3 art 0.881 0.798 0.885 0.825 0.944 0.907 MT3 infra 0.962 0.935 0.940 0.905 0.989 0.980 MT4 sci 0.933 0.891 0.966 0.944 0.974 0.957 MT4 health 0.826 0.719 0.937 0.898 0.990 0.983 Metafam 0.124 0.000 0.291 0.011 0.490 0.223 FBNELL 0.700 0.564 0.726 0.605 0.833 0.737 Inductive $e$ FB-v1 0.646 0.523 0.705 0.599 0.814 0.723 FB-v2 0.695 0.570 0.713 0.590 0.847 0.761 FB-v3 0.679 0.553 0.742 0.644 0.860 0.780 FB-v4 0.638 0.488 0.766 0.665 0.873 0.799 WN-v1 0.836 0.740 0.792 0.613 0.924 0.858 WN-v2 0.853 0.790 0.764 0.572 0.924 0.863 WN-v3 0.707 0.577 0.741 0.568 0.937 0.888 WN-v4 0.860 0.803 0.764 0.570 0.937 0.886 NL-v1 0.636 0.358 0.657 0.453 0.862 0.731 NL-v2 0.742 0.652 0.780 0.696 0.893 0.855 NL-v3 0.669 0.544 0.725 0.612 0.815 0.731 NL-v4 0.606 0.489 0.794 0.691 0.868 0.807 ILPC Small 0.905 0.843 0.919 0.872 0.955 0.921 ILPC Large 0.875 0.799 0.894 0.829 0.948 0.908 HM 1k 0.626 0.447 0.663 0.414 0.687 0.500 HM 3k 0.592 0.439 0.664 0.418 0.714 0.549 HM 5k 0.605 0.452 0.672 0.428 0.746 0.593 HM Indigo 0.681 0.559 0.852 0.765 0.956 0.921 Transductive NELL995 0.583 0.437 0.578 0.457 0.684 0.555 NELL23k 0.669 0.548 0.756 0.657 0.831 0.762 WDsinger 0.668 0.546 0.720 0.621 0.823 0.738 ConceptNet100k 0.181 0.083 0.650 0.469 0.795 0.658 CoDExSmall 0.900 0.820 0.961 0.935 0.982 0.970 CoDExLarge 0.892 0.824 0.902 0.837 0.973 0.950 YAGO310 0.646 0.403 0.783 0.598 0.971 0.943 AristoV4 0.254 0.201 0.389 0.265 0.597 0.496 DBpedia100k 0.650 0.509 0.717 0.582 0.919 0.861 Hetionet 0.634 0.524 0.809 0.707 0.940 0.890 FB15k-237(10) 0.688 0.550 0.795 0.711 0.918 0.876 FB15k-237(20) 0.695 0.558 0.834 0.758 0.952 0.923 FB15k-237(50) 0.717 0.591 0.876 0.812 0.968 0.946
+Table 12: Finetuned relation prediction results. Bold indicates the best score per row.
+Dataset Ultra Trix Flock MRR Hits@1 MRR Hits@1 MRR Hits@1 Inductive $e,r$ FB-25 0.684 0.563 0.805 0.724 0.909 0.857 FB-50 0.696 0.575 0.780 0.699 0.881 0.820 FB-75 0.754 0.638 0.822 0.699 0.911 0.854 FB-100 0.851 0.769 0.921 0.880 0.965 0.939 WK-25 0.897 0.834 0.905 0.860 0.968 0.954 WK-50 0.865 0.793 0.881 0.840 0.925 0.876 WK-75 0.911 0.875 0.937 0.910 0.965 0.948 WK-100 0.924 0.879 0.916 0.885 0.970 0.946 NL-0 0.632 0.502 0.655 0.518 0.731 0.602 NL-25 0.737 0.622 0.709 0.606 0.757 0.634 NL-50 0.808 0.704 0.774 0.683 0.814 0.721 NL-75 0.795 0.678 0.790 0.671 0.848 0.774 NL-100 0.803 0.678 0.885 0.793 0.937 0.887 MT1 tax 0.990 0.984 0.995 0.990 0.999 0.998 MT1 health 0.929 0.867 0.973 0.949 0.994 0.988 MT2 org 0.981 0.963 0.987 0.978 0.994 0.988 MT2 sci 0.977 0.961 0.990 0.984 0.995 0.992 MT3 art 0.907 0.851 0.887 0.828 0.950 0.916 MT3 infra 0.966 0.947 0.970 0.952 0.996 0.993 MT4 sci 0.954 0.929 0.972 0.952 0.983 0.968 MT4 health 0.951 0.919 0.986 0.979 0.995 0.991 Metafam 0.368 0.112 0.265 0.024 0.997 0.995 FBNELL 0.720 0.576 0.766 0.639 0.879 0.801 Inductive $e$ FB-v1 0.650 0.513 0.705 0.599 0.855 0.766 FB-v2 0.675 0.547 0.713 0.590 0.887 0.812 FB-v3 0.677 0.556 0.742 0.644 0.879 0.810 FB-v4 0.690 0.560 0.766 0.665 0.884 0.807 WN-v1 0.844 0.754 0.776 0.591 0.926 0.879 WN-v2 0.834 0.766 0.765 0.574 0.927 0.869 WN-v3 0.707 0.577 0.756 0.594 0.950 0.911 WN-v4 0.861 0.795 0.804 0.651 0.943 0.898 NL-v1 0.719 0.504 0.590 0.341 0.883 0.766 NL-v2 0.668 0.549 0.811 0.740 0.911 0.870 NL-v3 0.646 0.484 0.757 0.643 0.868 0.795 NL-v4 0.570 0.412 0.822 0.735 0.906 0.849 ILPC Small 0.922 0.876 0.919 0.872 0.953 0.918 ILPC Large 0.875 0.799 0.894 0.829 0.953 0.915 HM 1k 0.626 0.447 0.663 0.414 0.756 0.561 HM 3k 0.592 0.439 0.664 0.418 0.790 0.623 HM 5k 0.605 0.452 0.672 0.428 0.744 0.591 HM Indigo 0.726 0.614 0.835 0.746 0.946 0.903 Transductive NELL995 0.630 0.513 0.578 0.457 0.713 0.584 NELL23k 0.688 0.571 0.755 0.658 0.869 0.805 WDsinger 0.730 0.603 0.721 0.627 0.885 0.815 ConceptNet100k 0.612 0.488 0.712 0.551 0.885 0.813 CoDExSmall 0.942 0.900 0.964 0.943 0.981 0.967 CoDExLarge 0.907 0.850 0.908 0.845 0.973 0.950 YAGO310 0.930 0.891 0.826 0.666 0.970 0.942 AristoV4 0.254 0.201 0.498 0.381 0.651 0.547 DBpedia100k 0.650 0.509 0.780 0.665 0.923 0.869 Hetionet 0.737 0.646 0.922 0.862 0.942 0.897 FB15k-237(10) 0.688 0.550 0.795 0.711 0.940 0.905 FB15k-237(20) 0.695 0.558 0.846 0.778 0.958 0.931 FB15k-237(50) 0.728 0.618 0.903 0.858 0.970 0.948 Pretrained FB15k-237 0.795 0.709 0.924 0.870 0.976 0.957 WN18RR 0.914 0.871 0.783 0.634 0.982 0.968 CoDExMedium 0.919 0.870 0.931 0.886 0.974 0.952
 Table 13: Dataset statistics for inductive-$e,r$ link prediction datasets.
@@
 Hetionet 45158 24 2025177 112510 112510 $h/t$
-Table 19: Different graph pretraining mix shown in Section 5.3.
+Table 16: Different graph pretraining mix shown in Section˜5.3.
 1 2 3 4 5 6 8
@@
 AristoV4 $\checkmark$
-Table 20: Hyperparameter for Flock in pretraining and finetuning setup, for entity prediction and relation prediction.
+Table 17: Hyperparameter for Flock in pretraining and finetuning setups.
 Hyperparameter Entity prediction Relation prediction